### v4

# SDEs with critical general distributional drifts: sharp solvability and blow-ups

Abstract.
We establish weak well-posedness for SDEs having discontinuous diffusion coefficients and general distributional drifts that may introduce blow-up effects. Our drifts satisfy minimal assumptions, i.e. we assume only that the Cauchy problem for the Kolmogorov backward equation is well-posed in the standard Hilbert triple $W^{1,2}\hookrightarrow L^{2}\hookrightarrow W^{-1,2}$. By a result of Mazya and Verbitsky, these assumptions are precisely those drifts that can be represented as the sum of a form-bounded component (encompassing, for example, Morrey or Chang-Wilson-Wolff drifts) and a divergence-free distributional component in the ${\rm BMO}^{-1}$ space of Koch and Tataru.
We apply our results to finite particle systems with strong attracting interactions immersed in a turbulent flow. This includes particle systems of Keller-Segel type. Crucially, in dimensions $d\geq 3$, we cover almost the entire admissible range of attraction strengths, reaching nearly to the blow-up threshold.
As a further application of our results for SDEs and of the theory of Bessel processes, we obtain an improved upper bound on the constant in the many-particle Hardy inequality. Consequently, the lower bound previously derived by Hoffmann-Ostenhof, Hoffmann-Ostenhof, Laptev, and Tidblom is shown to be close to optimal.
Key words and phrases:
Stochastic differential equations, singular and distributional drifts, divergence-free drifts, form-boundedness, ${\rm BMO}^{-1}$, De Giorgi’s method, particle systems, Keller-Segel model, blow-ups
2020 Mathematics Subject Classification:
60H10, 47D07 (primary), 35J75 (secondary)
The research of D.K. is supported by NSERC grant (RGPIN-2024-04236)
section.2section.38section.46section.67section.74section.111section.142section.149appendix.312appendix.332appendix.339appendix.342appendix.354section*.357
1. Introduction
The subject of this paper is the stochastic differential equation (SDE)
$X_{t}=x-\int_{0}^{t}c(X_{s})ds+\sqrt{2}\int_{0}^{t}\sigma(X_{s})dB_{s},\quad x\in\mathbb{R}^{d},$ (1.1)
with critical general distributional drift $c$ and discontinuous diffusion coefficients $\sigma$ (see Section 5 and Section 7 for our precise setting). Here $\{B_{t}\}_{t\geq 0}$ denotes a $d$-dimensional Brownian motion. SDEs with singular drifts arise in various physical models, for instance, the passive-tracer model, where a time-dependent drift $c$ is the velocity field obtained from the Navier–Stokes equations [MK], or finite-particle approximations of the Keller–Segel model of chemotaxis [CP, FJ]. In the latter case one observes blow-ups, i.e. when even the weak existence for SDE (1.1) fails once one replaces $c$ by $(1+\varepsilon)c$, $\varepsilon>0$, in which case all particles collide a.s. in finite time and can stay “stuck” indefinitely. Both of these models are within the scope of the present work.
We postpone a survey of the recent literature on SDEs with singular drifts (much of it produced in the past decade) until after we have stated and discussed our assumptions on $c$.
Our main focus in this paper is on general drifts, i.e. not satisfying any special structural conditions such as control of the sign of ${\rm div\,}c$. So, on the one hand, our condition ($A_{1}$)-($A_{4}$) (or, equivalently, (1.6)) on $c$ includes the drifts with sufficiently small Morrey norm
$\|c\|_{M_{2+\varepsilon}}=\sup_{r>0,x\in\mathbb{R}^{d}}r\biggl(\frac{1}{|B_{r}|}\int_{B_{r}(x)}|c|^{2+\varepsilon}dx\biggr)^{\frac{1}{2+\varepsilon}},$ ($M_{2+\varepsilon}$)
or, more generally, the drifts in the Chang-Wilson-Wolff class: for some $\alpha>0$,
$\sup_{r>0,x\in\mathbb{R}^{d}}\frac{1}{|B_{r}|}\int_{B_{r}(x)}|c|^{2}\,r^{2}\bigl(1+(\log^{+}|c|^{2}\,r^{2})^{1+\alpha}\big)\text{ is sufficiently small}$ (1.2)
(actually, the definition is more general, see (3.3)). This, for example, provides us with flexible means to construct interaction kernels in particle systems. Crucially, we almost reach the blow-up threshold for $c$, i.e. multiplying $c$ by $1+\varepsilon$ takes us out of the weak existence regime. On the other hand, the same condition ($A_{1}$)-($A_{4}$) $\Leftrightarrow$ (1.6) includes drifts arising in some of the physical models mentioned above, notably of the form
[…]
where $\nabla$ is the row-divergence operator (see notations in Section 2). The class $\mathbf{BMO}^{-1}$ consists of divergence-free vector fields. It was identified by Koch-Tataru [KT] as a large class of initial conditions for which one can prove the existence and uniqueness of mild solution to 3D Navier-Stokes equations, and which provides natural scale and translation invariant version of $L^{2}$ boundedness of this solution. This class contains divergence-free drifts with entries in Besov space of distributions:
${\rm div\,}c=0\quad\text{ and }\quad c^{i}\in B^{-1+d/p}_{p,\infty}$
for some $p>d$, or Borel measurable divergence-free drifts with entries in the largest scaling-invariant Morrey class $M_{1}$:
${\rm div\,}c=0\quad\text{ and }\quad\langle|c|\mathbf{1}_{B_{r}(x)}\rangle\leq Cr^{d-1}$
for constant $C$ independent of $r$ or $x\in\mathbb{R}^{d}$. One particular example is
$c(x)=\biggl(\frac{x_{2}}{x_{1}^{2}+x_{2}^{2}},\frac{-x_{1}}{x_{1}^{2}+x_{2}^{2}},0,\dots,0\biggr),\quad x\in\mathbb{R}^{d},$ (1.3)
that, evidently, does not belong $L^{2}_{{\rm loc}}]^{d}$.
The diffusion coefficients $\sigma$ in Theorem 7.1 can have critical discontinuities. This in principle allows us to handle some systems of particles that interact via diffusion coefficients (Section 7).
Example 1.1 (Brownian particles in a turbulent flow).
We test our results for general SDE (1.1) against some interacting particle systems. That is, we aim at describing the dynamics of $N$ Brownian particles in $\mathbb{R}^{d}$, $d\geq 3$, subject to strong pair-wise attraction, while being advected by a turbulent flow whose divergence-free distributional velocity field belongs to $\mathbf{BMO}^{-1}(\mathbb{R}^{d})$. To this end, we work in $\mathbb{R}^{dN}$ where $N$ is large, and consider SDE
$X_{t}=x_{0}-\int_{0}^{t}c(X_{s})ds+\sqrt{2}B_{t},\quad x_{0}=(x_{0}^{1},\dots,x_{0}^{N})\in\mathbb{R}^{dN},$ (1.4)
where $X_{t}=(X_{t}^{1},\dots,X_{t}^{N})$, $X_{t}^{i}$ is the position of the $i$-th particle at time $t$, $B_{t}=(B_{t}^{1},\dots,B_{t}^{N})$, $\{B_{t}^{i}\}_{t\geq 0}$, $(i=1,\dots,N)$ are independent Brownian motions in $\mathbb{R}^{d}$. We further take
$c=b+q$
where:
  1. The first drift $b:\mathbb{R}^{dN}\rightarrow\mathbb{R}^{dN}$ is given component-wise by
    $b^{i}(x^{1},\dots,x^{N}):=\frac{1}{N}\sum_{j=1,j\neq i}^{N}\sqrt{\kappa}\frac{d-2}{2}e_{i,j}(x)\frac{x^{i}-x^{j}}{|x^{i}-x^{j}|^{2}},\quad x^{i},x^{j}\in\mathbb{R}^{d},$ (1.5)
    where $e_{ij}\in L^{\infty}(\mathbb{R}^{dN})$, $\|e_{ij}\|_{\infty}\leq 1$. Setting $e_{ij}=1$ introduces attraction between the particles arising, for example, in the finite particle approximation of the Keller-Segel model of chemotaxis [^1] In this example we assume $d\geq 3$, but will be able to include the case $d=2$ as well, albeit, at the moment, with additional conditions on the strength of attraction between the particles $\kappa>0$, see Appendix A..
  2. The second drift $q$ is described by external divergence-free velocity field,
    $q(x^{1},\dots,x^{N})=\big(q_{0}(x^{1}),\dots,q_{0}(x^{N})\big),\quad q_{0}\in\mathbf{BMO}^{-1}(\mathbb{R}^{d}),$
    so, it is easily seen, $q\in\mathbf{BMO}^{-1}(\mathbb{R}^{dN})$.
Either of our main results, Theorem 5.1 or Theorem 6.1 for $e_{ij}=1$, applies to $c=b+q$ and provide, in particular, weak existence and approximation uniqueness for particle system (1.4). Both theorems impose dimension-independent conditions, and so the resulting constraint on $\kappa$ does not degenerate as the number of particles $N$ goes to infinity. See Examples 5.1, 6.1 where we detail the particle system (1.4). In fact, in Example 6.1 we show that when all $e_{ij}=1$ Theorem 6.1 allows us to handle all $\kappa<16$ for all $N$, which is close to the blow-up threshold for (1.4), i.e. if $\kappa$ is greater than a constant that is slightly larger than 16, then all particles collide in finite time and stay glued to each other.
Section 4 compares our results (in the case $q_{0}=0$ and $e_{ij}=1$) with those of Cattiaux-Pédèches [CP], Fournier-Jourdain [FJ], Fournier-Tardy [FT] and Tardy [T] whose methods exploit the special structure of the drift (1.5) to obtain sharp, detailed results.
The search for the maximal admissible value of the strength of attraction $\kappa$ before a blow-up regime can be re-stated as the search for the minimal level of thermal excitation that prevents sticky collisions between the particles. The fact that the noise can turn local in time solutions into global ones can be viewed as another instance of regularization by noise (regarding the latter, see [Fl, FGP, FR]).
1.1. Result #1
In Theorem 5.1 and also in Theorem 7.1, we establish weak well-posedness of SDE (1.1) for the following class of $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ drifts:
$c=b+q,$ ($A_{1}$)
where $q$ is in general distribution-valued,
$q\in\mathbf{BMO}^{-1}\qquad(\Rightarrow\;\;{\rm div\,}q=0),$ ($A_{2}$)
and $b$ is a Borel-measurable drift satisfying
$b\in\mathbf{F}_{\delta},\quad\;\text{ i.e.}\;\quad|b|\in L^{2}_{{\rm loc}}\text{ and }\|b\varphi\|_{2}^{2}\leq\delta\|\nabla\varphi\|_{2}^{2}+c_{\delta}\|\varphi\|_{2}^{2}\quad\forall\,\varphi\in W^{1,2},$ ($A_{3}$)
for some finite constants $\delta$ (important) and $c_{\delta}$ (only its finiteness is important). (Here and below, $L^{2}_{{\rm loc}}=L^{2}_{{\rm loc}}(\mathbb{R}^{d})$ and $\|\cdot\|_{2}$ is the $L^{2}=L^{2}(\mathbb{R}^{d})$ norm, $W^{1,2}=W^{1,2}(\mathbb{R}^{d})$ is the Sobolev space of functions that are square integrable together with their first-order derivatives.) That is, $b$ is form-bounded. The last condition covers several important cases. It includes the Morrey class $M_{2+\varepsilon}$ (with form-bound $\delta$ depending on the Morrey norm $\|b\|_{M_{2+\varepsilon}}$), the larger Chang-Wilson-Wolff class recalled in (3.3) and, when we work in $\mathbb{R}^{dN}$, the many-particle drift (1.5) in Example 1.1 where one has $\delta=(N-1)^{2}N^{-2}\kappa$. A fuller discussion and more examples appear in Section 3.
The constant $\delta$ measures the size or the “strength” of singularities. In Theorem 5.1 we impose a completely dimension-free condition on $\delta$:
$\delta<4.$ ($A_{4}$)
A key consequence is that the corresponding restriction on the attraction parameter $\kappa=N^{2}(N-1)^{-2}\delta$ in Example 1.1 does not degenerate as the number of particles $N$ tends to infinity. (Theorem 6.1 relaxes this constraint on $\kappa$ by taking into account the divergence of $b$.)
Our proofs make use of specific properties of the classes $\mathbf{F}_{\delta}$ and $\mathbf{BMO}^{-1}$, such as the compensated compactness estimates and results on $\rm BMO$}-multipliers.
1.1.1. Optimality of our conditions
We claim that the class of drifts ($A_{1}$)-($A_{4}$) cannot be substantially enlarged. This is justified by the following two observations.
  1. Mazya and Verbitsky [MV] proved that conditions ($A_{1}$)–($A_{3}$) are equivalent to the estimate
    $|\langle c\cdot\nabla\varphi,\eta\rangle|\leq\alpha\|\nabla\varphi\|_{2}\|\nabla\eta\|_{2}$ (1.6)
    for all $\varphi,\eta\in C_{c}^{\infty}(\mathbb{R}^{d})$, for some constant $\alpha>0$ (assuming $c_{\delta}=0$ or up to replacing the homogeneous Sobolev spaces with their non-homogeneous counterparts, see Section 7). If $\alpha<1$, then inequality (1.6) ensures that the KLMN theorem [Ka, Ch. 6,§2] applies and so the Cauchy problem for the Kolmogorov backward equation is weakly well-posed in the standard Hilbert triple of Sobolev spaces $W^{1,2}\hookrightarrow L^{2}\hookrightarrow W^{-1,2}$. To the extent that one can view the latter as a minimal theory of the Kolmogorov backward equation, the present paper bridges the Eulerian and Lagrangian descriptions of diffusion:
[…]
    Here:
    1. $\bullet$
      Eulerian viewpoint: one studies averaged quantities, such as temperature or concentration, governed by the heat equation or, more generally, by transport-diffusion PDEs.
    2. $\bullet$
      Lagrangian viewpoint: one follows individual molecules whose trajectories solve SDEs (Brownian motion in the simplest case).
    Until quite recently, Lagrangian results required stronger regularity assumptions on the drift.
    We refer to the equivalence ($A_{1}$)-($A_{3}$) $\Leftrightarrow$ (1.6) as the “generalized form-boundedness”, although Mazya and Verbitsky [MV] call (1.6) simply the form-boundedness condition on $c$.
    Let us comment on why one would expect the Kolmogorov equation to be well-posed in the standard Hilbert triple. In fact, we can substantially relax the conditions on drift $b$ by requiring well-posedness of the Kolmogorov backward equation in the shifted triple of Bessel spaces $\mathcal{W}^{\frac{3}{2},2}\hookrightarrow\mathcal{W}^{\frac{1}{2},2}\hookrightarrow\mathcal{W}^{-\frac{1}{2},2}$, see Theorem A.1. Doing so, however, forces us to drop the distributional part $q$, lose at least for now the almost optimal dimension-independent condition ($A_{4}$) and precludes discontinuous diffusion coefficients. By contrast, under ($A_{1}$)-($A_{4}$) we can allow discontinuous diffusion coefficients $\sigma$ for which the associated non-divergence operator can still be rewritten in divergence form (Theorem 7.1). Such diffusion coefficients keep us within the standard Hilbert triple $W^{1,2}\hookrightarrow L^{2}\hookrightarrow W^{-1,2}$, as discussed in Section 7.
  2. Near-optimality of the bound ($A_{4}$). Consider the SDE
    $X_{t}=x-\sqrt{\delta}\frac{d-2}{2}\int_{0}^{t}\frac{X_{s}}{|X_{s}|^{2}}ds+\sqrt{2}B_{t},$ (1.7)
    where $B_{t}$ is the $d$-dimensional Brownian motion, $d\geq 3$. The drift here, $b(x)=\sqrt{\delta}\frac{d-2}{2}\frac{x}{|x|^{2}}$, belongs to $\mathbf{F}_{\delta}$, see Section 3 for a detailed explanation. Theorems 5.1 and 6.1 show that if ($A_{4}$) is satisfied, then there exists a strong Markov family of weak solutions to SDE (1.7), and these weak solutions are unique at least among those weak solutions that can be constructed via regularization of the drift. Moreover, in the critical case $\delta=4$ there is still a sufficiently rich theory of the corresponding Kolmogorov backward PDE, see Remark 5.1. On the other hand, taking advantage of the anti-symmetry of the drift, one can show that $R_{t}=|X_{t}|^{2}$ is a squared Bessel process (see, e.g. [BFGM]) and therefore:
    (a) If
    $\delta\geq 4\left(\frac{d}{d-2}\right)^{2},$
    then for every initial point solution arrives at the origin in finite time with probability $1$, and stays there indefinitely (that is, $R_{t}=0$ for all $t\geq\tau$ for some finite stopping time $\tau$). A simple argument shows that for such $\delta$ SDE (1.7) does not have a weak solution departing from $x=0$, see e.g. [BFGM].
    (b) If
    $4<\delta<4\left(\frac{d}{d-2}\right)^{2},$
    then solution still visits the origin infinitely many times, but does not stay there, i.e. $\int_{0}^{\infty}R_{t}dt<\infty$ a.s.
    In Section 4 we discuss similar counterexamples in the context of particle system introduced in Example 1.1.
Remark 1.1.
The fact that (1.6) implies ($A_{1}$)–($A_{3}$) is proved in [MV] by taking $b$ and $q$ from the Hodge-type decomposition of $c$:
$b:=-\nabla(1-\Delta)^{-1}{\rm div\,}c+(1-\Delta)^{-1}c,$ (1.8)
$q:=-(1-\Delta)^{-1}{\rm curl\,}c.$ (1.9)
Here, $b$ caputres the attractive or the repulsive part of the drift in the dynamics of $X_{t}$. However, much more is already known when the entire drift belongs to the form-bounded class $\mathbf{F}_{\delta}$: one has strong well-posedness for the SDE, as well as well-posedness for the associated stochastic transport equation, see Theorem 5.1. How far these results extend when a non-zero distributional component $q$ in $\mathbf{BMO}^{-1}$ is not yet clear. So, if we insist on the decomposition (1.8), (1.9) above, Theorem 5.1 would treat only gradient-type drifts, which we want to avoid.
1.1.2. Blow ups
In Section 6 we will also discuss drifts having critical (form-bounded) divergence, which allows to relax the assumptions on the drift to some super-critical conditions, i.e. passing to the small scales actually increases the norm of the drift. The sub-critical/critical/super-critical classification of the spaces of vector fields, widely used in the literature, is recalled in Appendix B. It should be added, however, that this classification is not so relevant to the main body of the present paper (except for Section 6) because:
(a) Our main focus is on general drifts, for which one only has the dichotomy sub-critical/critical.
(b) This classification does not distinguish between critical spaces that do, and those that do not, reach blow-ups.
(Regarding (b), for example, both $L^{d}$ space and weak $L^{d}$ space (we recall its definition in Section 2) are critical, but only the weak $L^{d}$ space contains drift $C\frac{x}{|x|^{2}}$ whose attracting singularity at the origin is strong enough to kill the weak well-posedness of the SDE if $C$ is too large. In other words, in the critical case, a lot depends on the definition of the norm of the drift.)
If a critical class of drifts is broad enough to contain blow-up examples, a well-posedness theorem must include a smallness condition on the drift norm, such as e.g. ($A_{4}$).
Compared with blow-ups for the Navier–Stokes equations, the mechanism of blow-ups in particle systems of the kind treated in Example 1.1 is much better understood [CP, CPZ, FJ, FT, JL]. That said, as was noted in [FJ], at the time of writing of their article there was still a substantial gap between (i) the drift singularities that the general theory of SDEs with singular drifts could handle and (ii) the even stronger singularities they themselves had to treat. One of the purposes of the present work, together with [KS2, K5, KS6], is to close this gap:
$\text{SDEs with general singular drift \qquad\qquad$\longrightarrow$ \qquad\qquad particle systems},$ (1.10)
i.e. to bring the general theory of SDEs “up to the task” so that it can handle blow-ups. See, in particular, Appendix A where we demonstrate how the weak well-posedness of the finite-particle Keller-Segel SDE in $\mathbb{R}^{2N}$ can be reached from an earlier result in [KS1] on SDEs with general drifts, albeit at the moment under additional rather restrictive conditions of the strength of attraction. Interestingly, the path that leads to this passes through non-local operators (Theorem A.1 and Corollary A.1).
The same connection (1.10) was already pursued by Krylov and Röckner [KrR], but with a different interaction kernel: its attractive part can be extremely singular, but it is always dominated on average (not pointwise) by the repulsive part, so no blow-up occurs.
1.2. Result #2
In Appendix B we will also discuss super-critical drifts (necessarily under additional assumptions on their divergence). Specifically, if the positive part of the divergence ${\rm div\,}b$ is a form-bounded potential, then this enables us to relax the form-boundedness condition on $b$ to a super-critical form-boundedness condition:
$|b|^{\frac{1+\nu}{2}}\in\mathbf{F}_{\delta},\quad\nu\in]0,1[,\quad\delta<\infty.$
In this case one still has weak existence for every initial point [KS2], moreover, as was mentioned there, one can combine such a drift with an ordinary form-bounded component, again reaching the blow-up threshold, and may also include discontinuous diffusion coefficients of the type treated in Section 7.
Generally speaking, in super-critical settings many standard regularity properties of the diffusion process are lost. Some of them can be saved, but to recover most of them one must impose additional critical conditions on the drift. In our framework, to establish e.g. the Markov property, Theorem 6.3 supplements the super-critical assumption with the multiplicative form-boundedness condition $b\in\mathbf{MF}_{\delta}$, i.e.
$\langle|b|\varphi,\varphi\rangle\leq\delta\|\nabla\varphi\|_{2}\|\varphi\|_{2}+c_{\delta}\|\varphi\|_{2}^{2},\quad\varphi\in W^{1,2},\quad\delta<\infty.$
This condition is critical and is substantially more general than $b\in\mathbf{F}_{\delta}$ because somehow it implicitly presumes the existence of ${\rm div\,}b$. In contrast to the form-bounded class $\mathbf{F}_{\delta}$, the multiplicative form-bounded class $\mathbf{MF}_{\delta}$ can be completely characterized in elementary terms: $b\in\mathbf{MF}_{\delta}$ if and only if
$\langle|b|\mathbf{1}_{B_{r}(x)}\rangle\leq Cr^{d-1}$
with constant $C$ independent of $r$ or $x\in\mathbb{R}^{d}$. The latter is the largest scaling-invariant Morrey [^2] For the theory of Morrey spaces and related estimates, see Adams-Xiao [AX] and Krylov [Kr6]. class $M_{1}$. In fact, we have
${\qquad|b|\in M_{1}}$${b\in\mathbf{MF}_{\delta}\text{ for some finite $\delta$ and }c_{\delta}=0}$${b\in\mathbf{BMO}^{-1}}$                      $\scriptstyle{\text{form-bounded }({\rm div\,}b)_{+}}$                                 $\scriptstyle{{\rm div\,}b=0}$
where the arrow $\rightarrow$ in this diagram means inclusion. (These inclusions were proved in [M, Theorem 1.4.7] and [MV2, Theorem V]. See discussion before Theorem 6.3 and the proof in Appendix D.) Choosing the inclusion into $\mathbf{MF}_{\delta}$ leads to an approach to studying the Kolmogorov backward equation based on “Caccioppoli’s iterations” for establishing the classical Caccioppoli’s inequality [KV], see Remark 6.1 for details; choosing the inclusion into $\mathbf{BMO}^{-1}$ amounts to absorbing the stream matrix $Q$ (for $b=\nabla Q$) into the diffusion coefficients and obtaining a Caccioppoli-type inequality from there [H, SSŠZ]. We discuss this in Section 6.
1.3. Result #3
As an application of our results on SDEs (Theorem 6.1) and of the theory of Bessel processes, we obtain an upper bound on the best possible constant $C_{d,N}$ in the many-particle Hardy inequality
$C_{d,N}\sum_{1\leq i<j\leq N}\int_{\mathbb{R}^{dN}}\frac{|\varphi(x)|^{2}}{|x^{i}-x^{j}|^{2}}dx\leq\int_{\mathbb{R}^{dN}}|\nabla\varphi(x)|^{2}dx\quad\forall\,\varphi\in C_{c}^{\infty}(\mathbb{R}^{dN})$
(Theorem 6.2). Our upper bound on $C_{d,N}$ improves the existing results by replacing factorial growth in the dimension with polynomial growth. It also shows that the lower bound on $C_{d,N}$ obtained earlier by Hoffmann-Ostenhof–Hoffmann-Ostenhof-Laptev-Tidblom [HHLT] is nearly optimal in high dimensions. To our knowledge, this is the first time a probabilistic argument is used in the analysis of the best possible constant in a Hardy inequality.
1.4. Literature
Let us now comment on the existing literature on SDEs with singular drifts, focusing mostly on general drifts. In our setting, even if we restrict our attention to Borel measurable drifts, then the Zvonkin transform is not applicable and our drifts are not of Girsanov type.
1) The present paper continues [KS2, K5, KS6] and also [KS7, KS1]. In [KS1], Semënov and the first-named author constructed weak solutions to SDEs whose drifts lie in a class even larger than $\mathbf{F}_{\delta}$, namely, the class of weakly form-bounded vector fields [K1, S]. This larger class contains, for example, the Morrey class $M_{1+\varepsilon}$, and therefore vector field (1.3); however, that result requires more restrictive condition $\delta<\frac{c}{d^{2}}$ (see Appendix A). In all those earlier works one has $q=0$ and $\sigma=I$, with the exception of [KS7] that dealt with the class of discontinuous diffusion coefficients treated in Section 7.
2) If we restrict attention to Borel-measurable drifts, the most closely related results are the recent works of Krylov [Kr1, Kr2, Kr3, Kr4, Kr5] and of Röckner and Zhao [RZ]. In [Kr1], Krylov treats a class of diffusion coefficients much larger than ours (his diffusion coefficients are in $\rm VMO$} class, or have small $\rm BMO$} norm), but a smaller class of drifts (namely, the Morrey class $M_{(\frac{d}{2}\vee 2)+\varepsilon}$). The irregular diffusion coefficients require estimates on second-order derivatives of solutions to the Kolmogorov backward equation, whereas we neither assume nor have such estimates under our assumptions on the drift. Krylov’s drifts can also produce blow-up phenomena, but he assumes that the size of the singularity, measured by the Morrey norm, remains below a small, dimension-dependent constant. By contrast, our goal is to reach the maximal admissible strength of singularity (in Example 1.1, the interaction strength $\kappa$; or the supremum (1.2)). Other recent results of Krylov are described at various points in this work.
3) Some techniques such as De Giorgi’s method also connect our paper to the works of Zhang-Zhao [ZZ1] and Hao-Zhang [HZ], which focus on super-critical divergence-free drifts (in [HZ], distributional). Some of their results deal with non-zero divergence, but they do not reach blow-ups. It should be added that in the study of Navier-Stokes equations (the SDEs connected to the N-S equations is the main interest of [ZZ1, HZ]) the blow ups of the type discussed above, i.e. attracting, do not appear.
4) There is rich literature on SDEs with general distributional drifts, also motivated by physical applications. In all the results we are aware of, the diffusion coefficients $\sigma$ have to be at least Hölder continuous. Many of these works treat time-inhomogeneous drifts, however, in this brief discussion we will specify these results to time-homogeneous drifts. Flandoli-Issoglio-Russo [FIR] and Zhang-Zhao [ZZ2] study drifts that belong to Bessel potential spaces with negative index and employ the Zvonkin transform to obtain weak well-posedness for SDE (1.1). Chaudru de Raynal-Menozzi [CM] proved, among other results, weak well-posedness with drift $b$ in Besov space $B_{p,q}^{\beta}$ with $\frac{1}{2}<\beta\leq 0$ and $p>d$. They addressed, in particular, the problem of defining the product of two distributions $b\cdot\nabla v$, where $v$ is a solution to the Kolmogorov backward equation. The latter, in turn, dictates their restriction $\beta>-\frac{1}{2}$. Earlier, Delarue-Diel [DD] and Cannizzaro-Chouk [CC] proved, in dimensions $d=1$ and $d\geq 2$, respectively, well-posedness of the martingale problem for general distributional drifts in Hölder-Besov space $B_{\infty,\infty}^{\beta}$ for all $\frac{2}{3}<\beta\leq 0$. In the more singular regime $\frac{2}{3}<\beta\leq-\frac{1}{2}$ they assume that the drift can be enhanced to a rough distribution, in order to apply the theory of paracontrolled distributions. This allowed them to consider random drifts of the form $b=\nabla h$, where $h$ is a solution of a KPZ-type equation, and thereby construct the polymer measure with white noise potential. As mentioned earlier, the class of divergence-free drifts $\mathbf{BMO}^{-1}$ contains drifts whose components lie in Besov space $B^{-1+d/p}_{p,\infty}$, $p>d$, i.e. the exponent $\beta=-1+d/p$ can go up to 1-1. This, however, cannot serve as a comparison with the previous cited results since by definition $\mathbf{BMO}^{-1}$ drifts are divergence-free, but this justifies to some extent why we do not address in this paper the problem of mutiplying distributions in the Kolmogorov equation, i.e. since our $\beta\not>-\frac{1}{2}$.
5) Let us also mention recent papers by Chaudru de Raynal-Jabir-Menozzi [CJM, CJM2] where the authors handle McKean-Vlasov SDEs with distributional Besov drifts. The regularization by convolution allows them to venture substantially farther in the assumptions on the drift, compared to the papers cited in 4). It is quite noteworthy, since they can start with a delta-function in the initial distribution.
The papers on general distributional drifts mentioned in 4) and 5) do not reach blow-ups, in contrast to our Theorems 5.1, 7.1, 6.1. That said, the cited works cover some other highly irregular drifts that fall outside the scope of our hypotheses.
Finally, we mention Bresch-Jabin-Wang [BJW] who consider gradient-form singular attracting drifts that allow blow-ups. Their focus is different, namely, it is quantitative propagation of chaos at the PDE level, but they are also interested in reaching the blow-up thresholds. Their class of drifts is not contained in our class, and vice versa, although there is a very substantial overlap between these classes.
6) Some singular drifts, as well as some degenerate diffusion coefficients, can be treated using the Ambrosio-Figalli-Trevisan superposition principle. The latter allows to conclude weak solvability of the SDE from some integrability and weak continuity properties of solution to the corresponding Fokker-Planck equation (see (A.7), (A.8)). See Trevisan [Tr] and Bogachev-Röckner-Shaposhnikov [BRS], see also Grube [Gru] and the survey of the literature therein. The theory of Fokker-Planck equations with locally unbounded drifts and degenerate diffusion coefficients is discussed in detail in [BKRS].
7) Regarding super-critical drifts, we refer again to Zhang-Zhao [ZZ1], as well as to recent papers by Hao-Zhang [HZ] and Gräfner-Perkowski [GP] where the authors consider super-critical distributional drifts. In [GP], the authors treat divergence-free Besov drifts $b\not\in B^{-1}_{2d(d+2)^{-1},2}$, provided that the initial density is absolutely continuous with respect to the Lebesgue measure; they also consider quite irregular attracting/repulsing component of the drift, although it does not reach the blow-ups.
Comprehensive surveys of the literature on SDEs with singular and distributional drifts can be found in [CM] and [HZ].
1.5. Main instruments: De Giorgi’s method, Trotter’s theorem and compensated compactness estimates
Since the drift $c$ satisfying (1.6) is in general distribution-valued ($c=b+q$) we have to give a proper meaning to term $c(X_{s})$ in SDE (5.1). We will do it in two ways:
(a) Theorem 5.1(iii): At the level of the martingale problem and with the Itô expansion, we have to define $(c\cdot\nabla v)(X_{s})$ for suitable test functions. The usual test functions in $C_{c}^{\infty}$ cannot be used, but it is still possible to find a sufficiently rich space of test functions (in particular, dense in $C_{\infty}$; see notations in Section 2) that will give us a continuous martingale. In fact, this space of test functions will be the domain of the Feller generator $\Lambda\supset-\Delta+c\cdot\nabla$.
(b) Theorem 5.1(iv): By constructing the “limiting drift” process (called “formal dynamics” in [CM]) for disperse initial data, assuming $\delta<1$. This is a result of Bass-Chen type [BC], see also Zhang-Zhao [ZZ2], Chaudru de Raynal-Menozzi [CM] and Hao-Zhang [HZ] who have similar results. The proof will use convergence of the martingale solutions of the approximating SDEs (5.2) that will follow as a by-product of (a).
To construct the sought Feller semigroup, we will have to employ some deep results from the operator theory, the theory of PDEs and harmonic analysis:
  1. Trotter’s approximation theorem,
  2. De Giorgi’s method in $L^{p}$ for $p>\frac{2}{2-\sqrt{\delta}}$, i.e. in the context of Example 1.1, $p$ will be an explicit function of the strength of attraction between the particles $\kappa$.
  3. Compensated compactness esimates and results on $\rm BMO$}-multipliers.
A crucial feature of Trotter’s approximation theorem (Theorem 8.1) is that it requires no a priori knowledge of the limiting object, i.e. in our case, the limiting Feller semigroup. By contrast, in some other of our results on singular SDEs (such as Theorem A.1) we first construct an explicit candidate for the Feller resolvent. This simplifies approximation arguments, but it also automatically provides strong gradient bounds that, in turn, introduce dimension-dependent restrictions on the form bound $\delta$. As the dimension grows, the admissible range of $\delta$ shrinks to the empty set, making problematic possible applications to many-particle systems that exist in spaces of very large dimension. (Example 5.1 explains how a bound on $\delta$ translates into a bound on the attraction strength $\kappa$.) Using De Giorgi’s method in the proof of Theorem 5.1 decouples gradient bounds from the bounds that provide the existence of the weak solution (i.e. tightness argument, construction of the Feller semigroup, etc). Consequently, we can impose dimension-free assumptions on $\delta$ (and on the entire drift), which is a crucial point for applications to many-particle systems.
Another way to view the relationship between the dimension-free assumptions on the drift and the Sobolev embedding is as follows. The Sobolev embedding becomes weaker as the dimension increases, and so if it is applied only once, then one arrives at dimension-dependent conditions on the drift. In De Giorgi’s method, however, the Sobolev embedding is applied infinitely many times, which allows to overcome this dependence on the dimension.
The present paper strengthens the results [K5, KS6] that dealt with Borel-measurable drifts. Allowing a non-zero distributional $q\neq 0$ drift, as in ($A_{1}$)-($A_{4}$), required substantial modifications of the arguments in those papers.
De Giorgi’s method for divergence-form operators with form-bounded and/or drift in $\mathbf{BMO}^{-1}$ has also been applied, in different contexts, by Hara [H] and Seregin-Silvestre-Šverak-Zlatoš [SSŠZ]. We also refer to recent paper by Liang-Wang-Zhao [LWZ] where the authors establish strong estimates on the modulus of continuity of solutions of elliptic equations with distributional coefficients.
1.6. Structure of the paper
  1. Section 4
    We begin with a preliminary discussion of blow-up thresholds for the particle system in 1.1. These thresholds both illustrate the sharpness of our SDE results and serve as input in Section 6, where they enter the proof of a sharper upper bound for the constant in the many-particle Hardy inequality (Theorem 6.2).
  2. Sections 5
    Theorem 5.1 is our main result for general drifts when the diffusion matrix is constant.
  3. Section 6
    Theorems 6.1 and 6.3 treat drifts whose divergence satisfies a critical (scaling-invariant) form-boundedness condition. Under this assumption we can relax the constraint on $|b|$, allowing some super-critical form-boundedness conditions. As an application, Theorem 6.2 gives an improved upper bound on the constant in the many particle Hardy inequality.
  4. Section 7
    Theorem 7.1 extends the analysis to diffusion matrices that may have critical discontinuities.
  5. Appendix A
    Corollary A.1 shows how well-posedness of the finite-particle Keller–Segel system can be derived from our earlier result on SDEs with general singular drifts (Theorem A.1). The proof of that theorem uses a more operator-theoretic approach, based on fractional resolvent representations, and covers the larger class of weakly form-bounded drifts that contains both the Morrey class $M_{1+\varepsilon}$ and the Kato class studied by Bass-Chen [BC].
  6. Appendix B
    discusses some super-critical drifts.
Acknowledgements
We are grateful to Galia Dafni, Jean-François Jabir and Kodjo Raphaël Madou for very useful discussions.
2. Notations and auxiliary results
Throughout this work we use the following notations.
1. $\mathcal{B}(X,Y)$ is the space of bounded linear operators $X\rightarrow Y$ between Banach spaces $X$, $Y$, endowed with the operator norm $\|\cdot\|_{X\rightarrow Y}$. Put $\mathcal{B}(X):=\mathcal{B}(X,X)$.
The space of $d$-dimensional vectors with entries in $X$ is denoted by $X]^{d}$. We reserve the upper index to denote the components $q^{i}\in X$ of $q\in[X]^{d}$.
The notation $\overset{w}{\rightarrow}$ in XX” stands for the weak convergence in $X$. Similarly for $X]^{d}$.
We write
$T=s\mbox{-}Y\mbox{-}\lim_{n}T_{n}$
for $T$, $T_{n}\in\mathcal{B}(X,Y)$ if
$\lim_{n}\|Tf-T_{n}f\|_{Y}=0\quad\text{ for every $f\in X$}.$
By $T\upharpoonright X$ we denote the restriction of operator $T$ to a subspace $X\subset D(T)$.
By $\big[T\upharpoonright X\big]_{Y\rightarrow Y}^{\rm clos}$ we denote the closure of the restriction $T\upharpoonright X$ (when it exists).
2. The space $L^{p}=L^{p}(\mathbb{R}^{d},dx)$, $W^{1,p}=W^{1,p}(\mathbb{R}^{d},dx)$ corresponds to the Lebesgue and to the Sobolev space, respectively. Let $L^{p}_{\rho}=L^{p}(\mathbb{R}^{d},\rho(x)dx)$ denote the weighted Lebesgue space with weight $\rho$ (defined by formula (2.1) below).
Set $\|\cdot\|_{p}:=\|\cdot\|_{L^{p}}$ and denote operator norm $\|\cdot\|_{p\rightarrow q}:=\|\cdot\|_{L^{p}\rightarrow L^{q}}$.
Given $1<p<\infty$, we set $p^{\prime}:=\frac{p}{p-1}$.
Put
$\langle f,g\rangle=\langle f\bar{g}\rangle:=\int_{\mathbb{R}^{d}}f\bar{g}dx.$
For vector fields $b$, $\mathsf{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, we put
$\langle b,\mathsf{f}\rangle:=\langle b\cdot\mathsf{f}\rangle\qquad\text{($\cdot$ is the scalar product in $\mathbb{R}^{d}$)}.$
$C_{c}$ (resp.  $C_{c}^{\infty}$) denotes the space of continuous (infinitely differentiable) functions on $\mathbb{R}^{d}$ having compact support.
$C_{b}$ is the space of bounded continuous functions on $\mathbb{R}^{d}$ endowed with the $\sup$-norm, and $C_{b}^{k}$ is the subspace of bounded continuous functions with bounded continuous derivatives up to order $k$.
$C_{\infty}$ is a closed subspace of $C_{b}$ consisting of functions vanishing at infinity.
We denote by $\mathcal{S}$ the Schwartz space, and by $\mathcal{S}^{\prime}$ the space of tempered distributions on $\mathbb{R}^{d}$.
Let $\nabla_{i}=\partial_{x_{i}}$, $1\leq i\leq d$.
Given a matrix field $a:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$, define the row divergence operator $a\mapsto\nabla a$ via
$(\nabla a)_{j}:=\sum_{i=1}^{d}\nabla_{i}a_{ij},\quad 1\leq j\leq d.$
We denote by $H_{\xi}$ ($\xi>0$) the set of bounded symmetric uniformly elliptic Borel measurable matrix fields $a:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$:
$a=a^{\scriptscriptstyle\top},\quad\xi I\leq a(x)\leq\xi^{-1}I\text{ for a.e.\,}x\in\mathbb{R}^{d},$
for $I$ the $d\times d$ identity matrix.
Put
$\Gamma_{c}(t,x)=\Gamma(ct,x):=(4c\pi t)^{-\frac{d}{2}}e^{-\frac{|x|^{2}}{4ct}},$
the Gaussian density.
Set
$\gamma(x):=\left\{\begin{array}[]{ll}c\exp\left(\frac{1}{|x|^{2}-1}\right)&\text{ if }|x|<1,\\ 0,&\text{ if }|x|\geqslant 1,\end{array}\right$.
where $c$ is adjusted to $\int_{\mathbb{R}^{d}}\gamma(x)dx=1$, and put $\gamma_{\varepsilon}(x):=\frac{1}{\varepsilon^{d}}\gamma\left(\frac{x}{\varepsilon}\right)$, $\varepsilon>0$, $x\in\mathbb{R}^{d}$. Define the De Giorgi mollifier of a function $h\in L^{1}_{{\rm loc}}$ (or a vector field with entries in $L^{1}_{{\rm loc}}$) by
$E_{\varepsilon}h:=e^{\varepsilon\Delta}h.$
$B_{r}(x)$ denotes the open ball of radius $r$ centered at $x\in\mathbb{R}^{d}$. If $x=0$, we simply write $B_{r}$.
Given a function $f\in L^{1}_{{\rm loc}}$, we denote by $(f)_{B_{r}(x)}$ its average over the ball $B_{r}(x)$:
$(f)_{B_{r}(x)}:=\frac{1}{|B_{r}|}\int_{B_{r}(x)}fdx.$
If $x=0$, then we write $(f)_{r}\equiv(f)_{B_{r}}$.
We denote the positive and negative parts of function $f$ by
$(f)_{+}:=f\vee 0,\quad(f)_{-}:=-(f\wedge 0).$
Define weight
$\rho(y)\equiv\rho_{\epsilon_{0}}(y)=(1+\sigma|y|^{2})^{-\frac{d+\epsilon_{0}}{2}},\quad\varepsilon_{0}>0,\quad\sigma>0.$ (2.1)
This weight has property
$|\nabla\rho|\leq\frac{d+\epsilon_{0}}{2}\sqrt{\sigma}\rho.$ (2.2)
In the same way, $|\nabla\rho^{-1}|\;(=\frac{|\nabla\rho|}{\rho^{2}})\;\leq\frac{d+\epsilon_{0}}{2}\sqrt{\sigma}\rho^{-1}$. (The last two inequalities will allow us to replace all occurrences of $\nabla_{i}\rho$ or $\nabla_{i}\rho^{-1}$ resulting from the integration by parts in the analysis of PDEs in weighted spaces by the weight $\rho$ itself or $\rho^{-1}$, respectively, times a constant that is proportional to $\sqrt{\sigma}$. This constant can be made arbitrarily small by fixing $\sigma$ sufficiently small. We will use this to get rid of the terms containing $\nabla_{i}\rho$ or $\nabla_{i}\rho^{-1}$.) Put
$\rho_{x}(y):=\rho(x-y).$
3. Let $\mathbf{C}$ and $\mathbf{D}$ denote the canonical spaces of continuous and càdlàg trajectories from $0,\infty[$ to $\mathbb{R}^{d}$ equipped with the uniform topology and the Skorohod topology, respectively, endowed with the natural filtration $\mathcal{B}_{t}=\sigma\{\omega_{s}\mid 0\leq s\leq t\}$, where $\omega_{t}$ is the coordinate process.
Let $\mathcal{P}(\mathbf{C})$ and $\mathcal{P}(\mathbf{D})$ denote the space of probability measures on $\mathbf{C}$ and $\mathbf{D}$, respectively.
Recall that a probability measure $\mathbb{P}_{s,x}$ ($s\geq 0$, $x\in\mathbb{R}^{d}$) on $\mathbf{C}$ is said to be a classical martingale solution to SDE
$X_{t}=x-\int_{s}^{t}b(r,X_{r})dr+\sqrt{2}(B_{t}-B_{s}),\quad x\in\mathbb{R}^{d},\quad t\geq s,$ (2.3)
with a time-inhomogeneous drift $b\in L^{1}_{{\rm loc}}(\mathbb{R}^{1+d})^{d}$ if $\mathbb{P}_{s,x}[\omega_{s}=x]=1$,
$\mathbb{E}_{\mathbb{P}_{s,x}}\int_{s}^{t}|b(r,\omega_{r})|dr<\infty$
and for every $v\in C_{c}^{2}$
$t\mapsto v(\omega_{t})-v(x)+\int_{s}^{t}(-\Delta+b\cdot\nabla)v(\omega_{r})dr,\quad t\geq s,$
is a continuous martingale with respect to $\mathbb{P}_{s,x}$.
If $b$ does not depend on time, then we take $s=0$, and in the above martingale problem write $\mathbb{P}_{x}$.
4. BMO functions. (a) A function $g\in L^{1}_{{\rm loc}}$ is in class ${\rm BMO}={\rm BMO}(\mathbb{R}^{d})$ if
$\|g\|_{\rm{\rm BMO}}:=\sup_{x\in\mathbb{R}^{d},R>0}\frac{1}{|B_{R}|}\int_{B_{R}(x)}|g-(g)_{B_{R}(x)}|dy<\infty.$
One can also define the $\rm BMO$} semi-norm as
$\|g\|_{\rm{\rm BMO}}=\biggl(\sup_{x\in\mathbb{R}^{d},R>0}\frac{1}{|B_{R}|}\int_{B_{R}(x)}\int_{0}^{R^{2}}|\nabla e^{t\Delta}g|^{2}dtdy\biggr)^{\frac{1}{2}}$ (2.4)
(this is Carleson’s characterization of $\rm BMO$} functions).
(b) We will need the compensated compactness estimate of Coifman-Lions-Meyer-Semmes:
Proposition 2.1 ([CLMS]).
There exists a constant $C_{d}$ such that for every anti-symmetric matrix field $Q$ with entries in $\rm BMO$} one has
$|\langle Q\cdot\nabla f,\nabla g\rangle|\leq C_{d}\|Q\|_{\rm BMO}\|\nabla f\|_{2}\|\nabla g\|_{2},\quad\forall\,f,g\in W^{1,2}.$
(c) Qian-Xi [QX] employed in their work the following modification of the previous proposition. We will need it as well.
Proposition 2.2.
There exists a constant $C_{d}$ such that, for each function $h\in{\rm BMO}$, $i=1,\dots,d$,
$|\langle h,g\nabla_{i}g\rangle|\leq C_{d}\|h\|_{\rm BMO}\|\nabla g\|_{2}\|g\|_{2},\quad\forall\,g\in W^{1,2}.$
(d) The following result on the multipliers in space $\rm BMO$} on $\mathbb{R}^{d}$ follows from the analysis of Nakai-Yabuta in [NY] (see, in particular, Sect. 5 in their paper).
Lemma 2.1.
Set $\xi(y):=\frac{y_{i}}{1+\sigma|y|^{2}}$ ($\sigma>0$). Then we have
$\|\xi h\|_{\rm BMO}\leq C_{d,\sigma}\|h\|^{\prime}_{\rm BMO}\quad\text{$\forall\,h\in{\rm BMO}$},$
where $\|h\|^{\prime}_{\rm BMO}:=\langle\mathbf{1}_{B_{1}(0)}h\rangle+\|h\|_{\rm BMO}$ is the $\rm BMO$}-norm.
We will apply Lemma 2.1 in the proof of Theorem 5.1, to $\xi(y)=\frac{\nabla_{i}\rho}{\rho}=\frac{2\sigma y_{i}}{1+\sigma|y|^{2}}$, where $\rho$ is the weight introduced above.
3. Classes of singular drifts
3.1. Definitions and examples
The following classes of singular or distributional drifts are covered by our Theorems 5.1, 6.1, 7.1.
3.1.1. Form-bounded vector fields
Definition 3.1.
A vector field $b\in[L^{2}_{{\rm loc}}]^{d}$ is said to be form-bounded with form-bound $\delta>0$ (abbreviated as $b\in\mathbf{F}_{\delta}$) if
$\|b\varphi\|_{2}^{2}\leq\delta\|\nabla\varphi\|_{2}^{2}+c_{\delta}\|\varphi\|_{2}^{2}\quad\forall\,\varphi\in W^{1,2}$ (3.1)
for some constant $c_{\delta}<\infty$.
The constant $c_{\delta}$ does not affect the well-posedness of SDE (5.1) (one needs $c_{\delta}>0$ to include $L^{\infty}$ drifts). By contrast, $\delta$ is crucial: if $\delta$ exceeds the critical threshold $\delta=4$, then in general SDE (5.1) ceases to have global in time weak solution (a blow-up occurs). On the other hand, as Theorem 5.1 shows that whenever $\delta<4$, weak well-posedness holds. Moreover, as $\delta\downarrow 0$, this theory becomes more detailed, e.g. one has conditional weak uniqueness, strong well-posedness, etc. The form-bound $\delta$ appears explicitly in the examples below.
Examples 3.1.
1. If $b\in[L^{d}]^{d}+[L^{\infty}]^{d}$ (== sums of vector fields in $L^{d}]^{d}$ and in $L^{\infty}]^{d}$), then $b\in\mathbf{F}_{\delta}$ with form-bound $\delta$ that can be chosen arbitrarily small at the expense of increasing the constant $c_{\delta}$. Indeed, for every $\epsilon>0$ one can represent $b=b_{1}+b_{2}$ with $\|b_{1}\|_{d}<\epsilon$ and $\|b_{2}\|_{\infty}<\infty$. By the Sobolev embedding theorem,
$\|b\varphi\|_{2}^{2}$ $\leq 2\|b_{1}\|_{d}^{2}\|\varphi\|_{\frac{2d}{d-2}}^{2}+2\|b_{2}\|_{\infty}^{2}\|\varphi\|_{2}^{2}$
$\leq C_{S}2\|b_{1}\|_{d}^{2}\|\nabla\varphi\|_{2}^{2}+2\|b_{2}\|_{\infty}^{2}\|\varphi\|_{2}^{2},$
so $b\in\mathbf{F}_{\delta}$ with $\delta=C_{S}2\epsilon$ and $c_{\delta}=2\|b_{2}\|_{\infty}^{2}$.
2. The class of form-bounded vector fields contains the weak $L^{d}$ class:
$\|b\|_{d,\infty}:=$ $\sup_{s>0}s|\{y\in\mathbb{R}^{d}\mid|b(y)|>s\}|^{1/d}<\infty$
$\Rightarrow\quad b\in\mathbf{F}_{\delta}\quad\text{ with }\delta=\|b\|_{d,\infty}|B_{1}(0)|^{-\frac{1}{d}}\frac{2}{d-2}$
with $c_{\delta}=0$. This was proved in [KPS, Lemma 2.7].
3. The weak $L^{d}$ class includes itself the Hardy drift:
$b(x)=\pm\sqrt{\delta}\frac{d-2}{2}|x|^{-2}x\in\mathbf{F}_{\delta}\text{ with }c_{\delta}=0\quad(\text{but }b\not\in\mathbf{F}_{\delta^{\prime}}\text{ with any }\delta^{\prime}<\delta,c_{\delta^{\prime}}<\infty).$ (3.2)
In fact, inclusion (3.2) is a re-statement of the usual Hardy inequality
$\||x|^{-1}\varphi\|_{2}^{2}\leq\frac{4}{(d-2)^{2}}\|\nabla\varphi\|_{2}^{2}.$
The plus sign in (3.2) corresponds in SDE (5.1) to the attraction towards the origin, the minus corresponds to the repulsion.
4. The previous example can be refined using the weighted Hardy inequality of Hoffmann-Ostenhof–Laptev [HL]. Fix
$0\leq\Phi\in L^{s}(S^{d-1})\quad\text{ for some }s\geq\frac{2(d-2)^{2}}{2(d-1)}+1,$
where $S^{d-1}$ is the unit sphere in $\mathbb{R}^{d}$. If
$|b(x)|^{2}\leq\delta\frac{(d-2)^{2}}{4}c\frac{\Phi(x/|x|)}{|x|^{2}},\qquad\text{where $c:=\frac{|S^{d-1}|^{\frac{1}{q}}}{\|\Phi\|_{L^{s}(S^{d-1})}}$},$
then $b\in\mathbf{F}_{\delta}$ with $c_{\delta}=0$. Using this example, one can e.g. cut holes in the drift (3.2) while still controlling the value of $\delta$.
5. One can also refine example (3.2) using a Hardy-type inequality of Felli-Marchini-Terracini [FMT, Lemma 3.5]. Assume that
$|b(x)|^{2}\leq\delta\frac{(d-2)^{2}}{4}\sum_{i=1}^{\infty}\frac{\mathbf{1}_{B_{r}(a_{i})}(x)}{|x-a_{i}|^{2}},\quad x\in\mathbb{R}^{d},$
where the loci of singularities $\{a_{i}\}_{i=1}^{\infty}$ are sufficiently “spread out”:
$\sum_{i=1}^{n}|a_{i}|^{-d+2}<\infty,\quad\sum_{k=1}^{\infty}|a_{i+k}-a_{i}|^{-d+2}\text{ is bounded uniformly in $i$,}$
and $|a_{i}-a_{m}|\geq 1$ for all $i\neq m$. Then there exists $r$ sufficiently small (so, the singularities are strictly local) so that $b\in\mathbf{F}_{\delta}$ with $c_{\delta}=0$.
6. The following simple lemma applies, in particular, to the multi-particle Hardy drift $b:\mathbb{R}^{dN}\rightarrow\mathbb{R}^{dN}$ defined by (1.5) in Example 1.1.
Lemma 3.1 ([K5, Lemma 1]).
If $K\in\mathbf{F}_{\kappa}(\mathbb{R}^{d})$, then the drift $b=(b_{1},\dots,b_{N}):\mathbb{R}^{dN}\rightarrow\mathbb{R}^{dN}$ with components defined by
$b_{i}(x^{1},\dots,x^{N}):=\frac{1}{N}\sum_{j=1,j\neq i}^{N}K(x^{i}-x^{j}),\quad x^{i},x^{j}\in\mathbb{R}^{d}$
is in $\mathbf{F}_{\delta}$ with
$\delta=\frac{(N-1)^{2}}{N^{2}}\kappa,$
i.e. there is almost equality between $\delta$ and $\kappa$. (In the context of Example 1.1 $\kappa$ is the strength of attraction between the particles.)
7. Critical Morrey classes. Every vector field $b\in M_{2+\varepsilon}$ for some $\varepsilon>0$ small, i.e.
$\|b\|_{M_{2+\varepsilon}}:=\sup_{r>0,x\in\mathbb{R}^{d}}r\biggl(\frac{1}{|B_{r}|}\int_{B_{r}(x)}|b(y)|^{2+\varepsilon}dy\biggr)^{\frac{1}{2+\varepsilon}}<\infty,$
is in $\mathbf{F}_{\delta}$ with $\delta=C(d,\varepsilon)\|b\|_{M_{2+\varepsilon}}$ and $c_{\delta}=0$. The constant $C=C(d,\varepsilon)$ depends on the constants in some fundamental inequalities of harmonic analysis [Fe]. There exist far-reaching and deep extensions of this inclusion due to Adams [A1] (Appendix C) and Chiarenza-Frasca [CF].
The Morrey class $M_{2+\varepsilon}$ is substantially larger than the weak $L^{d}$ class, e.g. it includes vector fields having strong hypersurface singularities. It is easily seen that the class $M_{2+\varepsilon}$ gets larger as $\varepsilon$ gets smaller. Note, however, that by passing through the Morrey class one to a large extent loses the control over the form-bound $\delta$.
On the other hand, the form-bounded class $\mathbf{F}_{\delta}$ (with $c_{\delta}=0$) is contained in the Morrey class $M_{2}$; this is not difficult to see by selecting cutoff functions as test function $\varphi$ in the definition of $\mathbf{F}_{\delta}$. Thus, to summarize,
$\cup_{\varepsilon>0}M_{2+\varepsilon}\quad\subsetneq\quad\cup_{\delta>0}\mathbf{F}_{\delta}\;(\text{with $c_{\delta}=0$})\quad\subsetneq\quad M_{2}.$
8. A larger class than $\cup_{\varepsilon>0}M_{2+\varepsilon}$ sub-class of $\mathbf{F}_{\delta}$ was found by Chang-Wilson-Wolff [CWW], that is, $|b|\in L^{2}_{{\rm loc}}(\mathbb{R}^{d})$ and
$\sup_{r>0,x\in\mathbb{R}^{d}}\frac{1}{|B_{r}|}\int_{B_{r}(x)}|b(y)|^{2}\,r^{2}\xi\big(|b(y)|^{2}\,r^{2}\big)dy<\infty,$ (3.3)
where $\xi:\mathbb{R}_{+}\rightarrow[1,\infty[$ is a fixed increasing function such that
$\int_{1}^{\infty}\frac{ds}{s\xi(s)}<\infty.$
For instance, one can take $\xi(s)=1+(\log^{+}s)^{1+\epsilon}$ or $\xi(s)=1+\log^{+}s(\log\log^{+}s)^{1+\epsilon}$ for some $\epsilon>0$ (but not $\xi(s)=1+\log^{+}s$).
The Lebesgue, weak Lebesgue, Morrey, and Chang–Wilson–Wolff classes are all elementary sub‐classes of $\mathbf{F}_{\delta}$. We regard estimating an integral over a ball as an “elementary” calculation. This is admittedly subjective, but in practice one can carry it out quite easily for a concrete vector field. By contrast, computing an operator norm for a given vector field (or equivalently checking a quadratic‐form inequality) often requires more advanced tools, such as various forms of Hardy’s inequality.
9. There are also deep necessary and sufficient conditions for ensuring that $b\in\mathbf{F}_{\delta}$, such as the criterion of Kerman-Sawyer [KSa] or the criterion of Mazya [M2]. The latter is
$\langle\mathbf{1}_{E}|b|^{2}\rangle\leq K{\rm cap}(E)\quad\forall\,\text{compact }E\subset\mathbb{R}^{d},$ (3.4)
where, recall,
${\rm cap}(E)=\inf\{\|\nabla v\|_{2}\mid v\in C_{c}^{\infty},v\geq 1\text{ on }E\}.$
So, criterion (3.4) requires computing the capacity of an arbitrary compact set (one cannot restrict to dyadic cubes only). Kerman-Sawyer’s inequality that needs to be verified is more complex than (3.4), but the calculations are confined to dyadic cubes, so it is in some sense more practical. That said, precisely because these are necessary and sufficient conditions, their verifications can be non-trivial (for example, verifying that the Chang-Wilson-Wolff class (3.3) satisfies these conditions, see [CWW]).
Efforts to characterize inequalities of type (3.1), known as trace inequalities, remains an active research area. The interest was originally motivated by the problems related to estimating the spectrum of Schrödinger operators, cf. [Fe, MV].
Finally, we note that we can combine the previous examples:
$b_{1}\in\mathbf{F}_{\delta_{1}},\quad b_{2}\in\mathbf{F}_{\delta_{2}}\quad\Rightarrow\quad b_{1}+b_{2}\in\mathbf{F}_{\delta},\quad\delta=(\sqrt{\delta_{1}}+\sqrt{\delta_{2}})^{2}.$
This extends to series of form-bounded drifts.
3.1.2. $\text{Bounded mean oscillation}^{-1}$
Definition 3.2.
A divergence-free vector field $q=(q_{i})_{i=1}^{d}\in[\mathcal{S}^{\prime}]^{d}$ is said to be in class $\mathbf{BMO}^{-1}$ if there exist functions $Q^{ij}=-Q^{ji}\in{\rm BMO}(\mathbb{R}^{d})$ such that
$q^{i}=\sum_{j=1}^{d}\nabla_{j}Q^{ij},\quad 1\leq i\leq d,$
i.e. $q=\nabla Q$, where $\nabla$ is the row-divergence operator.
Examples of $\rm BMO$} functions include
$g\in L^{\infty}\quad\text{ or }\quad g(x)=\log|p(x)|\text{ for a polynomial $p$.}$
In the last example the $\rm BMO$} semi-norm does not depend on the coefficients of $p$.
Examples 3.2.
1. A divergence-free vector field $q$ belongs to $\mathbf{BMO}^{-1}$ if and only if the caloric extensions of its components $q^{i}$, $i=1,\dots,d$, satisfy
$\sup_{x\in\mathbb{R}^{d},R>0}\frac{1}{|B_{R}|}\int_{[0,R^{2}]\times B_{R}(x)}|e^{t\Delta}q^{i}|^{2}dtdy<\infty,$
where $e^{t\Delta}$ is the heat semigroup, see [KT].
Equivalently, $q^{i}$ can be characterized as elements of homogeneous Triebel-Lizorkin space $\dot{F}^{-1}_{2,\infty}$.
2. The divergence-free vector fields with entries in Morrey space $M_{1}$, i.e. such that
$\langle|q|\mathbf{1}_{B_{r}(x)}\rangle\leq Cr^{d-1}$
with constant $C$ independent of $r$ or $x\in\mathbb{R}^{d}$, are in $\mathbf{BMO}^{-1}$ [M, Sect. 3.4.5].
3. The divergence-free vector fields with entries in Besov space $B^{-1+d/p}_{p,\infty}$, $p>d$, are in $\mathbf{BMO}^{-1}$. This follows by recalling that $B^{-1+d/p}_{p,\infty}$ consists of tempered distribution $h$ such that
$\|e^{t\Delta}h\|_{p}\leq Ct^{-\frac{1-\frac{d}{p}}{2}},\quad 0<t\leq 1,$
see [KT] for details.
In [KT], Koch and Tataru established, among other results, the existence and the uniqueness of global in time mild solution to Cauchy problem for the 3D Navier-Stokes equations in the critical space $Z$ of functions $v:\mathbb{R}_{+}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{3}$ satisfying
$\|v\|_{Z}:=\sup_{t\geq 0}t^{\frac{1}{2}}\|v(t)\|_{\infty}+\biggl(\sup_{x\in\mathbb{R}^{d},R>0}\frac{1}{|B_{R}|}\int_{[0,R^{2}]\times B_{R}(x)}|v(t,y)|^{2}dtdy\biggr)^{\frac{1}{2}}<\infty,$ (3.5)
provided that the initial data $v(0)\in\mathbf{BMO}^{-1}(\mathbb{R}^{3})$ have sufficiently small norm. (An even larger $\mathbf{BMO}^{-1}$ type space is considered in the recent paper [CE].)
3.1.3. Multiplicatively form-bounded vector fields
This class of drifts was mentioned in the introduction (class $\mathbf{MF}_{\delta}$). We postpone its discussion until Section 6.
3.2. “Physical” approximations
We now introduce the classes of bounded smooth approximations of vector fields in $\mathbf{F}_{\delta}$ and ${\mathbf{BMO}}^{-1}$ that preserve the structure constants of the latter.
Definition 3.3.
Given $b\in\mathbf{F}_{\delta}$, we denote by $[b]$ the set of sequences of vector fields $\{b_{n}\}\subset[C_{b}\cap C^{\infty}]^{d}$ such that
[…]
(note that we can always increase $c_{\delta}$, if needed; what matters is that $c_{\delta}$ does not depend on $n$).
For instance, the sequence $\{b_{n}\}$ defined by $b_{n}:=E_{\varepsilon_{n}}b$ (De Giorgi mollifier $E_{\varepsilon}$ is defined in Section 2) for any $\varepsilon_{n}\downarrow 0$ is in $[b]$. See [KS4] or [K5, Sect. 6] for the proof.
In the quantum-mechanical context, the form-boundedness condition on potential $|b|^{2}$ expresses smallness of the potential energy with respect to the kinetic energy in the system described by the Hamiltonian $\Delta-|b|^{2}$. Our conditions on regularizations $\{b_{n}\}$ is thus that they, essentially, do not increase the potential energy.
Given a potential $0\leq V\in L^{1}_{{\rm loc}}$, we write, with some abuse of notation, $V^{\frac{1}{2}}\in\mathbf{F}_{\delta_{+}}$ if
$\langle V,\varphi^{2}\rangle\leq\delta_{+}\langle|\nabla\varphi|^{2}\rangle+c_{\delta_{+}}\langle\varphi^{2}\rangle\quad\forall\,\varphi\in C_{c}^{\infty}.$
Definition 3.4.
Let
$b\in\mathbf{F}_{\delta},\quad({\rm div\,}b)^{\frac{1}{2}}_{+}\in\mathbf{F}_{\delta_{+}},\quad({\rm div\,}b)_{-}\in L^{1}+L^{\infty},$
we denote by $b]^{\prime}$ the set of sequences of vector fields $\{b_{n}\}\subset[C^{1}_{b}\cap C^{\infty}]^{d}$ that satisfy the previous inclusions with the same constants $\delta$, $c_{\delta}$, $\delta_{+}$, $c_{\delta_{+}}$ (so, independent of $n$) and such that
$b_{n}\rightarrow b\text{ in }[L^{2}_{{\rm loc}}]^{d},\quad{\rm div\,}b_{n}\rightarrow{\rm div\,}b\text{ in }L^{1}_{{\rm loc}}.$
Once again, we can take e.g. $b_{n}:=E_{\varepsilon_{n}}b$, see [K5, Sect. 6] for the proof.
Definition 3.5.
Given a matrix field $a\in H_{\xi}$ and a vector field $b$ such that $\nabla a+b\in\mathbf{F}_{\delta}$, we denote by $[a,b]$ the set of sequences $\{a_{n}\}\in H_{\xi}\cap[C_{b}\cap C^{\infty}]^{d\times d}$, $\{b_{n}\}\in[C_{b}\cap C^{\infty}]^{d}$ such that
$\nabla a_{n}+b_{n}\in\mathbf{F}_{\delta}\text{ with $c_{\delta}$ independent of $n$},$
and
$\nabla a_{n}+b_{n}\rightarrow\nabla a+b\text{ in }[L^{2}_{{\rm loc}}]^{d},\quad a_{n}\rightarrow a\quad\text{ a.e.\,on }\mathbb{R}^{d}$
as $n\rightarrow\infty$.
For example, we can take $a_{n}=E_{\varepsilon_{n}}a$, see [KS3, Sect. 4.4] for the proof.
Definition 3.6.
Given $q=\nabla Q\in\mathbf{BMO}^{-1}$, where $Q$ is the corresponding anti-symmetric matrix field with entries in $\rm BMO$} (so, ${\rm div\,}q=0$), we denote by $[q]$ the set of sequences
$\{q_{m}=\nabla Q_{m}\text{ for anti-symmetric }Q_{m}\in[C^{\infty}\cap W^{1,\infty}]^{d\times d}\}$
such that
[…]
For example, one possible choice is
$Q_{m}:=E_{\varepsilon_{m}}(Q\wedge U_{\varepsilon_{m}}\vee V_{\varepsilon_{m}}),$
where the maximum and the minimum are taken componentwise,
$U_{\varepsilon}:=(-c\log|x|+\varepsilon^{-1})\wedge\varepsilon^{-1}\vee 0,\quad V_{\varepsilon}:=(c\log|x|-\varepsilon^{-1})\wedge 0\vee(-\varepsilon^{-1}),\quad\varepsilon_{m}\downarrow 0$
with $c$ chosen so that $\|c\log|x|\|_{{\rm BMO}}\leq\|Q\|_{{\rm BMO}}$. The last two functions are compactly supported and are in $\rm BMO$}. Since $\rm BMO$} is a lattice, the components of the matrix fields $Q_{m}$ (still anti-symmetric) are in $\rm BMO$}. This regularization of $Q$ was employed earlier in [QX].
4. Preliminary discussion: blow-up thresholds for Brownian particles
We continue discussing the particle system in Example 1.1, i.e.
$X^{i}_{t}=x^{i}_{0}-\sqrt{\kappa}\frac{d-2}{2}\frac{1}{N}\sum_{j=1,j\neq i}^{N}\int_{0}^{t}\frac{X_{s}^{i}-X_{s}^{j}}{|X_{s}^{i}-X_{s}^{j}|^{2}}ds+\int_{0}^{t}q_{0}(X_{s}^{i})ds+\sqrt{2}B^{i}_{t},\quad i=1,\dots,N,$ (4.1)
where $q_{0}\in\mathbf{BMO}^{-1}(\mathbb{R}^{d})$, $d\geq 3$.
1. First, we present positive well-posedness results for (4.1) that follow from Theorems 5.1 and 6.1.
2. Next, we exhibit counterexamples to well-posedness of particle system (4.1) when the attraction strength $\kappa$ is too large. On one hand, these counterexamples test the sharpness of Theorems 5.1 and 6.1; on the other, they play a crucial role in the proof of Theorem 6.2, which provides an improved upper bound on the constant in the many-particle Hardy inequality.
3. Finally, we discuss the two- and the one-dimensional cases.
4.1. Positive results
Theorem 6.1 covers a large portion of the admissible range for the attraction parameter $\kappa$ in (4.1):
$\kappa\in\bigg[0,16\biggl(1\vee\frac{N}{1+\sqrt{1+\frac{3(d-2)^{2}}{(d-1)^{2}}(N-1)(N-2)}}\biggr)^{2}\bigg[.$ (4.2)
Here, the right endpoint arises from Hoffmann-Ostenhof, Hoffmann-Ostenhof, Laptev and Tidblom’s lower bound on the constant in the many-particle Hardy inequality [HHLT]. Since the optimal constant in that inequality is not yet known, one expects the true admissible interval for $\kappa$ to be strictly larger. Note that in dimensions $d\geq 7$ the interval (4.2) reduces to $\kappa\in[0,16[$.
By contrast, Theorem 5.1 handles only $\kappa\in[0,4\frac{N^{2}}{(N-1)^{2}}[$, but it offers greater flexibility in modifying the interaction kernel in (4.1). For example, one can multiply the attracting interaction kernel by any function of $L^{\infty}$ norm at most one, without altering the assumption on $\kappa$. In particular, one can “cut holes” in the interaction kernel so that the particles do not interact along certain directions (cf. Examples 3.1.4).
One can also describe how the particles behave as they approach collision via estimates on the heat kernel of the corresponding Kolmogorov operator that, necessarily, involve a desingularizing weight, see [BK].
4.2. Counterexamples
These were the positive results for (4.1) that we now balance with counterexamples, i.e. analogues of (a), (b) in Section 1.1.1. For simplicity, assume there is no divergence-free distributional component of the drift (i.e. $q_{0}=0$). The right endpoint of the interval (4.2) lies just below the first blow-up threshold for (4.1), or the “non-sticky collisions threshold”. More precisely, following [F], set
$R_{t}:=\frac{1}{4N}\sum_{i,j=1}^{N}|X_{t}^{i}-X_{t}^{j}|^{2}.$
It is not difficult to see that $R_{t}$ is a local squared Bessel process, i.e.
$R_{t}=R_{0}+2\int_{0}^{t}\sqrt{|R_{t}|}dW_{t}+\mu t,$ (4.3)
where $W_{t}$ is a one-dimensional Brownian motion, and
$\mu=(N-1)\left(d-\sqrt{\kappa}\frac{d-2}{4}\right)$
is its “dimension”; the value of $\mu$ controls how often $R_{t}$ hits zero. In what follows, by a collision of particles in (4.1) we mean a collision of all $N$ particles at the same time, i.e. when $R_{t}=0$. So, by standard theory of Bessel processes (see  [RY, Ch. XI, §1]):
(a’) If $\kappa\geq 16(\frac{d}{d-2})^{2}$, then there are a.s. sticky collisions in (4.1), i.e. the particles collide in finite time and stay clumped up. Furthermore, one can show that for $\kappa>16(\frac{d}{d-2})^{2}$ the particle system ceases to have a weak solution, cf. (a) in Section 1.1.1.
(b’) If $16\big(\frac{d}{d-2}\big)^{2}\bigl(1-\frac{2}{d(N-1)}\bigr)<\kappa<16(\frac{d}{d-2})^{2}$, then there are a.s. non-sticky collisions in (4.1), i.e. particles collide infinitely many times, but $\int_{0}^{\infty}R_{t}dt<\infty$ a.s.
Comparing (a’), (b’) with the admissible range (4.2) in Theorem 6.1, we see that the latter provides a result that is close to optimal.
4.3. Two- and one-dimensional cases
For the two-dimensional counterpart of (4.1), namely, the finite-particle approximation of the Keller-Segel model,
$X^{i}_{t}=x^{i}-\sqrt{\varkappa}\int_{0}^{t}\frac{X_{s}^{i}-X_{s}^{j}}{|X_{s}^{i}-X_{s}^{j}|^{2}}ds+\sqrt{2}B^{i}_{t},\quad x=(x^{1},\dots,x^{N})\in\mathbb{R}^{2N},$
the attracting interaction kernel is no longer locally square intergrable and therefore is not form-bounded, so one cannot apply our Theorems 5.1, 6.1 (see, however, Appendix A where we to some extent address $d=2$). Cattiaux-Pédèches [CP] and Fournier-Jourdain [FJ], Fournier-Tardy [FT] exploited the special structure of the drift and constructed process $(X_{t}^{1},\dots,X_{t}^{N})$. They work either via a suitable Dirichlet form or by exhibiting a weak solution to the SDE, respectively. In particular, they cover the full critical range $\varkappa\in[0,16[$, where $16$ is the sticky collisons threshold in dimension two. In this context, note:
1) The arguments of [FJ, FT] work in dimensions $d\geq 3$ as well, and allow to handle $16\leq\kappa<16(\frac{d}{d-2})^{2}$ in (4.1), which includes non-sticky collisions, see (b’) above.
2) Ohashi-Russo-Texeira [ORT] consider squared Bessel processes in the low-dimensional regime $0<\nu<1$ (i.e. non-sticky collisions). They characterize the process as the unique solution of SDE
$dX_{t}=\frac{1-\nu}{2}\frac{dt}{X_{t}}+dB_{t},\quad X_{0}=x_{0}>0,\quad 0<\nu<1,$
whose drift $b(x)=\frac{1}{x}$ is not locally in $L^{1}_{{\rm loc}}$, and thus has to be considered as a distributional drift.
It is not yet clear whether Theorems 5.1, 6.1 can be extended in some form to the non-sticky collisions part of the interval of admissible values of $\kappa$. However, taking into account the positive results in 1) and in 2), such extension is conceivable.
We refer to Cattiaux [C] and Fournier [F] for recent surveys (among new results) on the Keller-Segel model and its finite particle approximations.
5. General drifts
In this section we consider SDE
$X_{t}=x-\int_{0}^{t}\bigl(b(X_{s})+q(X_{s})\bigr)ds+\sqrt{2}B_{t},\quad x\in\mathbb{R}^{d},\quad t\geq 0.$ (5.1)
with
$b\in\mathbf{F}_{\delta},\quad q\in\mathbf{BMO}^{-1}\;\;(\Rightarrow{\rm div\,}q=0),$
with the form-bound $\delta$ of $b$ going all the way up to (but staying strictly less) the critical threshold $\delta=4$. As mentioned in the introduction, by a result of Mazya and Verbitsky, this assumption on $b$ and $q$ is equivalent to having generalized form-boundedness (1.6) of $c=b+q$.
We fix bounded smooth approximations $\{b_{n}\}\in[b]$, $\{q_{m}\}\in[q]$ (as in Section 3.2). Consider the approximating SDEs
$X^{n,m}_{t}=x-\int_{0}^{t}\big(b_{n}(X^{n,m}_{s})+q_{m}(X^{n,m}_{s})\big)ds+\sqrt{2}B_{t},$ (5.2)
on a complete probability space $\mathfrak{F}=(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbf{P})$, with $B_{t}$ being a $\mathcal{F}_{t}$-Brownian motion. By classical theory, for every $x\in\mathbb{R}^{d}$ and every $n=1,2,\dots,$ there exists a pathwise unique strong solution $\{X_{t}^{n,m}\}_{t\geq 0}$ to (5.2). The corresponding Kolmogorov operators
$\Lambda(b_{n},q_{m}):=-\Delta+(b_{n}+q_{m})\cdot\nabla,\quad D\big(\Lambda(b_{n},q_{m})\big)=(1-\Delta)^{-1}C_{\infty}.$
generate strongly continuous Feller semigroups on $C_{\infty}$ such that
$e^{-t\Lambda(b_{n},q_{m})}f(x)=\mathbf{E}[f(X^{n,m}_{t})].$
Set $\mathbb{P}_{x}^{n,m}:=\mathbf{P}(X_{t}^{n,m})^{-1}$.
Theorem 5.1.
Let $d\geq 3$. Assume that $b$ and $q$ are, respectively, Borel measurable and distribution-valued vector fields $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that satisfy
$\left\{\begin{array}[]{l}b\in\mathbf{F}_{\delta}\text{ with }\delta<4,\\ q\in\mathbf{BMO}^{-1}.\end{array}\right$. (5.3)
Let $\{b_{n}\}\in[b]$, $\{q_{m}\}\in[q]$ as in Definitions 3.3 and 3.6. The following are true:
  1. (i)
    (Feller semigroup) The limit
    $s\mbox{-}C_{\infty}\mbox{-}\lim_{n}\lim_{m}e^{-t\Lambda(b_{n},q_{m})}\text{ (loc.\,uniformly in $t\geq 0$)}$
    exists and determines a strongly continuous Feller semigroup, say, $e^{-t\Lambda}=e^{-t\Lambda(b,q)}$. (The order in which we take the limits is essential.) The generator $\Lambda$ of $e^{-t\Lambda}$ is an operator realization of the formal differential expression $\Delta+(b+q)\cdot\nabla$ in $C_{\infty}$.
  2. (ii)
    (A relaxed approximation uniqueness) The limit in (i) does not depend on the choice of $\{b_{n}\}\in[b]$ and $\{q_{m}\}\in[q]$. Furtermore, if $\delta<1$ and $\{b_{n}\}\in[b]\cap[L^{2}]^{d}$, then already weak convergence
    $b_{n}\overset{w}{\rightarrow}b\quad\text{in $[L^{2}]^{d}$},$
    yields convergence of the approximating Feller semigroups to the same limit from (i):
    $e^{-t\Lambda(b,q)}=s\mbox{-}C_{\infty}\mbox{-}\lim_{n}\lim_{m}e^{-t\Lambda(b_{n},q_{m})}\quad\text{(loc.\,uniformly in $t\geq 0$)}.$
    Thus, when $\delta<1$, we can extend quite substantially the class of admissible approximations of $b$.
  3. (iii)
    (Generalized martingale solution) There exists a strong Markov family of probability measures $\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$ on the canonical space $\mathbf{C}$ of continuous trajectories such that
    $e^{-t\Lambda(b,q)}f(x)=\mathbb{E}_{\mathbb{P}_{x}}[f(\omega_{t})],\quad f\in C_{\infty},\quad x\in\mathbb{R}^{d},\;t\geq 0,$
    $\mathbb{P}_{x}=w\mbox{-}\mathcal{P}(\mathbf{C})\mbox{-}\lim_{n}\lim_{m}\mathbb{P}_{x}^{n,m},$
    and for every test function $v$ in the domain $D\big(\Lambda(b,q)\big)$, a dense subspace of $C_{\infty}$, the process
    $t\mapsto v(\omega_{t})-v(x)+\int_{0}^{t}\Lambda(b,q)v(\omega_{s})ds$
    is a continuous martingale under $\mathbb{P}_{x}$. Selecting test functions $v$ from the domain of the Feller generator allows to address the problem of defining the term $\int_{0}^{t}q(X_{s})\cdot\nabla v(X_{s})ds$ in the martingale problem.
  4. (iv)
    (Weak solution for disperse initial data) Let $\delta<1$ and let us also assume that $b$, $q$, $b_{n}$, $q_{m}$ have supports in a ball of fixed radius. Given an initial (smooth) probability density $\nu_{0}$ satisfying $\langle\nu_{0}^{2r}\rangle<\infty$ for some $1<r<\frac{1}{\sqrt{\delta}}$, there exist a probability space $\mathfrak{F}^{\prime}=(\Omega^{\prime},\mathcal{F}^{\prime},\{\mathcal{F}^{\prime}_{t}\}_{t\geq 0},\mathbf{P}^{\prime})$ and a continuous process $X_{t}$ on this space such that the limit
    $A_{t}:=L^{2}(\Omega^{\prime})\mbox{-}\lim_{n}\lim_{m}\int_{0}^{t}\big(b_{n}(X_{s})+q_{m}(X_{s})\big)ds$
    exists, and we have a.s.
    $X_{t}=X_{0}-A_{t}+\sqrt{2}B_{t},\quad t>0,$
    for a $\mathcal{F}_{t}^{\prime}$-Brownian motion $B_{t}$, for $\mathbf{P}^{\prime}X_{0}^{-1}$ having density $\nu_{0}$. (The compact support assumption can be removed with a few additional efforts at expense of requiring $\langle\nu_{0}^{2r}\rho^{-\alpha}\rangle<\infty$, where the weight $\rho$ is defined by (2.1), for appropriate $\alpha>0$, see Remark 8.4.)
  5. (v)
    (Classical weak solution) If $q=0$, then, for every $x\in\mathbb{R}^{d}$,
    $\mathbb{E}_{\mathbb{P}_{x}}\int_{0}^{1}|b(\omega_{s})|ds<\infty$
    and, for every test function $v\in C_{c}^{2}$, the process
    $t\mapsto v(\omega_{t})-v(x)+\int_{0}^{t}(-\Delta+b\cdot\nabla)v(\omega_{s})ds$
    is a continuous martingale with respect to $\mathbb{P}_{x}$. Moreover,
    $B_{t}(\omega):=\frac{1}{\sqrt{2}}\bigg(\omega_{t}-x+\int_{0}^{t}b(\omega_{s})ds\bigg),\quad t\geq 0,$
    is a Brownian motion, so we have a weak solution to SDE (5.1).
    If, in addition to $q=0$, we have $\delta<\frac{C}{d^{2}}$ for sufficiently small constant $C$, then more can be said:
  6. (vi)
    (Strong solvability) The strong solutions $X_{t}^{n}$ of the approximating SDEs
    $X^{n}_{t}=x-\int_{0}^{t}b_{n}(X^{n}_{s})ds+\sqrt{2}B_{t},$
    considered on a fixed complete probability space, converge a.s. after passing to a subsequence independent of $x$, to a strong solution $X_{t}$ to the SDE
    $X_{t}=x-\int_{0}^{t}b(X_{s})ds+\sqrt{2}B_{t}.$ (5.4)
    That is, $X=\Phi(B)$ for some Borel measurable function $\Phi$.
  7. (vii)
    (Another kind of approximation uniqueness) Let $\delta<\frac{4}{(d-2)^{2}}\wedge 1$. Let $\{\mathbb{Q}_{x}\}_{x\in\mathbb{R}^{d}}$ be a family of solutions to the martingale problem in (v) that are constructed via approximation, i.e. are such that
    $\mathbb{Q}_{x}=w{\mbox{-}}\mathcal{P}(\mathbf{C})\mbox{-}\lim_{n}\mathbb{P}_{x}(\tilde{b}_{n})\quad\text{for every $x\in\mathbb{R}^{d}$},$
    where $\tilde{b}_{n}\in\mathbf{F}_{\delta}\cap[C_{b}\cap C^{\infty}]^{d}$ with $c_{\delta}$ independent of $n$. Then
    $\{\mathbb{Q}_{x}\}_{x\in\mathbb{R}^{d}}=\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}},$
    where $\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$ were constructed in (iii). Here we do not require any convergence of $\tilde{b}_{n}$ to $b$.
  8. (viii)
    (Stochastic transport equation) Let $b\in\mathbf{F}_{\delta}$ with $\delta<(1+4rd)^{-2}$ for a given $r=1,2,\dots$ Then, for every $v_{0}\in W^{1,4r}$, there exists a unique weak solution to Cauchy problem for the stochastic transport equation
    $dv+b\cdot\nabla vdt+\sqrt{2}\nabla v\circ dB_{t}=0,\quad v|_{t=0}=f,$ (5.5)
    with $\circ$ denoting the Stratonovich multiplication. It satisfies
    $\sup_{0\leq\alpha\leq 1}\bigl\|\mathbb{E}|\nabla v|^{2r}\bigr\|_{L^{\frac{2}{1-\alpha}}([0,t],L^{\frac{2d}{d-2+2\alpha}})}\leq C_{1}e^{C_{2}t}\|\nabla f\|^{2r}_{4r}.$
    In particular, if $2r>d$, then by the Sobolev embedding theorem for a.e. $\omega\in\Omega$ the function $x\mapsto v(t,x,\omega)$ is Hölder continuous, possibly after modification on a set of measure zero in $\mathbb{R}^{d}$ (in general, depending on $\omega$).
The novelty of Theorem 5.1 is in assertions (i)-(iv). Assertions (v)-(viii) were proved in [KS2], [KM1], [KS7], [KSS], respectively. (The strong solvability (vi) is proved in [KM1] via a modification of the method of Röckner-Zhao, under the additional compact support hypothesis on $b$, but it can be removed with a few additional efforts.)
Since the uniqueness results play an important role in the theory of SDE (5.1), we included the proof of assertion (vii), i.e. another kind of approximation uniquness.
In the case when $q=0$ and the form-bound $\delta$ of $b$ is sufficiently small, one can furhermore prove uniqueness of the constructed weak or strong solutions in the classes of solutions satisfying Krylov-type bounds. For details, see [KM2], [K4] (conditional weak uniqueness) and [KM1] (conditional strong uniqueness).
In the case distributional $q\neq 0$, we do not establish conditional uniqueness or strong solvability. Nevertheless, already the existence of the Feller semigroup and its approximation uniqueness cover the following rather common scenario arising in the study of physical models: one mollifies a singular drift and studies the corresponding SDE as an approximation of the true dynamics (for instance, when investigating long‐term behavior such as sub‐ or super‐diffusivity, see e.g. [ABK]). Although the mollified drift is bounded and smooth, and so the corresponding SDE is well-posed, one still must ensure that:
– the resulting stochastic dynamics does not depend on the choice of the regularization of the drift (in particular, on the choice of the mollifier);
– in the limit one still gets a non-pathological diffusion process.
The former is addressed by assertion (ii). The latter is addressed by assertions (i), (iii), (iv).
Example 5.1.
Let us return to the problem of describing the dynamics $X_{t}=(X_{t}^{1},\dots,X_{t}^{N})$ of $N$ interacting particles immersed in a velocity field in $\mathbf{BMO}^{-1}$, $d\geq 3$. That is, we are in the setting of Example 1.1, where, recall,
$X_{t}=x_{0}-\int_{0}^{t}\bigl(b(X_{s})+q(X_{s})\bigr)ds+\sqrt{2}B_{t},\quad x_{0}=(x_{0}^{1},\dots,x_{0}^{N})\in\mathbb{R}^{dN},\quad B_{t}=(B_{t}^{1},\dots,B_{t}^{N}),$ (5.6)
and $b(x)=(b^{1}(x),\dots,b^{N}(x))$, $q(x)=(q_{0}(x^{1}),\dots,q_{0}(x^{N}))$ ($x=(x^{1},\dots,x^{N})\in\mathbb{R}^{dN}$),
$b^{i}(x^{1},\dots,x^{N}):=\frac{1}{N}\sum_{j=1,j\neq i}^{N}\sqrt{\kappa}\frac{d-2}{2}\frac{x^{i}-x^{j}}{|x^{i}-x^{j}|^{2}},\quad q_{0}\in\mathbf{BMO}^{-1}(\mathbb{R}^{d}).$
Then, by Lemma 3.1, $b\in\mathbf{F}_{\delta}(\mathbb{R}^{dN})$ with $\delta=\frac{(N-1)^{2}}{N^{2}}\kappa$. Also, as mentioned in the introduction, $q\in\mathbf{BMO}^{-1}(\mathbb{R}^{dN})$. Therefore, if the strength of attraction between the particles $\kappa$ satisfies
$\kappa<4\frac{N^{2}}{(N-1)^{2}}$ ($\kappa_{\rm hyp}$)
(so that $\delta<4$), then Theorem 5.1 applies and ensures the existence and the approximation uniqueness for this particle system. Importantly, the assumption on $\kappa$ basically does not depend on the number of particles $N$ (assumed to be large).
In the case $d=2$, which is of interest e.g. in the Keller-Segel model, the drift $b:\mathbb{R}^{2N}\rightarrow\mathbb{R}^{2N}$ defined above is not form-bounded. However, it is weakly form-bounded, which still allows us to say something about the corresponding SDE (5.6), see Appendix A.
Remark 5.1 (On the critical threshold $\delta=4$).
1. In the construction of the Feller semigroup in Theorem 5.1, we run De Giorgi’s method in $L^{p}$ for
$p>\frac{2}{2-\sqrt{\delta}}$ (5.7)
(and so we need $\delta<4$; the counterexamples discussed in the introduction show that $\delta<4$ is sharp at least in high dimensions). The condition (5.7) comes from the following elementary calculation for the Kolmogorov backward equation. Let $b\in\mathbf{F}_{\delta}$ and $q\in\mathbf{BMO}^{-1}$ be additionally bounded and smooth so that the manipulations with the equations are justified, but the constants in the estimated will not depend on the smoothness of $b$ and $q$. Let us also assume for simplicity that $c_{\delta}=0$ (if not, then we need to add a constant term in the Kolmogorov equation to absorb $c_{\delta}>0$). Consider Cauchy problem
$(\partial_{t}-\Delta+(b+q)\cdot\nabla)v=0,\quad v|_{t=0}=v_{0}\in C_{c}^{\infty}.$
Without loss of generality, $v_{0}\geq 0$, and so $v\geq 0$. Multiply equation by $v^{p-1}$ and integrate by parts. Since ${\rm div\,}q=0$, one finds
$\frac{1}{p}\langle\partial_{t}v^{p}\rangle+\frac{4(p-1)}{p^{2}}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle+\frac{2}{p}\langle b\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\rangle=0,$
or, equivalently,
$\partial_{t}\langle v^{p}\rangle+\frac{4(p-1)}{p}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle=-2\langle b\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\rangle.$
Applying the Cauchy-Schwarz inequality in the last term gives
$\langle\partial_{t}v^{p}\rangle+\frac{4(p-1)}{p}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle\leq 2\biggl(\alpha\langle|b|^{2},v^{p}\rangle+\frac{1}{4\alpha}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle\biggr).$
Now, applying $b\in\mathbf{F}_{\delta}$ and selecting $\alpha=\frac{1}{2\sqrt{\delta}}$, we obtain
$\langle\partial_{t}v^{p}\rangle+\biggl[\frac{4(p-1)}{p}-2\sqrt{\delta}\biggr]\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle\leq 0.$ (5.8)
To keep the dispersion term positive, one needs $\frac{4(p-1)}{p}-2\sqrt{\delta}>0$, i.e. $p>\frac{2}{2-\sqrt{\delta}}$. This calculation reappears (in slightly different form, e.g. for sub-solutions of the Kolmogorov equation) in the proof of Theorem 5.1(i).
The observation that one should work in $L^{p}$, $p>\frac{2}{2-\sqrt{\delta}}$, was made already in [KoS] in the context of the $L^{p}$ semigroup theory of $\Delta+b\cdot\nabla$, $b\in\mathbf{F}_{\delta}$. See also [CPZ, JL] where the authors use energy methods in $L^{p}$ to study regularity of solutions of the elliptic-parabolic Keller-Segel model of chemotaxis (i.e. of the corresponding McKean-Vlasov PDE), although the optimal choice of $p$ is not really discussed in these papers.
2. It is interesting to abstract away the previous calculation to see if there are other test functions that allow for a similar energy analysis of Cauchy problem $(\partial_{t}-\Delta+b\cdot\nabla)v=0$, $v|_{t=0}=v_{0}$, where $0\leq v_{0}\in C_{c}^{\infty}$. We seek test functions $\varphi(v)=\varphi(v(t))$ such that $\varphi\geq 0$, $\varphi^{\prime}\geq 0$, and so that there exists a “conjugate” function $\psi$ satisfying $\psi(0)=0$ and
$\psi^{\prime}=\sqrt{\varphi^{\prime}},\quad\varphi=a\psi\psi^{\prime}$ (5.9)
for some constant $a>0$ to be chosen (hence $\psi\geq 0$). Multiplying the parabolic equation by $\varphi(v)$ and integrating by parts, we obtain
$\langle\partial_{t}v,\varphi(v)\rangle$ $=a\langle\partial_{t}v,\psi(v)\psi^{\prime}(v)\rangle=\frac{a}{2}\langle\partial_{t}(\psi(v))^{2}\rangle,$
$\langle-\Delta v,\varphi(v)\rangle$ $=\langle\nabla v,\varphi^{\prime}(v)\nabla v\rangle=\langle|\nabla\psi(v)|^{2}\rangle,$
and
$\langle b\cdot\nabla v,\varphi(v)\rangle$ $=\langle b\cdot\nabla\psi(v),a\psi(v)\rangle$
$\leq a\biggl(\alpha\langle|b|^{2},(\psi(v))^{2}\rangle+\frac{1}{4\alpha}\langle|\nabla\psi(v)|^{2}\rangle\biggr)$
(take $\alpha=1/2\sqrt{\delta}$ and apply $b\in\mathbf{F}_{\delta}$ (with $c_{\delta}=0$))
$\leq a\sqrt{\delta}\langle|\nabla\psi(v)|^{2}\rangle.$
Thus, we obtain an energy inequality of the form
$\frac{a}{2}\partial_{t}\langle(\psi(v))^{2}\rangle+\bigl(1-a\sqrt{\delta}\bigr)\langle|\nabla\psi(v)|^{2}\rangle\leq 0.$
In order for the disperson term to remain non-negative, we need to take in (5.9) $a=\frac{1}{\sqrt{\delta}}$ (or smaller, but the equality is least restrictive on $\varphi$). So far, no constraint on $\delta$ has appeared, but it will appear once we solve the resulting (from (5.9)) system
$\left\{\begin{array}[]{l}\psi^{\prime}=\sqrt{\varphi^{\prime}},\\ \varphi=\frac{\psi\psi^{\prime}}{\sqrt{\delta}}\end{array}\right.\qquad\Rightarrow\qquad\varphi^{\prime}=\frac{1}{\sqrt{\delta}}(\psi\psi^{\prime\prime}+(\psi^{\prime})^{2}),$
so that $\psi$ must satisfy
$\psi\psi^{\prime\prime}=(\sqrt{\delta}-1)(\psi^{\prime})^{2}.$
Assuming $\psi>0$ for $v>0$, we can reduce order to obtain
$\psi^{\prime}=C\psi^{\sqrt{\delta}-1}.$ (5.10)
We look for non-trivial solutions $\psi$ that are defined globally. Hence the right-hand side of (5.10) must grow at most linearly, which forces
$\sqrt{\delta}-1\leq 1\quad\Rightarrow\quad\delta\leq 4.$
If $1<\sqrt{\delta}-1<1$, i.e. $0<\delta<4$, then we find $\psi(v)=C_{1}v^{\frac{1}{2-\sqrt{\delta}}}$, among other possible solutions of this type (note that there is no uniqueness here). Then, using equation $\psi^{\prime}=\sqrt{\varphi^{\prime}}$, we find
$\varphi(v)=cv^{\frac{2}{2-\sqrt{\delta}}-1}=cv^{p-1},\quad p:=\frac{2}{2-\sqrt{\delta}},$
covering the classical power-type test function discussed in the beginning of this remark.
If $\delta=4$, then ODE (5.10) becomes linear. Taking, for instance, $C=\frac{1}{2}$ in (5.10) yields the non-trivial solution
$\psi(v)=2e^{\frac{v}{2}},\quad\varphi(v)=e^{v}.$
This is essentially the test function that allows us to treat the critical threshold $\delta=4$, see below. Although $\psi(0)\neq 0$, which fails the original requirement $\psi(0)=0$ needed above to apply the form-boundedness of $b$, one can overcome this by working with test functions $\varphi(v)=e^{v}-1$ or $\varphi(v)=e^{v}-e^{-v}$ multiplied by a cutoff function, or working on torus instead of $\mathbb{R}^{d}$ [K2, KS6].
3. Let $b\in\mathbf{F}_{4}$. Assume additionally that $b$ and the approximating vector fields $\{b_{n}\}\in[b]$ have supports in a fixed ball (or, more generally, decay sufficiently rapidly at infinity uniformly in $n$); since here we are interested in admissible local singularities of $b$, this is not a particularly restrictive assumption. Let $q\in\mathbf{BMO}^{-1}$. By following closely the proof of [K2, Theorem 1] (see also [KS6, Theorem 2]) i.e. using test function $e^{v}-e^{-v}$ in the analysis of Cauchy problem for the Kolmogorov PDE $(\partial_{t}-\Delta+b\cdot\nabla)v=0$, one can show that the limit
$s\mbox{-}L_{{\rm\cosh-1}}\mbox{-}\lim_{n}\lim_{m}e^{-t\Lambda(b_{n},q_{m})}\text{ (loc.\,uniformly in $t\geq 0$)}$
exists and determines a strongly continuous Markov semigroup $e^{-t\Lambda(b)}$ on the Orlicz space
[…]
$\|f\|_{\cosh-1}=\inf\left\{c>0\mid\langle\cosh\frac{f}{c}-1\rangle\leq 1\right\}.$
On the torus, in the case $q=0$ (i.e. no distributional component of the drift), [K2] established the following energy inequality for $v(t)=e^{-t\Lambda(b_{n})}v_{0}$:
$\frac{1}{2}\sup_{s\in[0,t]}\langle e^{v^{p}(s)}\rangle+4\frac{(p-1)}{p}\int_{0}^{t}\langle(\nabla v^{\frac{p}{2}})^{2}e^{v^{p}}\rangle ds\leq\langle e^{v_{0}^{p}}\rangle,\quad p=2,4,\dots,$
provided $\frac{c_{\delta}}{\sqrt{\delta}}t<\frac{1}{2}$. The small time restriction can be removed using the semigroup property. One can compare this to the usual $L^{p}$ energy inequality (5.8) when $\delta<4$. At first sight, letting $\delta\uparrow 4$ seems to eliminate the dispersion term; however, it turns out that one retains an energy inequality once appropriate exponential factors are included. One also obtains uniqueness of weak solution to Cauchy problem for Kolmogorov PDE at least for sufficiently regular initial functions [K2].
In some sense, Orlicz space $L_{\cosh-1}$ can be viewed as the limit of $L^{p}$ spaces, i.e. as $p>\frac{2}{2-\sqrt{\delta}}$ tends to $\infty$ as $\delta\uparrow 4$.
In fact, a moment of reflection after inspecting the test function $\varphi(v)=e^{v}-1$ suggests that the theory of the Kolmogorov equation in the critical regime $\delta=4$ should be viewed as the limit $p\rightarrow\infty$ of the asymptotic $L^{p}$ theory, namely, with the non-standard $L^{p}$ test function
$\varphi(v)=\biggl(1+\frac{v}{p}\biggr)^{p}-1.$
Accordingly, one needs to study solutions $v$ of the Kolmogorov backward equation around $1$. We will address this in a subsequent paper.
Remark 5.2 (Krylov-type bound in the distributional case).
Consider the assumptions of Theorem 5.1. Then the following a priori Krylov-type bound holds. Let $U\in\mathbf{F}_{\delta_{1}}$, $\delta_{1}<\infty$, be a form-bounded function, i.e. $U\in L^{2}_{{\rm loc}}$ and
$\langle U^{2},\varphi^{2}\rangle\leq\delta_{1}\langle|\nabla\varphi|^{2}\rangle+c_{\delta_{1}}\langle\varphi^{2}\rangle,\quad\varphi\in W^{1,2}.$ (5.11)
Let $W={\rm div\,}w$ for some vector field $w$ whose components lie in $\rm BMO$}. Fix some smooth approximations $\{U_{n}\}\in[U]$, $\{W_{m}\}\in[W]$, defined as in Section 2. Fix $1<\theta<\frac{d}{d-2}$ and $p\geq 2$ such that $p>\frac{2}{2-\sqrt{\delta}}$. Then, for all $f\in\mathcal{S}$,
$\sup_{x\in\mathbb{R}^{d}}\biggl|\,\mathbb{E}_{\mathbb{P}_{x}}\int_{0}^{1}(U_{n}+W_{m})(\omega_{s})f(\omega_{s})ds\,\biggr|\leq K\|A_{m}\|_{p\theta}\vee\|A_{m}\|_{p\theta^{\prime}},$ (5.12)
where
$A_{m}=|w^{i}_{m}||\nabla f|+(1+|w^{i}_{m}|)|f|,$
and constant $K$ does not depend on $n$, $m$ or $f$. Informally, this bound shows that a solution of SDE (5.1) cannot spend too much time near the singularities of $V$ and $W$. The proof is essentially given in Proposition 8.1. (The latter is an elliptic estimate, so one needs to use identity $1=e^{\mu s}e^{-\mu s}$ to arrive at (5.12).) There we take as $U$ and $W$ the components of vector fields $b$ and $q$, but the proof extends to $U$ and $W$ right away since it does not exploit any interaction between the coefficients in the right-hand side and the drift term.
By running parabolic De Giorgi’s iterations, one refines (5.12) to
$\biggl|\,\mathbb{E}_{\mathbb{P}^{n,m}_{x}}\int_{0}^{\varepsilon}(V_{n}+W_{m})(\omega_{s})f(\omega_{s})ds\,\biggr|\leq H(\varepsilon),$ (5.13)
where $H(\varepsilon)\downarrow 0$ ($\varepsilon\downarrow 0$) is independent of $n$, $m$ and $f$ (and $x\in\mathbb{R}^{d}$). If we could place the absolute value under the integral, then, after taking $U=b^{i}$ and $W=q^{i}$, a standard argument would allow to conclude tightness of $\{\mathbb{P}_{x}^{n}\}$. However, since $W$ is a distribution, we cannot do this. Still, an argument of Hao-Zhang [HZ] shows that one can conclude tightness of $\{\mathbb{P}_{x}^{n}\}$ from, basically, (5.13), by applying Itô’s formula to $\sqrt{\sigma+|x-x_{0}|^{2}}$ with $\sigma>0$ small, which allows to control the smallness of the incremenents of solutions of the approximating SDEs after taking $\sigma\downarrow 0$. That said, [HZ] need a tightness argument since they are dealing with divergence-free super-critical drifts, while we are dealing with general critical drifts and obtain stronger convergence results for $\{\mathbb{P}_{x}^{n}\}$ provided by the theory of Feller semigroups.
Remark 5.3 (On strong solutions).
There is a well known link between the stochastic transport equation (5.5) and the SDE
$X_{t}=x-\int_{0}^{t}b(X_{r})dr+\sqrt{2}B_{t}.$ (5.14)
Namely, when $b$ is bounded and smooth, the solution $v$ to the STE (5.5) can be represented as
$v(t)=f(\Psi_{t}^{-1}),\quad t\geqslant 0,$ (5.15)
where $\Psi_{t}:\mathbb{R}^{d}\times\Omega\rightarrow\mathbb{R}^{d}$ is the stochastic flow for the SDE (5.14) i.e. there exists $\Omega_{0}\subset\Omega$, $\mathbb{P}(\Omega_{0})=1$, such that, for all $\omega\in\Omega_{0}$, $\Psi_{t}(\cdot,\omega)\Psi_{s}(\cdot,\omega)=\Psi_{t+s}(\cdot,\omega)$, $\Psi_{0}(x,\omega)=x$, and
1) for every $x\in\mathbb{R}^{d}$, the process $t\mapsto\Psi_{t}(x,\omega)$ is a strong solution to (5.14),
2) $\Psi_{t}(x,\omega)$ is continuous in $(t,x)$, $\Psi_{t}(\cdot,\omega):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ are homeomorphisms and $\Psi_{t}(\cdot,\omega)$, $\Psi_{t}^{-1}(\cdot,\omega)\in C^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$.
Beck-Flandoli-Gubinelli-Maurelli [BFGM] reversed this connection when $b$ is singular (for time-homogeneous drifts their condition reads as $|b|\in L^{d}+L^{\infty}$). They used the stochastic transport equation (5.5) to construct, for a.e. initial point $x\in\mathbb{R}^{d}$, a strong solution to SDE (5.14). Having Theorem 5.1(viii), one can extend the argument of [BFGM] to $b\in\mathbf{F}_{\delta}$, see [KSS, Remark 1]. However, since this approach excludes a measure zero set of initial points, it does not imply Theorem 5.1(vi).
Regarding recent progress strong solutions of SDEs with singular drifts, we also refer to Krylov [Kr5] who develops a different approach based on Itô-Duhamel series that can be viewed, to some extent, as the Duhamel series for the stochastic transport equation.
Remark 5.4 (Dispersion estimate, local maximum principle and gradient bounds).
We now make a few remarks regarding the theory of the Kolmogorov operator $\Delta+(b+q)\cdot\nabla$ behind SDE (5.1).
1. Under the assumptions of Theorem 5.1(i), we can descend from $C_{\infty}$ to $L^{p}$ and show that for every $p>\frac{2}{2-\sqrt{\delta}}$ the operators
$e^{-t\Lambda_{p}(b,q)}:=\biggl[e^{-t\Lambda(b,q)}\upharpoonright C^{\infty}\cap L^{p}\biggr]^{\rm clos}_{L^{p}\rightarrow L^{p}}$
are bounded on $L^{p}$ and constitute a strongly continuous semigroup. Moreover, for all $\frac{2}{2-\sqrt{\delta}}<p\leq r<\infty$,
$\|e^{-t\Lambda_{p}(b,q)}f\|_{r}\leq C_{\delta,d}e^{\omega_{p}t}t^{-\frac{d}{2}(\frac{1}{p}-\frac{1}{r})}\|f\|_{p},\quad f\in L^{p},\quad\omega_{p}=\frac{c_{\delta}}{2(p-1)}.$
The latter and the Dunford-Pettis theorem yields that $e^{-t\Lambda(b,q)}$, $t\geq 0$, are integral operators.
If additionally $\delta<1$, then $v(t):=e^{-t\Lambda_{2}(b,q)}f$, $f\in L^{2}$, is the unique weak solution to Cauchy problem
$(\partial_{t}-\Delta+(b+q)\cdot\nabla)v=0,\quad v|_{t=0}=f,$
in the standard Hilbert triple $W^{1,2}\hookrightarrow L^{2}\hookrightarrow W^{-1,2}$.
The proof of this dispersion estimates uses Nash’s argument, see e.g. [KS3, proof of Theorem 4.2]. The construction of the semigroup in $L^{p}$ follows closely [KS3, proof of Theorem 4.2]. In fact, we basically construct this semigroup in the proof of Proposition 8.3. The proof of the uniqueness of the weak solution follows the classical Lions’s argument and the compensated compactness estimate (Proposition 2.1), see [QX] for details.
2. Also under the assumptions of Theorem 5.1(i), for every $f\in L^{p\theta}\cap L^{p\theta^{\prime}}$, $u:=(\mu+\Lambda(b,q))^{-1}f$ satisfies for each $x\in\mathbb{R}^{d}$ the following local maximum principle
$\sup_{B_{\frac{1}{2}}(x)}|u|\leq K\biggl(\langle|f|^{p\theta}\rho_{x}\rangle^{\frac{1}{p\theta}}+\langle|f|^{p\theta^{\prime}}\rho_{x}\rangle^{\frac{1}{p\theta^{\prime}}}\biggr),\quad\mu>\mu_{0}>0,$
for fixed $1<\theta<\frac{d}{d-2}$ and $p\geq 2$ such that $p>\frac{2}{2-\sqrt{\delta}}$. Here $\rho_{x}(y)=(1+\sigma|y-x|)^{-\frac{d}{2}+}$ for some $\sigma>0$. The constants $K$ and $\mu_{0}$ do not depend on $f$ or $x$. This is the content of Proposition 8.4.
3. Let now $q=0$.
Assume that $\delta<\frac{4}{(d-2)^{2}}\wedge 1$. Then the unique weak solution $u$ to the elliptic equation
$(\mu-\Delta+b\cdot\nabla)u=f$
satisfies, for every $r\in[2,\frac{2}{\sqrt{\delta}}[$,
$\|\nabla u\|_{r}\leq K_{1}(\mu-\mu_{0})^{-\frac{1}{2}}\|f\|_{r},\quad\|\nabla|\nabla u|^{\frac{r}{2}}\|_{2}\leq K_{2}(\mu-\mu_{0})^{-\frac{1}{2}+\frac{1}{r}}\|f\|_{r},$ ([KoS])
$\|(\mu-\Delta)^{\frac{1}{2}+\frac{1}{s}}u\|_{r}\leq K\|(\mu-\Delta)^{-\frac{1}{2}+\frac{1}{\ell}}f\|_{r},\quad\text{ for all }2\leq\ell<r<s$ ([K6])
for all $\mu$ greater than some generic constant $\mu_{0}$.
The corresponding parabolic gradient bounds [KS8] impose more restrictive conditions on $\delta$. Namely, assume that form-bound $\delta$ satisfies, for some $r=d+\varepsilon$ (with this choice of $r$ the Sobolev embedding theorem will give Hölder continuity of solution)
$\sqrt{\delta}<\left\{\begin{array}[]{ll}\big(\sqrt{r-1}-\frac{r-2}{2}\big)\frac{2}{r}&\text{ in dimensions }d=3,4,\\ (1-\mu)\frac{r-1}{r-2}\frac{1}{r}&\text{ in dimensions }d\geq 5,\end{array}\right$.
where $0<\mu<1$, $16\mu>(1-\mu)^{4}\frac{(r-1)^{2}}{(r-2)^{4}}$. (For instance, these assumptions on $\delta$ are satisfied if $\delta<\frac{1}{d^{2}}$, $d\geq 3$.) Then the unique weak solution $v$ to Cauchy problem
$(\partial_{t}-\Delta+b\cdot\nabla)v=0,\quad v|_{t=0}=f,$
satisfies
$\sup_{0\leq s\leq t}\|\nabla v(s)\|_{r}^{r}$ $+C_{1}\int_{0}^{t}\||\nabla v|^{\frac{r-2}{2}}\partial_{s}v\|_{2}^{2}ds+C_{2}\int_{0}^{t}\langle|\nabla|\nabla v|^{\frac{r}{2}}|^{2}\rangle ds\leq e^{C_{3}t}\|\nabla f\|_{r}^{r}$ (5.16)
for constants $C_{i}>0$ ($i=1,2,3$) that depend only on $d$, $\delta$ and $c_{\delta}$.
6. Critical divergence and the constant in many-particle Hardy inequality
1. We can substantially relax condition ($\kappa_{\rm hyp}$) on the strength of attraction between the particles in Example 5.1 by employing the many-particle Hardy inequality of [HHLT] and the following variant of Theorem 5.1.
Let $b\in[L^{1}_{{\rm loc}}]^{d}$ be a vector field with divergence ${\rm div\,}b\in L^{1}_{{\rm loc}}$. Let $({\rm div\,}b)_{+}$ denote the positive part of ${\rm div\,}b$.
Definition 6.1.
We say that “potential” $({\rm div\,}b)_{+}$ is form-bounded, and write $({\rm div\,}b)^{\scriptscriptstyle 1/2}_{+}\in\mathbf{F}_{\delta_{+}}$, if
$\langle({\rm div\,}b)_{+},\varphi^{2}\rangle\leq\delta_{+}\langle|\nabla\varphi|^{2}\rangle+c_{\delta_{+}}\langle\varphi^{2}\rangle\quad\forall\,\varphi\in C_{c}^{\infty}$
for some constants $\delta_{+}$ and $c_{\delta_{+}}$.
Theorem 6.1.
Let
$\left\{\begin{array}[]{l}b\in\mathbf{F}_{\delta}\text{ with }\delta<\infty,\quad({\rm div\,}b)^{\frac{1}{2}}_{+}\in\mathbf{F}_{\delta_{+}}\text{ with }\delta_{+}<4,\quad({\rm div\,}b)_{-}\in L^{1}+L^{\infty},\\ q\in\mathbf{BMO}^{-1}.\end{array}\right$.
Let $\{b_{n}\}\in[b]^{\prime}$ (see Definition 3.4), $\{q_{m}\}\in[q]$. Then assertions (i)-(v) of Theorem 5.1 remain valid.
Example 6.1.
Let us establish weak well-posedness of particle system (5.6) using Theorem 6.1 rather than Theorem 5.1. The difference between the two theorems is in the assumptions on the drift $b(x)=(b^{1}(x),\dots,b^{N}(x))$,
$b^{i}(x):=\frac{1}{N}\sum_{j=1,j\neq i}^{N}\sqrt{\kappa}\frac{d-2}{2}\frac{x^{i}-x^{j}}{|x^{i}-x^{j}|^{2}},\quad x=(x^{1},\dots,x^{N}).$
We already know that this drift is form-bounded, but what matters in Theorem 6.1 is the form-bound of potential
$({\rm div\,}b(x))_{+}={\rm div\,}b(x)=\sqrt{\kappa}\frac{(d-2)^{2}}{N}\sum_{1\leq i<j\leq N}\frac{1}{|x^{i}-x^{j}|^{2}}.$
To verify the form-boundedness of $({\rm div\,}b)_{+}$, we invoke the many-particle Hardy inequality: for $d\geq 3$, all $N\geq 2$,
$C_{d,N}\sum_{1\leq i<j\leq N}\int_{\mathbb{R}^{dN}}\frac{|\varphi(x)|^{2}}{|x^{i}-x^{j}|^{2}}dx\leq\int_{\mathbb{R}^{dN}}|\nabla\varphi(x)|^{2}dx$ (6.1)
for all $\varphi\in W^{1,2}(\mathbb{R}^{dN})$, where, from now on, $C_{d,N}$ denotes the best possible constant in (6.1). Hence
$({\rm div\,}b)_{+}^{\frac{1}{2}}\in\mathbf{F}_{\delta_{+}},\quad\delta_{+}=\sqrt{\kappa}\frac{(d-2)^{2}}{N}C^{-1}_{d,N}.$
To the best of our knowledge, the problem of finding the exact value of $C_{d,N}$ is still open. It is not difficult to obtain a crude lower bound on $C_{d,N}$ by summing up the ordinary Hardy inequalities for the inverse square potential $x^{i}\mapsto|x^{i}-x^{j}|^{-2}$, each in its own copy of $\mathbb{R}^{d}$. However, as is pointed out by Hoffmann-Ostenhof, Hoffmann-Ostenhof, Laptev and Tidblom in [HHLT], this lower bound on $C_{d,N}$ is quite suboptimal. They provided a finer argument that gives a much better lower bound
$C_{d,N}\geq(d-2)^{2}\max\bigg\{\frac{1}{N},\frac{1}{1+\sqrt{1+\frac{3(d-2)^{2}}{2(d-1)^{2}}(N-1)(N-2)}}\bigg\}.$ (6.2)
Therefore, it suffices for us to require
$\kappa<16$ ($\kappa_{\rm hyp2}$)
which guarantees $\delta_{+}<4$ and allows us to apply Theorem 6.1.
2. We argue that the relationship between the many-particle Hardy inequality and the particle system
$X^{i}_{t}=x^{i}_{0}-\sqrt{\kappa}\frac{d-2}{2}\frac{1}{N}\sum_{j=1,j\neq i}^{N}\int_{0}^{t}\frac{X_{s}^{i}-X_{s}^{j}}{|X_{s}^{i}-X_{s}^{j}|^{2}}ds+\sqrt{2}B^{i}_{t}$ (6.3)
goes both ways. Namely, we can use the counterexample in (a’) of Section 4 to the weak well-posedness of (6.3), i.e. when the strength of attraction $\kappa$ is too large, to obtain an upper bound on the best possible constant $C_{d,N}$ in (6.1).
Theorem 6.2 (An upper bound on the constant in the many particle Hardy inequality (6.1)).
$C_{d,N}\leq\frac{d(d-2)}{N}.$
Proof.
By Theorem 6.1 and the calculation in the previous example, (6.3) has a weak solution for every initial configuration of the particles provided that
$\sqrt{\kappa}\frac{(d-2)^{2}}{N}C^{-1}_{d,N}<4.$
On the other hand, by the counterexample in (a’) of Section 4, if $\kappa>16(\frac{d}{d-2})^{2}$, then (6.3) does not have a weak solution, so we must have
$4\frac{d}{d-2}\frac{(d-2)^{2}}{N}C^{-1}_{d,N}\geq 4,$
otherwise a weak solution would exist. This gives the sought upper bound on $C_{d,N}$. ∎
In [HHLT], the authors also provided, among other results, the following upper bound:
$C_{d,N}\leq\frac{2d}{2(N-1)}\pi^{\frac{d}{2}}\Gamma\left(\frac{d}{2}\right).$ (6.4)
Their argument uses particular test functions in (6.1). Theorem 6.2 improves the dependence on the dimension $d$ in (6.4), i.e. we now have polynomial growth versus factorial growth in $d$. Of course, the simplicity of the proof of Theorem 6.2 is only seeming: we apply Theorem 6.1 whose proof uses De Giorgi’s method.
Theorem 6.2 shows that the lower bound (6.2) of [HHLT] is close to optimal, at least in high dimensions.
3. We can relax the assumptions on $b$ in Theorem 6.1 as follows.
Definition 6.2 (Multiplicative form-boundedness).
A vector field $b\in[L^{1}_{{\rm loc}}(\mathbb{R})]^{d}$ is said to be multiplicatively form-bounded if
$\langle|b|\varphi,\varphi\rangle\leq\delta\|\nabla\varphi\|_{2}\|\varphi\|_{2}+c_{\delta}\|\varphi\|^{2}_{2}\qquad\forall\,\varphi\in W^{1,2}$
for some constants $\delta$ and $c_{\delta}$ (we will see below that only their finiteness is important). This will be abbreviated as $b\in\mathbf{MF}_{\delta}$.
Once again, the constant $c_{\delta}>0$ plays a secondary role when it comes to handling local singularities of $b$ (e.g. $c_{\delta}>0$ allows to include $L^{\infty}$ drifts).
Mazya [M, Sect. 1.4.7] proved that
$\langle|b|\varphi,\varphi\rangle\leq\delta\|\nabla\varphi\|_{2}\|\varphi\|_{2}\;\;\forall\varphi\in W^{1,2}\text{ for some $\delta<\infty$}\quad\Leftrightarrow\quad\sup_{r>0,x\in\mathbb{R}^{d}}\langle|b|\mathbf{1}_{B_{r}(x)}\rangle\leq Cr^{d-1}$
for some $C<\infty$, i.e. there is a complete characterization of $\mathbf{MF}_{\delta}$ in terms of Morrey spaces:
$\cup_{\delta>0}\mathbf{MF}_{\delta}\;(\text{with $c_{\delta}=0$})\quad=\quad M_{1}$ (6.5)
(see Appendix D for the proof). Let us emphasize that for the class of form-bounded vector fields one only has inclusions
$M_{2+\varepsilon}\quad\subset\quad\cup_{\delta>0}\mathbf{F}_{\delta}\;(\text{with $c_{\delta}=0$})\quad\subset\quad M_{2},$ (6.6)
where $\varepsilon>0$ is fixed arbitrarily small, i.e. there is no complete characterization of $\mathbf{F}_{\delta}$ in terms of Morrey spaces. See discussion in Section 3.
Comparing (6.5) and (6.6), one sees that one gains quite a lot in admissible singularities of $b$ by passing from form-bounded drifts to multiplicatively form-bounded drifts. Of course, this comes at expense of imposing conditions on ${\rm div\,}b$.
Theorem 6.3.
The following are true:
  1. (i)
    (Classical martingale solutions) If
    $|b|^{\frac{1+\nu}{2}}\in\mathbf{F}_{\delta}\quad\nu\in]0,1],\quad\delta<\infty,$
    and
    $({\rm div\,}b)_{+}^{1/2}\in\mathbf{F}_{\delta_{+}},\quad\delta_{+}<4,\qquad({\rm div}\,b)_{-}\in L^{1}+L^{\infty},$ (6.7)
    then, for every $x\in\mathbb{R}^{d}$, SDE
    $X_{t}=x-\int_{0}^{t}b(X_{s})ds+\sqrt{2}B_{t}$
    has a martingale solution $\mathbb{P}_{x}$.
  2. (ii)
    (Approximation uniqueness and Markov property) If, in addition to the assumptions of (i), $b\in\mathbf{MF}_{\delta}$ for some $\delta<\infty$, then there exists $0<\gamma<1$ such that, regardless of the choice of $\{b_{n}\}\in[b]$ (defined in the same way as in Section 3.2, i.e. to preserve the structure constants of $b$), provided that $\{b_{n}\}$ additionally satisfies
    $b_{n}\rightarrow b\quad\text{ in $[L^{1+\gamma}]^{d}$,}$
    we have convergence
    $\mathbb{P}_{x}^{n}\rightarrow\mathbb{P}_{x}\quad\text{ weakly in $\mathcal{P}(\mathbf{C})$},$
    of the martingale solutions $\{\mathbb{P}_{x}^{n}\}$ to the approximating SDEs
    $X^{n}_{t}=x-\int_{0}^{t}b_{n}(X^{n}_{r})dr+\sqrt{2}B_{t}.$ (6.8)
    Furthermore, $\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$ is a Markov family.
  3. (iii)
    (Feller semigroup) Under the assumptions of (ii),
    $T_{t}f(x):=\mathbb{E}_{\mathbb{P}_{x}}[f(\omega_{t})],\quad f\in C_{\infty}$
    is a strongly continuous Feller semigroup on $C_{\infty}$, say, $T_{t}=:e^{-t\Lambda}$, where the generator $\Lambda$ is thus appropriate operator realization of the formal operator $\Delta+b\cdot\nabla$ in $C_{\infty}$.
The first two statements were proved in [KS2] and [K5], respectively. The novelty is in assertion (iii). Its proof uses the Trotter approximation theorem in the same way as the proof of Theorem 5.1.
Remark 6.1 ($L^{2}$ vs $L^{1+\gamma}$ for some $\gamma<1$).
If, in the setting of Theorem 6.3(i), we additionally require $b\in[L^{2}_{{\rm loc}}(\mathbb{R}^{d})]^{d}$, then it is also possible to prove a.e. approximation uniqueness. The last condition is actually satisfied if time-inhomogeneous $b$ is a Leray-Hopf solution of the 3D Navier-Stokes equations, i.e. then one has $b\in L^{\infty}([0,1],[L^{2}_{{\rm loc}}(\mathbb{R}^{d})]^{d})$. This was explored by a number of authors, see Appendix B. There are, however, other classes of solutions to 3D N-S equations that are not uniformly in $t$ square integrable, such as the critical class (3.5) of Koch and Tataru. So, we are interested in finding different additional conditions on $b$ that do not require square integrability, but still allow us to prove, among other results, the approximation uniqueness. This is the condition $b\in\mathbf{MF}_{\delta}$ in Theorem 6.3(ii).
The proof of the approximation uniqueness in Theorem 6.3(ii) uses an $L^{\frac{1+\gamma}{\gamma}}(\mathbb{R}^{d})$ gradient bound on solutions of the corresponding elliptic Kolmogorov equation (Lemma 8.3). It is proved by means of the Gehring-Giaquinta-Modica’s lemma (Lemma 16.1), so $\gamma$ can be estimated explicitly, see Remark 16.1.
To use Gehring-Giaquinta-Modica’s lemma, we need Caccioppoli’s inequality. The proof of Caccioppoli’s inequality for multiplicatively form-bounded drifts employs an extra iteration procedure (“Caccioppoli’s iterations”) which was introduced in our previous paper [KV] to study regularity of solutions of Dirichlet problem for the drift-diffusion equation. Namely, for $v=(u-k)_{+}$ and cutoff function $\eta$ one has
$\langle b\cdot\nabla u,\eta v\rangle$ $=\frac{1}{2}\langle b\cdot\nabla v^{2},\eta\rangle$
$=-\frac{1}{2}\langle b\cdot\nabla\eta,v^{2}\rangle$
$\leq\frac{1}{2}\langle|b|,\psi^{2}\rangle,\quad\text{ where }\psi:=\sqrt{|\nabla\eta|v}.$
By $b\in\mathbf{MF}_{\delta}$ (for simplicity, take $c_{\delta}=0$),
$\langle|b|,\psi^{2}\rangle$ $\leq\delta\|\nabla(v\sqrt{|\nabla\eta|})\|_{2}\|v\sqrt{|\nabla\eta|}\|_{2}$
$\leq\delta\bigg(\|(\nabla v)\sqrt{|\nabla\eta|}\|_{2}+\|v\nabla\sqrt{|\nabla\eta|}\|_{2}\bigg)\,\|v\sqrt{|\nabla\eta|}\|_{2},$
so
$\langle|b|,\psi^{2}\rangle\leq\frac{C_{1}}{r_{2}-r_{1}}\|(\nabla v)\mathbf{1}_{B_{r_{2}}}\|_{2}\|v\mathbf{1}_{B_{r_{2}}}\|_{2}+\frac{C_{1}}{(r_{2}-r_{1})^{2}}\|v\mathbf{1}_{B_{r_{2}}}\|^{2}_{2},$
provided that $\eta$ is equal to $1$ on $B_{r_{1}}$, is zero outside of $B_{r_{2}}$, and its derivatives satisfy appropriate estimates. The first term in the RHS contains both $\nabla v$ and the indicator function of the ball of larger radius, so we cannot simply apply Cauchy-Schwarz’ inequality to obtain the Caccioppoli inequality. But it is possible to arrive at the Caccioppoli inequality by iterating over a sequence of intermediate balls with radii between $r_{1}$ and $r_{2}$.
Remark 6.2 (Heat kernel bounds).
Although in the results mentioned so far it is the singular positive part of ${\rm div\,}b$ that presents an obstacle to the well-posedness of the SDE, something nice that can be said about the case of positive divergence. Namely, assume that $a\in H_{\xi}$, i.e. we have a bounded symmetric uniformly elliptic matrix field. Let
$b\in\mathbf{F}_{\delta},\quad\delta<4\xi^{2},$
and
${\rm div\,}b\geq 0.$
Then the heat kernel $p(t,x,y)$ of $\nabla\cdot a\cdot\nabla+b\cdot\nabla$, defined as the integral kernel of the corresponding $C_{0}$ semigroup in $L^{p}$, $p>\frac{2}{2-\nu^{-1}\sqrt{\delta}}$, constructed via a suitable regularization of $a$ and $b$, satisfies, possibly after a modification on a measure zero set, the Gaussian lower bound
$c_{1}\Gamma_{c_{2}}(t,x-y)e^{-c_{3}t}\leq p(t,x,y),$ (6.9)
where $\Gamma_{c}(t,x):=(4\pi ct)^{-\frac{d}{2}}e^{-\frac{|x|^{2}}{ct}}$ and $c_{1},c_{2}>0$, $c_{3}\geq 0$, see [KS4].
Under the above assumptions on $b$ there is no Gaussian upper bound on $p(t,x,y)$. In fact, the counterexample is given by the Brownian particles considered in Example 1.1, see the end of Section 4. Consequently, the proof of the Gaussian lower bound (6.9) does not use the Gaussian upper bound. As is mentioned to [KS4], to the best of authors’ knowledge, this is the first result of this type.
Remark 6.3 (More on the case ${\rm div\,}b=0$).
1. Let $b=b(x)$. In the case ${\rm div\,}b=0$, there is an alternative approach to the proof of the approximation uniqueness for $b\in\mathbf{MF}_{\delta}$. Namely, Mazya-Verbitsky [MV, Theorem 5] proved equivalence
$|\langle b\varphi,\varphi\rangle|\leq\delta\|\nabla\varphi\|_{2}\|\varphi\|_{2}\quad\forall\,\varphi\in C_{c}^{\infty}\quad\Leftrightarrow\quad b=\nabla Q\text{ for some }\quad Q\in[{\rm BMO}]^{d\times d},$ (6.10)
where the LHS is, clearly, more general that $b\in\mathbf{MF}_{\delta}$. So, one can use this representation for divergence-free $b$, put the anti-symmetric matrix $Q$ in the diffusion coefficients, and then prove uniqueness of the weak solution to Cauchy problem for the Kolmogorov parabolic equation by working in the standard Hilbert triple $W^{1,2}\hookrightarrow L^{2}\hookrightarrow W^{-1,2}$, see [QX].
2. Assuming that $b\in\mathbf{MF}_{\delta}$, ${\rm div\,}b=0$, Semënov [S] proved two-sided Gaussian bounds
$C_{1}\Gamma_{c_{2}}(t-s,x-y)\leq p(t,s,x,y)\leq C_{3}\Gamma_{c_{4}}(t-s,x-y),$ (6.11)
where $p(t,s,x,y)$ is the heat kernel of the Kolmogorov operator $\nabla\cdot a\cdot\nabla+b\cdot\nabla$ with measurable symmetric uniformly elliptic $a$. He used Moser’s method to prove the upper bound. His proof of the lower bound is based on a substantial modification of Nash’s method. Next, Qian-Xi [QX] established two-sided Gaussian bounds for all $q=\nabla Q\in\mathbf{BMO}^{-1}$.
Having two-sided Gaussian bounds greatly simplifies the analysis of the corresponding diffusion process, e.g.  one obtains right away the Feller propagator, the continuity of trajectories follows easily from the Kolmogorov continuity criterion. Let us show how the tightness of the martingale solutions $\mathbb{P}_{x}^{n}$ of the approximating SDEs (6.8) follows from the upper Gaussian bound (our underlying goal here is to demonstrate how natural condition $b\in\mathbf{MF}_{\delta}$ is). The tightness argument requires, as its point of departure, the estimate
$\mathbb{E}_{\mathbb{P}^{n}_{x}}\int_{t_{0}}^{t_{0}+\varepsilon}|b_{n}(\omega_{r})|dr\leq H(\varepsilon),\quad 0\leq t_{0}\leq 1,0<\varepsilon<1,$ (6.12)
for a continuous function $H$ such that $H(\varepsilon)\downarrow 0$ as $\varepsilon\downarrow 0$. By Itô’s formula, $\mathbb{E}_{\mathbb{P}^{n}_{x}}\int_{t_{0}}^{t_{1}}|b_{n}(\omega_{r})|dr=u_{n}(t_{0},X^{n}_{t_{0}}),$ where $u_{n}$ solves $\partial_{t}u_{n}+\Delta u_{n}-b_{n}\cdot\nabla u_{n}+|b_{n}|=0$, $u_{n}(t_{1})=0.$ After reversing the direction of time, we can deal instead with the initial-value problem
$(\partial_{t}-\Delta+b_{n}\cdot\nabla)v_{n}=|b_{n}|,\quad v_{n}|_{t=0}=0.$ (6.13)
(with some abuse of notation, we continue denoting the drift by $b_{n}$). By the Duhamel formula,
$v(t,x)=\int_{0}^{t}\langle p(t,s,x,\cdot)|b_{n}(\cdot)|\rangle ds,$
so, applying the upper Gaussian bound on the heat kernel $p=p_{n}$, we obtain (put $\Gamma_{t,x}:=\Gamma_{c_{4}}(t-s,x-\cdot)$):
$|v(t,x)|$ $\leq C_{3}\int_{0}^{t}\langle|b|\sqrt{\Gamma_{t,x}},\sqrt{\Gamma_{t,x}}\rangle ds$
$(\text{we use $b\in\mathbf{MF}_{\delta}$})$ (6.14)
$\leq C_{3}\int_{0}^{t}\biggl(\delta\|\nabla\sqrt{\Gamma_{t,x}}\|_{2}\|\sqrt{\Gamma_{t,x}}\|_{2}+c_{\delta}\|\sqrt{\Gamma_{t,x}}\|^{2}_{2}\biggr)ds$
$=C_{3}\int_{0}^{t}\biggl(\delta\|\nabla\sqrt{\Gamma_{t,x}}\|_{2}+c_{\delta}\biggr)ds$
$=C_{3}\int_{0}^{t}\biggl(\delta\sqrt{\frac{d}{8c_{4}}\frac{1}{t-s}}+c_{\delta}\biggr)ds,$
which yields (6.12) for $H(t)=C\sqrt{t}+c_{\delta}t$. Let us note that the upper Gaussian bound that we assumed above is valid e.g. if $b\in\mathbf{MF}_{\delta}$ and ${\rm div\,}b\leq 0$ or, more generally, $({\rm div\,}b)_{+}$ is in the Kato class, see [KS4].
Remark 6.4.
The following is a variant of the particle system in Examples 1.1, 5.1 and 6.1 that additionally pushes the center of mass of the interacting particles towards the origin:
$X^{i}_{t}=x^{i}_{0}-\frac{1}{N}\sum_{j=1,j\neq i}^{N}\sqrt{\kappa}\frac{d-2}{2}\int_{0}^{t}\frac{X_{s}^{i}-X_{s}^{j}}{|X_{s}^{i}-X_{s}^{j}|^{2}}-\sqrt{\kappa}\int_{0}^{t}\frac{(d-2)^{2}}{4}\frac{1}{N-2}\frac{\frac{1}{N}\sum_{k=1}^{N}X_{s}^{k}}{|\frac{1}{N}\sum_{k=1}^{N}X_{s}^{k}|^{2}}ds+\sqrt{2}B^{i}_{t},$
where $i=1,\dots,N$. Theorem 6.1 applies whenever $\kappa<16$. Indeed, the corresponding drift $b=(b^{1},\dots,b^{N})$ is given by
$b^{i}(x^{1},\dots,x^{N}):=\frac{\sqrt{\kappa}}{N}\sum_{j=1,j\neq i}^{N}\frac{d-2}{2}\frac{x^{i}-x^{j}}{|x^{i}-x^{j}|^{2}}+\sqrt{\kappa}\frac{(d-2)^{2}}{4}\frac{1}{N-2}\frac{\frac{1}{N}\sum_{k=1}^{N}x^{k}}{|\frac{1}{N}\sum_{k=1}^{N}x^{k}|^{2}},$
so it suffices for us to apply to ${\rm div\,}b$, in the same way as it was done in Example 6.1, the following many-particle Hardy-type inequality, also proved in [HHLT]:
$\frac{(d-2)^{2}}{N}\sum_{1\leq i<j\leq N}\int_{\mathbb{R}^{dN}}\frac{\varphi^{2}(x)}{|x^{i}-x^{j}|^{2}}dx+\frac{(d-2)^{2}N}{4}\int_{\mathbb{R}^{dN}}\frac{\varphi^{2}(x)}{\big|\sum_{k=1}^{N}x^{k}\big|^{2}}dx\leq\int_{\mathbb{R}^{dN}}|\nabla\varphi(x)|^{2}dx.$ (6.15)
The proof of (6.1) in [HHLT] departs from (6.15). In turn, the proof of the latter is rather close to the proof of the usual Hardy inequalty (i.e. via representing the potential as the divergence of appropriately chosen vector field and integrating by parts, see [HHLT, proof of Theorem 2.1] for details). Despite that, to our knowledge, the question of the optimality of the constants in inequality (6.15) has not been addressed yet.
7. Diffusion coefficients with form-bounded $\nabla a$
By [MV, Theorem 6.1], for a distributional vector field $c$ on $\mathbb{R}^{d}$, one has
$\Delta+c\cdot\nabla\in\mathcal{B}(W^{1,2},W^{-1,2}),$ (7.1)
if and only if $c$ admits a decomposition
$c=b+q\quad\text{ for some $b\in\mathbf{F}_{\delta}$ and $q\in\mathbf{BMO}^{-1}_{\sharp}$.}$ (7.2)
Here $\mathbf{BMO}^{-1}_{\sharp}$ consists of divergence-free vector fields $q=\nabla Q$ whose $n\times n$ anti-symmetric primitive $Q=(Q^{ij})_{i,j=1}^{d}$ satisfies
$\|Q^{ij}\|_{{\rm BMO}_{\sharp}}=\sup_{x\in\mathbb{R}^{d},0<R\leq 1}\frac{1}{|B_{R}|}\int_{B_{R}(x)}|Q^{ij}-(Q^{ij})_{B_{R}(x)}|dy<\infty.$
It follows that Theorem 5.1 covers exactly the same class of drifts that guarantee the embedding (7.1), up to replacing $\mathbf{BMO}^{-1}_{\sharp}$ with $\mathbf{BMO}^{-1}$, i.e. imposing rather mild assumptions on the growth of $Q$ at infinity. In fact, [MV, Theorem 1] yields a similar necessary and sufficient condition for the embedding (7.1) for the homogeneous Sobolev spaces, in which case $\mathbf{BMO}^{-1}_{\sharp}$ in (7.2) gets replaced with $\mathbf{BMO}^{-1}$, and one has $c_{\delta}=0$ in the form-boundedness condition for $b$.
As noted above, in light of the result of Mazya and Verbitsky [MV, Theorem 6.1], the drift conditions in Theorem 5.1 are essentially optimal if one expects the Kolmogorov operator $\Delta+c\cdot\nabla$ to be $W^{1,2}\rightarrow W^{-1,2}$ bounded. In Theorem A.1, however, we will be dealing with a larger than $\mathbf{F}_{\delta}$ class of Borel measurable drifts such that the Kolmogorov operator is $\mathcal{W}^{\frac{3}{2},2}\rightarrow\mathcal{W}^{-\frac{1}{2},2}$ (Bessel spaces) bounded. This apparent ambiguity and the question why one might expect Kolmogorov operator to be $W^{1,2}\rightarrow W^{-1,2}$ bounded is clarified by noting that both Theorem 5.1 and [MV, Theorem 6.1] are valid in a greater generality. Let $a$ be a bounded uniformly elliptic symmetric matrix field on $\mathbb{R}^{d}$, i.e. $a\in H_{\xi}$ for some $\xi>0$. The cited result of Mazya and Verbitsky states that one has
[…]
where $\nabla\cdot a\cdot\nabla$ does not let us deviate from the embedding $W^{1,2}\rightarrow W^{-1,2}$. In Theorem 7.1, we treat non-divergence form operators
$a\cdot\nabla^{2}+c\cdot\nabla,$
where $\nabla a\in[L^{2}_{{\rm loc}}]^{d\times d}$ is in $\mathbf{F}_{\delta}$. Since
$a\cdot\nabla^{2}+c\cdot\nabla=-\nabla\cdot a\cdot\nabla+(\nabla a+c)\cdot\nabla,$
[MV, Theorem 6.1] again applies. For such diffusion coefficients, which are considered in the next theorem, our condition on the drift is, arguably, close to being optimal.
Let $a\in H_{\xi}$. Put $\sigma=\sqrt{a}$. In Theorem 7.1 we consider SDE
$X_{t}=x-\int_{0}^{t}\big(b(X_{s})+q(X_{s})\big)ds+\sqrt{2}\int_{0}^{t}\sigma(X_{s})dB_{s},$ (7.4)
with $\nabla a+b\in\mathbf{F}_{\delta}$ and $q\in\mathbf{BMO}^{-1}$.
Example 7.1.
For example, let
$a(x)=I+c\frac{x\otimes x}{|x|^{2}},\quad x\in\mathbb{R}^{d},\quad c>-1.$
Then $\nabla a(x)=c(d-1)\frac{x}{|x|^{2}},$ so, in view of Example 3.2,
$\nabla a\in\mathbf{F}_{\delta_{1}},\quad\delta_{1}=\frac{4c^{2}(d-1)^{2}}{(d-2)^{2}}.$
This matrix field produces diffusion coefficients $\sigma$ with discontinuity at the origin that is strong enough to make the weak solution to SDE (7.4) (even with $b=q=0$) arrive at the origin with positive probability, see e.g. [B, Ch. V, Sect. 3]. This example can be extended to a matrix field $a$ in $\mathbb{R}^{dN}$ similar to the many-particle drift (1.5), i.e. corresponding to $N$ particles in $\mathbb{R}^{d}$ interacting via diffusion coefficients.
Let $\{a_{n}\},\{b_{n}\}\in[a,b]$ (Section 2), $\{q_{m}\}\in[q]$ and $\sigma_{n}=\sqrt{a_{n}}$. Define the approximating Kolmogorov operators
$\Lambda(a_{n},b_{n},q_{m}):=-a_{n}\cdot\nabla^{2}+(b_{n}+q_{m})\cdot\nabla,\quad D\big(\Lambda(a_{n},b_{n},q_{m})\big)=(1-\Delta)^{-1}C_{\infty}.$
By classical theory, these are generators of strongly continuous Feller semigroups on $C_{\infty}$. One has
$e^{-t\Lambda(a_{n},b_{n},q_{m})}f(x)=\mathbf{E}[f(X^{n,m}_{t})],$
where $X^{n,m}_{t}$ is the unique strong solution to SDE
$X^{n,m}_{t}=x-\int_{0}^{t}\big(b_{n}(X^{n,m}_{s})+q_{m}(X^{n,m}_{s})\big)ds+\sqrt{2}\int_{0}^{t}\sigma_{n}(X^{n,m}_{s})dB_{s},$ (7.5)
considered on a fixed complete probability space $\mathfrak{F}=(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbf{P})$, $B_{t}$ is a $d$-dimensional Brownian motion on this space. Set $\mathbb{P}_{x}^{n,m}:=\mathbf{P}(X_{t}^{n,m})^{-1}$.
Theorem 7.1.
Let $d\geq 3$. Let $a\in H_{\xi}$ and let $b$ and $q$ be, respectively, Borel measurable and distribution-valued vector fields $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that satisfy
$\left\{\begin{array}[]{l}\nabla a+b\in\mathbf{F}_{\delta}\text{ with }\delta<\xi^{2},\\ q\in\mathbf{BMO}^{-1}.\end{array}\right$.
Let $\{a_{n}\},\{b_{n}\}\in[a,b]$, $\{q_{m}\}\in[q]$. The following are true:
  1. (i)
    (Feller semigroup) The limit
    $s\mbox{-}C_{\infty}\mbox{-}\lim_{n}\lim_{m}e^{-t\Lambda(a_{n},b_{n},q_{m})}\text{ (loc.\,uniformly in $t\geq 0$)}$
    exists and determines a strongly continuous Feller semigroup on $C_{\infty}$, say, $e^{-t\Lambda}=e^{-t\Lambda(a,b,q)}$, where thus $\Lambda\supset-a\cdot\nabla^{2}+(b+q)\cdot\nabla$ in $C_{\infty}$.
  2. (ii)
    (A relaxed approximation uniqueness) The limit in (i) does not depend on the choice of $\{a_{n}\}$, $\{b_{n}\}$ and $\{q_{m}\}$. In fact, we can replace the strong convergence of $b_{n}$ to $b$ in $L^{2}]^{d}$ by the weak convergence
    $b_{n}\overset{w}{\rightarrow}b\quad\text{ in }[L^{2}]^{d}$
    to have convergence of the Feller semigroups $e^{-t\Lambda(a_{n},b_{n},q)}\overset{s}{\rightarrow}e^{-t\Lambda(a,b,q)}$ in $C_{\infty}$ (loc. uniformly in $t\geq 0$).
  3. (iii)
    (Generalized martingale solution) There exists a strong Markov family of probability measures $\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$ on the canonical space $\mathbf{C}$ of continuous trajectories $\omega$ such that
    $e^{-t\Lambda(a,b,q)}f(x)=\mathbb{E}_{\mathbb{P}_{x}}[f(\omega_{t})],\quad f\in C_{\infty},\quad x\in\mathbb{R}^{d},\;t\geq 0,$
    $\mathbb{P}_{x}=w\mbox{-}\mathcal{P}(\mathbf{C})\mbox{-}\lim_{n}\lim_{m}\mathbb{P}_{x}^{n,m},$
    and for every $v$ in the domain $D\big(\Lambda(a,b,q)\big)\subset C_{\infty}$ of operator $\Lambda(a,b,q)\supset-a\cdot\nabla^{2}+(b+q)\cdot\nabla$, a dense subspace of $C_{\infty}$, the process
    $t\mapsto v(\omega_{t})-v(x)+\int_{0}^{t}\Lambda(a,b,q)v(\omega_{s})ds$
    is a continuous martingale with respect to $\mathbb{P}_{x}$.
  4. (iv)
    (Weak solution for disperse initial data) Assume additionally that $b$, $q$ and $b_{n}$, $q_{m}$ have supports in a fixed ball of finite radius. Given an initial (smooth) probability density $\nu_{0}$ satisfying $\langle\nu_{0}^{2r}\rangle<\infty$ for some $1<r<\delta^{-\frac{1}{2}}$, there exist a probability space $\mathfrak{F}^{\prime}=(\Omega^{\prime},\mathcal{F}^{\prime},\{\mathcal{F}^{\prime}_{t}\}_{t\geq 0},\mathbf{P}^{\prime})$ and a continuous process $X_{t}$ on this space such that the limit
    $A_{t}:=L^{2}(\Omega^{\prime})\mbox{-}\lim_{n}\lim_{m}\int_{0}^{t}\big(b_{n}(X_{s})+q_{m}(X_{s})\big)ds$
    exists, we have a.s.
    $X_{t}=X_{0}-A_{t}+\sqrt{2}\int_{0}^{t}\sigma(X_{s})dB_{s},\quad t>0,$
    for a $\mathcal{F}_{t}^{\prime}$-Brownian motion $B_{t}$, and $\mathbf{P}^{\prime}X_{0}^{-1}$ has density $\nu_{0}$.
Theorem 7.1, when applied to $a=I$ (so $\xi=1$), imposes a more restrictive condition on the form-bound $\delta$ than Theorem 5.1, i.e. $\delta<1$ instead of $\delta<4$ (the reason for this is that in the proof of Theorem 7.1 we pass to the limit in $L^{2}$, and thus we need its $L^{2}$ theory). In fact, under the assumptions of Theorem 7.1 but with $\delta<4\xi^{2}$, after passing to a subsequence $\{n_{k}\}$ we still get Feller semigroup
$e^{-t\Lambda(a,b,q)}=s\mbox{-}C_{\infty}\mbox{-}\lim_{k}\lim_{m}e^{-t\Lambda(a_{n_{k}},b_{n_{k}},q_{m})}\text{ (loc.\,uniformly in $t\geq 0$)}.$
That is, all assertions of Theorem 7.1 with the exception of the approximation uniqueness are also valid for $\delta<4\xi^{2}$.
Further, assuming that $q=0$ and $\delta<\xi^{2}$, some other assertions of Theorem 5.1 remain valid in the setting of Theorem 7.1. See [KS2] and [KS7] regarding the weak solutions.
Earlier, Veretennikov [V] and Zhang [Z], Zhang-Zhao [ZZ2] established strong well-posedness for diffusions coefficients having derivatives in $L^{p}$ with $p$ strictly larger than $2d$ or $d$, respectively. These assumptions, however, make diffusion coefficients Hölder continuous, so they exclude e.g. Example 7.1.
Similar conditions on diffusion coefficients on the scale of Morrey spaces are considered by Krylov, see [Kr1, Kr2].
Remark 7.1.
We have a counterpart of Remark 5.4. Namely, for every $p>\frac{2}{2-\xi^{-1}\sqrt{\delta}}$, the operators
$e^{-t\Lambda_{p}(a,b,q)}:=\biggl[e^{-t\Lambda(a,b,q)}\upharpoonright C^{\infty}\cap L^{p}\biggr]^{\rm clos}_{L^{p}\rightarrow L^{p}}$
are bounded on $L^{p}$, constitute a strongly continuous semigroup, and for all $\frac{2}{2-\xi^{-1}\sqrt{\delta}}<p\leq r<\infty$,
$\|e^{-t\Lambda_{p}(a,b,q)}f\|_{r}\leq C_{\delta,d}e^{\omega_{p}t}t^{-\frac{d}{2}(\frac{1}{p}-\frac{1}{r})}\|f\|_{p},\quad f\in L^{p},$
and so, by the Dunford-Pettis theorem, $e^{-t\Lambda(a,b,q)}$, $t\geq 0$, are integral operators.
Furthermore, $v(t):=e^{-t\Lambda_{2}(a,b,q)}f$, $f\in L^{2}$, is the unique weak solution to Cauchy problem
$(\partial_{t}-a\cdot\nabla^{2}+(b+q)\cdot\nabla)v=0,\quad v|_{t=0}=f,$
in the standard Hilbert triple $W^{1,2}\hookrightarrow L^{2}\hookrightarrow W^{-1,2}$.
The local maximum principle: for every $f\in L^{p\theta}\cap L^{p\theta^{\prime}}\cap C_{\infty}$, solution $u=(\mu+\Lambda(a,b,q))^{-1}f$ to the elliptic Kolmogorov equation $(\mu-a\cdot\nabla^{2}+(b+q)\cdot\nabla)u=f$ satisfies for each $x\in\mathbb{R}^{d}$:
$\sup_{B_{\frac{1}{2}}(x)}|u|\leq K\biggl(\langle|f|^{p\theta}\rho_{x}\rangle^{\frac{1}{p\theta}}+\langle|f|^{p\theta^{\prime}}\rho_{x}\rangle^{\frac{1}{p\theta^{\prime}}}\biggr),$
for fixed $1<\theta<\frac{d}{d-2}$ and $p>\frac{2}{2-\xi^{-1}\sqrt{\delta}}$, for all $\mu$ strictly greater than certain $\mu_{0}$. The constants $K$ and $\mu_{0}$ do not depend on $f$ or $x$.
Also, a priori Krylov-type bound analogous to (5.12) holds.
The elliptic gradient bounds in the case $q=0$ and $\delta<\frac{c}{d^{2}}$ are proved in [KS7].
8. Proof of Theorem 5.1
Let us write, to shorten notations,
$\Lambda_{n,m}:=\Lambda(b_{n},q_{m})\equiv-\Delta+(b_{n}+q_{m})\cdot\nabla,\quad D(\Lambda_{n,m})=(1-\Delta)^{-1}C_{\infty}.$
Proof of (i)
This assertion will follow from the Trotter approximation theorem (see e.g. [Ka, IX.2.5]). Applied to contraction semigroups $\{e^{-t\Lambda_{n,m}}\}_{n,m\geq 1}$ in $C_{\infty}$, this theorem is stated as follows:
Theorem 8.1 (Trotter’s approximation theorem).
Assume that there exists constant $\mu_{0}>0$ independent of $n$, $m$ such that
$1^{\circ}$) $\sup_{n,m\geq 1}\|(\mu+\Lambda_{n,m})^{-1}f\|_{\infty}\leq\mu^{-1}\|f\|_{\infty}$, $\mu\geq\mu_{0}$.
$2^{\circ}$) there exists $\text{\small$s\text{-}C_{\infty}\text{-}$}\lim_{n}\lim_{m}(\mu+\Lambda_{n,m})^{-1}$ for some $\mu\geq\mu_{0}$.
$3^{\circ}$) $\mu(\mu+\Lambda_{n,m})^{-1}\rightarrow 1$ in $C_{\infty}$ as $\mu\uparrow\infty$ uniformly in $n$, $m$.
Then there exists a contraction strongly continuous semigroup $e^{-t\Lambda}$ on $C_{\infty}$ such that
$e^{-t\Lambda}=s\mbox{-}C_{\infty}\mbox{-}\lim_{n}\lim_{m}e^{-t\Lambda_{n,m}}$
locally uniformly in $t\geq 0$.
8.0.1. Main PDE results
Proposition 8.1 (Embedding property).
Let $w_{n,m}$ is the classical solution to elliptic equation
$\big(\mu-\Delta+(b_{n}+q_{m})\cdot\nabla\big)w_{n,m}=(b_{n}^{i}+q_{m}^{i})f,\quad f\in\mathcal{S}.$ (8.1)
where we have fixed $1\leq i\leq d$ and have denoted $q^{i}_{m}=\sum_{j=1}^{d}\nabla_{j}Q_{m}^{ij}.$ Put
[…]
Then, for every $p>\frac{2}{2-\sqrt{\delta}}$, $p\geq 2$ and $1<\theta<\frac{d}{d-2}$, there exist constants $\mu_{0}>0$, $0<\beta<1$ and $K_{j}$, $j=1,2$, independent of $n$, $m$, such that
$\|w_{n,m}\|_{\infty}$ $\leq K_{1}(\mu-\mu_{0})^{-\frac{\beta}{p}}\|A_{m}\|_{p\theta^{\prime}}$
$+K_{2}(\mu-\mu_{0})^{-\frac{1}{p\theta}}\|A_{m}\|_{p\theta},$ (8.2)
for all $\mu>\mu_{0}$.
We prove Proposition 8.1 in Section 12.
Set $u_{n,m}$ be the classical solutions to elliptic equation
$\big(\mu-\Delta+(b_{n}+q_{m})\cdot\nabla\big)u_{n,m}=f,\quad\mu>0.$ (8.3)
Proposition 8.2 (A priori Hölder continuity).
For every $\mu>0$, $\{u_{n,m}\}$ are locally Hölder continuous uniformly in $n$, $m$.
That is, $u_{n,m}$ are Hölder continuous, in every unit ball, with constants that do not depend on $n$, $m$ or the center of the ball. These constants are, however, allowed to depend on $\|f\|_{\infty}$. We prove Proposition 8.2 in Section 13.
Proposition 8.3 (Convergence).
There exists $\mu_{0}>0$ such that for every $\mu\geq\mu_{0}$, for all $p>\frac{2}{2-\sqrt{\delta}}$, $p\geq 2$, and any $x\in\mathbb{R}^{d}$, there exists the limit
$u:=L_{\rho_{x}}^{p}\mbox{-}\lim_{n}\lim_{m}u_{n,m},$
where $\rho_{x}(y):=\rho(y-x)$, provided that constant $\sigma$ in the definition of weight $\rho$ (this is (2.1)) is chosen sufficiently small (independently of $x$).
We prove Proposition 8.3 in Section 15.
Remark 8.1.
The proof of Proposition 8.3 can be extended to show the existence of the limit $L^{p}\mbox{-}\lim_{n,m}u_{n,m}$, but at expense of imposing additional assumptions on drifts $b$ or $q$, such as form-bound $\delta$ of $b$ being strictly less than $1$, or the stream matrix $Q$ of $q$ having entries in $\rm VMO$}. See Remark 15.2 in the end of the proof of Proposition 8.3.
Proposition 8.4 (Separation property/local maximum principle).
Fix some $1<\theta<\frac{d}{d-2}$ and $p>\frac{2}{2-\sqrt{\delta}}$, $p\geq 2$. There exists constants $K$, $\mu_{0}>0$ and $\sigma$ (in the definition of weight $\rho$) independent of $n$, $m$ such that for all $\mu\geq\mu_{0}$, for every $x\in\mathbb{R}^{d}$,
$\sup_{B_{\frac{1}{2}}(x)}|u_{n,m}|\leq K\biggl(\langle|f|^{p\theta}\rho_{x}\rangle\rangle^{\frac{1}{p\theta}}+\big\langle|f|^{p\theta^{\prime}}\mathbf{1}_{B_{1}(x)}\big\rangle^{\frac{1}{p\theta^{\prime}}}\biggr).$ (8.4)
We prove Proposition 8.4 in Section 14.
We are in position to verify conditions of Trotter’s theorem for $u_{n,m}=(\mu+\Lambda_{n,m})^{-1}f$:
8.0.2. Proof of $1^{\circ}$)
This condition is a direct consequence of the fact that, by the classical theory, $e^{-t\Lambda_{n,m}}$ are $L^{\infty}$ contractions.
8.0.3. Proof of $2^{\circ}$)
By $1^{\circ}$), it suffices to verify the existence of the limit for all $f$ belonging to a countable dense subset of $C_{c}^{\infty}$. Proposition 8.2 and the Arzelà-Ascoli theorem yield: for every $r>0$, $\{u_{n,m}\}$ is relatively compact in $C(\bar{B}_{r})$. Proposition 8.3 allows to further conclude that $\{u_{n,m}\}$ converges in $C(\bar{B}_{r})$, for every $r>0$:
$u\upharpoonright\bar{B}_{r}=s\mbox{-}C(\bar{B}_{r})\mbox{-}\lim_{n}\lim_{m}u_{m,n}\upharpoonright\bar{B}_{r}.$ (8.5)
Remark 8.2.
We need Proposition 8.3 that any two partial limits of $u_{n,m}$ in $C(\bar{B}_{r})$ coincide. The choice of the topology, and thus the weight, is secondary.
We need to improve (8.5) to global uniform convergence:
$u=s\mbox{-}C_{\infty}\mbox{-}\lim_{n}\lim_{m}u_{n,m}.$ (8.6)
To this end, we combine Proposition 8.4 and convergence (8.5). Namely, since $f\in C_{c}^{\infty}$ and weight $\rho_{x}$ vanishes at infinity, it follows from (8.4) that solution $u_{n,m}$ is small uniformly in $n$, $m$ when considered far away from the support of $f$. (Hence the name “separation property” for (8.4).)
We have verified condition $2^{\circ})$ of Trotter’s theorem.
Remark 8.3.
Using Proposition 8.4, it is easy to obtain the preservation of probability, i.e. that
$e^{-t\Lambda(b_{n},q_{m})}(1-\mathbf{1}_{B_{R}})\rightarrow 0\quad\text{ as $R\uparrow\infty$ uniformly in $n$, $m$.}$
From here it follows easily that $e^{-t\Lambda}1=1$. (We use the fact that $e^{-t\Lambda(b_{n},q_{m})}$, $e^{-t\Lambda(b,q)}$ are semigroups of integral operators, so the expressions $e^{-t\Lambda(b_{n},q_{m})}(1-\mathbf{1}_{R})$, $e^{-t\Lambda(b,q)}1$ are well-defined.)
8.0.4. Proof of $3^{\circ}$)
In view of $1^{\circ}$), it suffices to verify $3^{\circ}$) on a dense subset of $C_{\infty}$, e.g. $C_{c}^{\infty}$. Fix $g\in C_{c}^{\infty}$. By the resolvent identity,
$\mu(\mu+\Lambda_{n,m})^{-1}g-\mu(\mu-\Delta)^{-1}g$ $=\mu(\mu+\Lambda_{n,m})^{-1}(b_{n}+q_{m})\cdot\nabla(\mu-\Delta)^{-1}g$
$=(\mu+\Lambda_{n,m})^{-1}(b_{n}+q_{m})\cdot\mu(\mu-\Delta)^{-1}\nabla g.$
Since $\mu(\mu-\Delta)^{-1}g\rightarrow g$ uniformly on $\mathbb{R}^{d}$ as $\mu\rightarrow\infty$, it suffices to show the convergence
$\|(\mu+\Lambda_{n,m})^{-1}(b_{n}+q_{m})\cdot\mu(\mu-\Delta)^{-1}\nabla g\|_{\infty}\rightarrow 0\text{ as }\mu\rightarrow\infty\quad\text{ uniformly in $n$, $m$}.$ (8.7)
To that end, we apply Proposition 8.1 to $w_{n,m}:=(\mu+\Lambda_{n,m})^{-1}(b^{i}_{n}+q^{i}_{m})f$ with $f$ taken to be
$f:=\mu(\mu-\Delta)^{-1}\nabla_{i}g.$ (8.8)
Our goal is thus to prove
$\|w_{n,m}\|_{\infty}\rightarrow 0\quad\text{ as $\mu\rightarrow\infty$ uniformly in $n$, $m$.}$ (8.9)
If we can obtain bound
[…]
then the convergence (8.9) will follow thanks to the factors $(\mu-\mu_{0})^{-\frac{\beta}{p}}$, $(\mu-\mu_{0})^{-\frac{1}{p\theta}}$ in (8.2). The only slightly non-trivial aspect of proving (8.10) is that $|Q_{m}^{i}|$ can grow at infinity. But since the entries of $Q_{m}^{i}$ are $\rm BMO$} functions, this growth cannot be arbitrary, see Lemma 8.2.
Proof of (8.10). So, from now on, let $f$ be given by (8.8) where, recall, $g$ has compact support (we will need this when we apply Lemma 8.1). In what follows, for brevity, $s=p\theta$ or $s=p\theta^{\prime}$. We have
$\||Q^{i}_{m}||f|\|^{s}_{s}$ $=\langle|Q^{i}_{m}|^{s}|\mu(\mu-\Delta)^{-1}\nabla_{i}g|^{s}\rangle$
$(\text{write $1=(1+|x|)^{-(d+\epsilon_{0})s}(1+|x|)^{(d+\epsilon_{0})s}$)}$
$\leq\big\langle|Q^{i}_{m}|^{s\nu}(1+|x|)^{-(d+\epsilon_{0})s\nu}\big\rangle^{\frac{1}{\nu}}\big\langle(1+|x|)^{(d+\epsilon_{0})s\nu^{\prime}}|\mu(\mu-\Delta)^{-1}\nabla_{i}g|^{s\nu^{\prime}}\big\rangle^{\frac{1}{\nu^{\prime}}},\quad\nu>1.$ (8.11)
We will need the following elementary lemma.
Lemma 8.1.
For every $r>0$,
$|(1+|x|^{r})\mu(\mu-\Delta)^{-1}\nabla_{i}g(x)|\leq c_{R}\mu(c\mu-\Delta)^{-1}|\nabla_{i}g|(x),$ (8.12)
for some positive constants $c$, $c_{R}$ that depend only on $d$, $r$ and $R$, i.e. the radius of a fixed ball that contains the support of $|\nabla_{i}g|$.
Proof.
For reader’s convenience, we provide the proof. We estimate
$(1+|x|^{r})\mu(\mu-\Delta)^{-1}\nabla_{i}g(x)$ $=\mu\int_{0}^{\infty}\int_{\mathbb{R}^{d}}e^{-\mu s}(4\pi s)^{-\frac{d}{2}}(1+|x|^{r})e^{-\frac{|x-y|^{2}}{4s}}\nabla_{i}g(y)dyds$ (8.13)
as follows. For all $\mu\geq 1$, provided that $1\leq s<\infty$,
$e^{-\mu s}(4\pi s)^{-\frac{d}{2}}\big|(1+|x|^{r})e^{-\frac{|x-y|^{2}}{4s}}\nabla_{i}g(y)\big|\leq C_{R}e^{-c_{1}\mu s}(4\pi s)^{-\frac{d}{2}}e^{-\frac{|x-y|^{2}}{c_{2}s}}|\nabla_{i}g(y)|$ (8.14)
for some $c_{2}>4$, $0<c_{1}<1$. To see this, it suffices to show that
$e^{-\mu s}(4\pi s)^{-\frac{d}{2}}(1+|x|^{r})e^{-\frac{|x|^{2}}{4s}}\leq Ce^{-c_{1}\mu s}(4\pi s)^{-\frac{d}{2}}e^{-\frac{|x|^{2}}{c_{2}s}},\quad x\in\mathbb{R}^{d},s\geq 1,$ (8.15)
since $y$ varies only in $B_{R}$, and the sought estimate is non-trivial when $|x|\gg R$. Inequality (8.15) reduces to
$e^{-\gamma_{1}\mu s}(1+|x|^{r})e^{-\frac{|x|^{2}}{\gamma_{2}s}}\leq C$
for $0<\gamma_{1}<1$ and $\gamma_{2}>4$ ($\gamma_{1}=1-c_{1}$, $\frac{1}{\gamma_{2}}=\frac{1}{4}-\frac{1}{c_{2}}$). So, denoting $x^{\prime}=\frac{x}{\sqrt{s}}$, we obtain for $\mu\geq 1$, for all $x^{\prime}\in\mathbb{R}^{d}$, $s\geq 1$,
$e^{-\gamma_{1}\mu s}(1+|x|^{r})e^{-\frac{|x|^{2}}{\gamma_{2}s}}$ $\leq e^{-\gamma_{1}\mu s}s^{\frac{r}{2}}(1+|x^{\prime}|^{d+1})e^{-\frac{|x^{\prime}|^{2}}{\gamma_{2}}}$
$\leq Ce^{-\frac{\gamma_{1}}{2}\mu s}e^{-\frac{|x^{\prime}|^{2}}{2\gamma_{2}}}\leq C,$
which gives us the previous estimate and hence (8.15). We thus have (8.14) for all $1\leq s<\infty$. Note that (8.14) is trivial for $0<s<1$. So, armed with (8.14) for all $0<s<\infty$, we estimate (8.13):
$\mu\int_{0}^{\infty}\int_{\mathbb{R}^{d}}e^{-\mu s}(4\pi s)^{-\frac{d}{2}}(1+|x|^{r})e^{-\frac{|x-y|^{2}}{4s}}|\nabla_{i}g(y)|dyds\leq c_{R}\mu(c\mu-\Delta)^{-1}|\nabla_{i}g|(x),\quad c=4c_{1}c_{2}^{-1},$
i.e. we have obtained (8.12). ∎
We will also need
Lemma 8.2 (see e.g. [Gr, Prop. 7.1.5]).
For every $f\in{\rm BMO}(\mathbb{R}^{d})$, for all $1\leq r<\infty$,
$\langle|f-(f)_{B_{1}}|^{r}\rho^{r}\rangle\leq C_{d,r,\varepsilon}\|f\|_{{\rm BMO}}^{r},$ (8.16)
where $\rho=\rho_{\varepsilon}$ is defined by (2.1). Hence
$\langle|f|^{r}\rho^{r}\rangle\leq C^{\prime}_{d,r,\varepsilon}\bigl(\|f\|_{{\rm BMO}}^{r}+\|f\mathbf{1}_{B_{1}}\|^{r}_{L^{1}}\bigr)<\infty.$ (8.17)
By Lemma 8.1 (with $r:=d+\epsilon_{0}$), we have pointwise estimate $(1+|x|)^{(d+\epsilon_{0})}|\mu(\mu-\Delta)^{-1}\nabla_{i}g(x)|\leq c_{R}\mu(c\mu-\Delta)^{-1}|\nabla_{i}g|(x)$ with constants $_{R}, cc$ independent of $\mu\geq 1$, so we can estimate in the second multiple in (8.11):
$\big\langle(1+|x|)^{(d+\epsilon_{0})s\nu^{\prime}}|\mu(\mu-\Delta)^{-1}\nabla_{i}g|^{s\nu^{\prime}}\big\rangle\leq C\|\mu(c\mu-\Delta)^{-1}|\nabla_{i}g|\|_{s\nu^{\prime}}^{s\nu^{\prime}}.$
In turn, by the contractivity of the heat semigroup in Lebesgue spaces, for all $\mu\geq 1$,
$\|\mu(c\mu-\Delta)^{-1}|\nabla_{i}g|\|_{s\nu^{\prime}}\leq c^{-1}\|\nabla_{i}g\|_{s\nu^{\prime}}.$
Thus, the second multiple in (8.11) can be estimated as follows: for all $\mu\geq 1$,
$\big\langle(1+|x|)^{(d+\epsilon_{0})s\nu^{\prime}}|\mu(\mu-\Delta)^{-1}\nabla_{i}g|^{s\nu^{\prime}}\big\rangle\leq C\|\nabla_{i}g\|_{s\nu^{\prime}},$
where $C$ depends on the support of $g$, but does not depend on $\mu$.
On the other hand, regarding the first multiple in (8.11), we have by Lemma 8.2 (that is, (8.17) with $r=s\nu$) after taking into account that the $\rm BMO$} semi-norms of $Q^{i}_{m}$ are uniformly in $m$ bounded,
$\sup_{m}\big\langle|Q^{i}_{m}|^{s\nu}(1+|x|)^{-(d+\epsilon_{0})s\nu}\big\rangle<\infty.$
We can thus conclude from (8.11):
$\sup_{\mu\geq 1,m}\||Q^{i}_{m}||f|\|_{s}<\infty.$
In the same way, since $\nabla f=\mu(\mu-\Delta)^{-1}\nabla_{i}\nabla g$ and the entries of $\nabla g$ have compact supports,
$\sup_{\mu\geq 1,m}\||Q^{i}_{m}||\nabla f|\|_{s}<\infty.$
The last two bounds yield (8.10). In view of the previous discussion, condition $3^{\circ}$) of Trotter’s theorem is thus verified.
Now, Trotter’s theorem applies and gives us assertion (i) of Theorem 5.1.
Proof of (ii)
The first statement, i.e. the approximation uniqueness, is immediate from the existence of the limit in (i) and the fact that $[b]$, $[q]$ are closed with respect to passing to a sub-sequence. Let us prove the second statement, i.e. a relaxed approximation uniqueness. It suffices for us to replace Proposition 8.3 with the following: for all sufficiently large $\mu$ and every $f\in\mathcal{S}$, $u_{n}=(\mu+\Lambda(b_{n},q))^{-1}f$ converge to the same limit:
$u_{n}\rightarrow u=(\mu+\Lambda(b,q))^{-1}f\quad\text{ in }L^{2}_{\rho}.$ (8.18)
The rest repeats the proof of assertion (i).
Proof of (8.18). Since $b\in\mathbf{F}_{\delta}$, $\delta<1$ and ${\rm div\,}q=0$ provide, via Cauchy-Schwarz inequality and the compensatd compactness estimate, coercivity and boundedness of the quadratic form of $\Delta+(b+q)\cdot\nabla$ in $L^{2}$, it is readily seen that function $u=(\mu+\Lambda(b,q))^{-1}f$, $f\in C_{\infty}\cap L^{2}$, constructed via approximation in (i), is a weak solution to elliptic equation $(\mu-\Delta+(b+q)\cdot\nabla)u=f$, i.e. $\mu\langle u,\varphi\rangle+\langle\nabla u,\nabla\varphi\rangle+\langle\nabla u,b\varphi\rangle+\langle\nabla u,q\varphi\rangle=\langle f,\varphi\rangle$ for all $\varphi\in W^{1,2}$.
Put $h_{n}:=u_{n}-u$. Our goal is to show convergence $h_{n}\rightarrow 0$ in $L^{2}_{\rho}$. Notice that, in contrast to the proof of assertion (i), we already have the limiting object $u$, which greatly simplifies the analysis.
Step 1. Proposition 8.4 (“separation property”) extends to $u$, $u_{n}$ and $f\in\mathcal{S}$ and yields: for a fixed $1<\theta<\frac{d}{d-2}$ there exist constants $K$ and $\mu_{0}>0$ independent of $n$ such that for all $\mu\geq\mu_{0}$, for every $x\in\mathbb{R}^{d}$,
$\sup_{n\geq 1}\sup_{B_{\frac{1}{2}}(x)}|h_{n}|\leq K\biggl(\langle|f|^{2\theta}\rho_{x}\rangle\rangle^{\frac{1}{2\theta}}+\big\langle|f|^{2\theta^{\prime}}\mathbf{1}_{B_{1}(x)}\big\rangle^{\frac{1}{2\theta^{\prime}}}\biggr),$
provided constant $\sigma$ in the definition of $\rho$ is fixed sufficiently small. Since $f\in\mathcal{S}$, it follows that for every $\varepsilon>0$ there exists sufficiently large $R>0$ such that
$\sup_{n}\sup_{x\in\mathbb{R}^{d}\setminus B_{R}}|h_{n}(x)|<\varepsilon.$ (8.19)
Step 2. The difference $h_{n}=u_{n}-u$ satisfies
$\mu\langle h_{n},\varphi\rangle+\langle\nabla h_{n},\nabla\varphi\rangle+\langle b_{n}\cdot\nabla h_{n},\varphi\rangle+\langle q\cdot\nabla h_{n},\varphi\rangle=\langle(b-b_{n})\cdot\nabla u,\varphi\rangle\quad\text{ for all }\varphi\in W^{1,2}.$
Hence, taking $\varphi=h_{n}\rho$ and repeating the proof of the energy inequality of Proposition 11.1(ii) (take there $s=2$, which is possible since now $b$, $b_{n}$ have form-bound $\delta<1$), with $\sigma$ in the definition of $\rho$ fixed sufficiently small, we obtain
$(\mu-\mu_{0})\langle|h_{n}|^{2}\rho\rangle+C_{1}\langle|\nabla h_{n}|^{2}\rho\rangle\leq\langle(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle,$
where $C_{1}>0$ is independent of $n$. Thus, it suffices to show that $\langle(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle\rightarrow 0$ as $n\rightarrow\infty$.
Step 3. We represent
$\langle(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle$ $=\langle\mathbf{1}_{\mathbb{R}^{d}\setminus B_{R}}(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle+\langle\mathbf{1}_{B_{R}}(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle$ (8.20)
$=:I_{1}+I_{2}$ (8.21)
By Step 1 and $\rho\leq\sqrt{\rho}$, term $I_{1}$ can be estimated as follows: for every $\varepsilon>0$, for all $R=R(\varepsilon)>0$ sufficiently large
$|I_{1}|=|\langle\mathbf{1}_{\mathbb{R}^{d}\setminus B_{R}}(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle|\leq K_{b}\|\nabla u\|_{2}\varepsilon,$
[…]
where $\|\nabla u\|_{2}<\infty$ by the energy inequality (Proposition 11.1(i)).
Let $R$ be as above (some some small $\varepsilon>0$). Let us deal with the second term $\langle\mathbf{1}_{B_{R}}(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle$ in (8.20) . Here we work over a compact set, so the weight $\rho$ plays no role. By the energy inequality of Proposition 11.1(i), $\sup_{n}\|\nabla u_{n}\|_{2}<\infty$, and thus $\sup_{n}\|\nabla h_{n}\|_{2}<\infty$. Therefore, by the Rellich-Kondrashov theorem, there is a subsequence of $\{h_{n}\}$ (without loss of generality, $\{h_{n}\}$ itself) such that
$h_{n}\upharpoonright B_{R}\rightarrow g\text{ in $L^{2}(B_{R})$ for some $g\in(L^{2}\cap L^{\infty})(B_{R})$.}$
(Here we have used a priori estimate $\|h_{n}\|_{\infty}\leq 2\mu^{-1}\|f\|_{\infty}$.) So,
$|I_{2}|=|\langle\mathbf{1}_{B_{R}}(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle|\leq|\langle\mathbf{1}_{B_{R}}(b-b_{n})\cdot\nabla u,g\rho\rangle|+|\langle\mathbf{1}_{B_{R}}(b-b_{n})\cdot\nabla u,(h_{n}-g)\rho\rangle|,$
where
$\langle\mathbf{1}_{B_{R}}(b-b_{n})\cdot\nabla u,g\rho\rangle\rightarrow 0\quad\text{since $\mathbf{1}_{B_{R}}(\nabla u)g\rho\in L^{2}$ and $b_{n}\rightarrow b$ weakly in $L^{2}$},$
and
$|\langle\mathbf{1}_{B_{R}}(b-b_{n})\cdot\nabla u,(h_{n}-g)\rho\rangle|$ $\leq\|\mathbf{1}_{B_{R}}(b-b_{n})\sqrt{\rho}\|_{2}\|(h_{n}-g)\nabla u\|_{2}\quad(\text{we have used $\rho\leq 1$})$
$\leq K_{b}\|\mathbf{1}_{B_{R}}\nabla u\|_{2+\varepsilon_{1}}\|\mathbf{1}_{B_{R}}(h_{n}-g)\|_{\frac{2(2+\varepsilon_{1})}{\varepsilon_{1}}},\qquad\varepsilon_{1}>0.$
Next, we apply
Lemma 8.3.
$\|\nabla u\|_{2+\varepsilon_{1}}<\infty$ provided that $\varepsilon_{1}>0$ is sufficiently small.
We prove Lemma 8.3, which is of interest on its own, in Section 16. Its proof uses Gehring-Giaquinta-Modica’s lemma.
In turn,
$\|\mathbf{1}_{R}(h_{n}-g)\|_{\frac{2(2+\varepsilon_{1})}{\varepsilon_{1}}}\rightarrow 0\text{ as }n\rightarrow\infty$
follows by interpolating between $\mathbf{1}_{R}(h_{n}-g)\rightarrow$ in $L^{2}$ and $\sup_{n}\|\mathbf{1}_{R}(h_{n}-g)\|_{\infty}\leq 4\mu^{-1}\|f\|_{\infty}<\infty$.
Combining the above estimate on $I_{1}$ and the convergence $I_{2}\rightarrow 0$ as $n\rightarrow\infty$, we obtain convergence (8.18), which ends the proof of assertion (iii).
Proof of (iii)
Since $e^{-t\Lambda(b,q)}$ is a strongly continuous Feller semigroup on $C_{\infty}$, there exist probability measures $\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$ on the canonical space $\mathbf{D}_{T}$ of càdlàg trajectories $\omega$ such that
$e^{-t\Lambda(b,q)}f(x)=\mathbb{E}_{\mathbb{P}_{x}}[f(\omega_{t})],\quad f\in C_{\infty}$
and, for every $v\in D\big(\Lambda(b,q)\big)$, the process
$t\mapsto v(\omega_{t})-x+\int_{0}^{t}\Lambda(b,q)v(\omega_{s})ds$
is a martingale with respect to $\mathbb{P}_{x}$, see e.g. [RY, Ch. VII, §1]. Since both $e^{-t\Lambda(b_{n},q_{m})}$, $e^{-t\Lambda(b,q)}$ are strongly continuous Feller semigroups, the convergence of their finite-dimensional distributions, provided by assertion (i), yields
$\mathbb{P}_{x}=w\mbox{-}\mathcal{P}(\mathbf{D})\mbox{-}\lim_{n}\lim_{m}\mathbb{P}_{x}^{n,m},$
see e.g. [EK, Ch.4, Theorem 2.5]. Since $\mathbf{C}$ is closed in $\mathbf{D}$ and $\mathbb{P}_{x}^{n,m}(\mathbf{C})=1$, it follows the weak convergence that $\mathbb{P}_{x}(\mathbf{C})=1$. Thus, probability measures $\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$ are concentrated on $\mathbf{C}$, and we have
$\mathbb{P}_{x}=w\mbox{-}\mathcal{P}(\mathbf{C})\mbox{-}\lim_{n}\lim_{m}\mathbb{P}_{x}^{n,m},$ (8.22)
as claimed.
Proof of (iv)
We argue as in [HZ]. Set $\mathbb{P}_{\nu_{0}}:=\int_{\mathbb{R}^{d}}\mathbb{P}_{x}\nu_{0}(x)dx.$ Define in the same way $\mathbb{P}_{\nu_{0}}^{n,m}$. Then, in view of (8.22),
$\mathbb{P}_{\nu_{0}}=w\mbox{-}\mathcal{P}(\mathbf{C})\mbox{-}\lim_{n}\lim_{m}\mathbb{P}_{\nu_{0}}^{n,m}.$
By the Skorohod representation theorem, there exists a probability space $\mathfrak{F}^{\prime}=\{\Omega^{\prime},\mathcal{F}^{\prime},\{\mathcal{F}^{\prime}_{t}\}_{t\geq 0},\mathbf{P}^{\prime}\}$ and continuous processes $X^{n,m}_{t}$, $X_{t}$ defined on this space such that $\mathbb{P}^{n,m}_{\nu_{0}}$, $\mathbb{P}_{\nu_{0}}$ are the laws of $X^{n,m}_{t}$, $X_{t}$ and
$X_{t}^{n,m}(\omega^{\prime})\rightarrow X_{t}(\omega^{\prime}),\quad(t\geq 0,\;\omega^{\prime}\in\Omega^{\prime})$ (8.23)
(see e.g. [Bil, Ch. 1, Sect. 6]). In particular, $\mathbf{P}^{\prime}(X^{n,m}_{0})^{-1},\mathbf{P}^{\prime}X_{0}^{-1}$ have density $\nu_{0}$.
Fix $1\leq i\leq d$. Our goal is to show that
$\lim_{n_{1},n_{2}}\lim_{m_{1},m_{2}}\mathbf{E}^{\prime}\biggl|\int_{0}^{t}\big(b_{n_{1}}^{i}+q_{m_{1}}^{i}-(b^{i}_{n_{2}}+q^{i}_{m_{2}})\big)(X_{r})dr\biggr|^{2}=0.$ (8.24)
It suffices to show that
$\lim_{n_{1},n_{2}}\lim_{m_{1},m_{2}}\mathbf{E}^{\prime}\biggl|\int_{0}^{t}\big(b_{n_{1}}^{i}+q_{m_{1}}^{i}-(b_{n_{2}}^{i}+q_{m_{2}}^{i})\big)(X_{r}^{n,m})dr\biggr|^{2}=0\quad\text{ uniformly in $n$, $m$}.$ (8.25)
Indeed, having (8.25), we can appeal to (8.23) and the Dominated convergence theorem to show that, for any fixed $n_{1}$, $m_{1}$, $n_{2}$, $m_{2}$,
$\lim_{n}\lim_{m}\mathbf{E}^{\prime}\biggl|\int_{0}^{t}\big(b_{n_{1}}^{i}+q_{m_{1}}^{i}-(b_{n_{2}}^{i}+q_{m_{2}}^{i})\big)(X_{r}^{n,m})dr-\int_{0}^{t}\big(b_{n_{1}}^{i}+q_{m_{1}}^{i}-(b_{n_{2}}^{i}+q_{m_{2}}^{i})\big)(X_{r})dr\biggr|^{2}=0,$
which then yields (8.24).
So, let us prove (8.25). Put for brevity $F:=b_{n_{1}}^{i}+q_{m_{1}}^{i}-(b_{n_{2}}^{i}+q_{m_{2}}^{i})$, so (8.25) becomes
$\lim_{n_{1},n_{2}}\lim_{m_{1},m_{2}}\mathbf{E}^{\prime}\biggl|\int_{0}^{t}F(X_{r}^{n,m})dr\biggr|^{2}\rightarrow 0.$ (8.26)
Let us rewrite the expression under the limit signs as follows. Let $u$ be the classical solution to the terminal-value problem
$\big(\partial_{s}+\Delta-(b_{n}+q_{m})\cdot\nabla\big)u(s)=-F,\quad s<t,\quad u(t,\cdot)=0.$ (8.27)
Then
$\mathbf{E}^{\prime}\biggl|\int_{0}^{t}F(X_{r}^{n,m})dr\biggr|^{2}$ $=2\mathbf{E}^{\prime}\int_{0}^{t}F(X_{s}^{n,m})\int_{s}^{t}F(X_{r}^{n,m})drds$
$=2\mathbf{E}^{\prime}\int_{0}^{t}F(X_{s}^{n,m})\mathbf{E}^{\prime}\biggl[\int_{s}^{t}F(X_{r}^{n,m})dr\mid\mathcal{F}^{\prime}_{s}\biggr]ds$
$=2\mathbf{E}^{\prime}\int_{0}^{t}F(X_{s}^{n,m})u(s,X_{s}^{n,m})ds=2\int_{0}^{t}\langle Fu,\nu\rangle ds,$
where $\nu$ is the probability density of $X^{n,m}$, i.e. solution to Cauchy problem for the Fokker-Planck equation
$\partial_{t}\nu(t)-\Delta\nu(t)-{\rm div\,}\big((b_{n}+q_{m})\nu(t)\big)=0,\quad t>0,\quad\nu(0,\cdot)=\nu_{0}.$ (8.28)
Now it is seen that convergence (8.26) will follow once we prove the next lemma.
Lemma 8.4.
$I:=\int_{0}^{T}\langle(b^{i}_{n_{1}}-b^{i}_{n_{2}})u,\nu\rangle ds\rightarrow 0,\quad J:=\int_{0}^{T}\langle(q^{i}_{m_{1}}-q^{i}_{m_{2}})u,\nu\rangle ds\rightarrow 0$
as $n_{1},n_{2},m_{1},m_{2}\rightarrow\infty$ uniformly in $n$, $m$.
Proof.
Below we will use the following bound: provided that $r>1$ is chosen sufficiently close to $1$, we have
$\sup_{n,m,n_{1},n_{2},m_{1},m_{2}}\biggl[\int_{0}^{T}\langle|\nabla u|^{2}\rangle ds,\int_{0}^{T}\langle u^{4}\rangle ds,\int_{0}^{T}\langle\nu^{2r}\rangle ds,\int_{0}^{T}\langle|\nabla\nu|^{2}\rangle ds\biggr]<\infty.$ (8.29)
For the last two terms this bound follows right away from Corollary 11.2 with $s=2r$, $s=2$, respectively (since $\delta$ can be close to $1$, $s=2r$ in Corollary 11.2 must be close to $2$, hence the condition that $r$ must be close to $1$). For the first two terms this bound is obtained as follows. Since $b_{n}$, $q_{m}$ have supports in the same ball $B_{R}$ independent of $n$, $m$, we can rewrite $F$ as
$F=\big[b_{n_{1}}^{i}+q_{m_{1}}^{i}-(b_{n_{2}}^{i}+q_{m_{2}}^{i})\big]f,\quad f\in C_{c}^{\infty}\text{ is identically $1$ on }B_{R}.$
We now invoke the energy inequality of Proposition 11.2(i) for $s=2$, $s=4$, and then estimate $\langle F,u|u|^{s-2}\rangle$ in exactly the same way as in Step 1 of the proof of Proposition 8.1. (In fact, strictly speaking, in $F$ there we have $b^{i}_{n}$, $q^{i}_{m}$ instead of $b^{i}_{n_{1}}-b^{i}_{n_{2}}$ and $q^{i}_{m_{1}}-q^{i}_{m_{2}}$, but what matters in the proof is that the form-bound of $b^{i}_{n_{1}}-b^{i}_{n_{2}}$ can be chosen independently of $n_{1}$, $n_{2}$, and that the $\rm BMO$} semi-norm of $Q_{m_{1}}^{i}-Q_{m_{2}}^{i}$ is bounded from above by a constant independent of $m_{1}$, $m_{2}$, which is obviously true.) From here the bound on the first two terms in (8.29) follows.
Armed with (8.29), we estimate
$I\leq\biggl(\int_{0}^{T}\|b^{i}_{n_{1}}-b^{i}_{n_{2}}\|_{2}^{2}ds\biggr)^{\frac{1}{2}}\biggl(\int_{0}^{T}\|u\|_{2r^{\prime}}^{2r^{\prime}}ds\biggr)^{\frac{1}{2r^{\prime}}}\biggl(\int_{0}^{T}\|\nu\|_{2r}^{2r}ds\biggr)^{\frac{1}{2r}},$
with $1<r<\infty$ selected close to $1$. The first term tends to 0 as $n_{1},n_{2}\rightarrow\infty$, while the other two terms are uniformly (in $n,m,n_{1},n_{2},m_{1},m_{2}$) bounded by (8.29). In turn,
$J=-\int_{0}^{T}\langle(Q_{m_{1}}^{i}-Q_{m_{2}}^{i}),(\nabla u)\nu\rangle ds$ $-\int_{0}^{T}\langle(Q_{m_{1}}^{i}-Q_{m_{2}}^{i}),u\nabla\nu\rangle ds=:J_{1}+J_{2},$
where the stream matrices $Q_{m_{\cdot}}$ can and will be chosen to satisfy
$|Q_{m_{\cdot}}|\leq C(1+|x|)^{-d+2}\quad\forall\,|x|\geq 2R\gg 1,$ (8.30)
where $R$ is chosen so that ${\rm sprt\,}q_{m_{\cdot}}\subset B_{R}(0)$ (Appendix E). The constant $C$ does not depend on $m_{1}$, $m_{2}$. We have
$|J_{1}|^{2}$ $\leq\int_{0}^{T}\langle|\nabla u|^{2}\rangle ds\biggl(T\langle|Q_{m_{1}}^{i}-Q_{m_{2}}^{i}|^{2r^{\prime}}\rangle\biggr)^{\frac{1}{r^{\prime}}}\biggl(\int_{0}^{T}\langle\nu^{2r}\rangle ds\biggr)^{\frac{1}{r}}.$ (8.31)
The first and the last factors are uniformly bounded in view of (8.29). Therefore, since $|Q_{m_{1}}^{i}-Q_{m_{2}}^{i}|\rightarrow 0$, in particular, in $L^{2r^{\prime}}_{{\rm loc}}$, we obtain
$\langle|Q_{m_{1}}^{i}-Q_{m_{2}}^{i}|^{2r^{\prime}}\rangle\rightarrow 0\quad\text{ as }m_{1},m_{2}\rightarrow\infty.$ (8.32)
Next,
$|J_{2}|^{2}$ $\leq\biggl(T\langle|Q_{m_{1}}^{i}-Q_{m_{2}}^{i}|^{4}\rangle\biggr)^{\frac{1}{2}}\biggl(\int_{0}^{T}\langle u^{4}\rangle ds\biggr)^{\frac{1}{2}}\int_{0}^{T}\langle|\nabla\nu|^{2}\rangle ds,$
where, by the same argument as above, $\langle|Q_{m_{1}}^{i}-Q_{m_{2}}^{i}|^{4}\rangle\rightarrow 0$ as $m_{1},m_{2}\rightarrow\infty$ (note that, taking into account the above estimate on the polynomial rate of vanishing of $Q_{m_{\cdot}}$ at infinity, we have 4(d2)>d4(d-2)>d even if $d=3$, so the integrals are finite). The other two factors are uniformly bounded. This ends the proof of Lemma 8.4. ∎
Remark 8.4.
It is not difficult to remove the compact support assumption on $b$ and $q$ by inserting identity $\rho\rho^{-1}=1$ in the definitions of $I$ and $J$, i.e.
$I=\int_{0}^{T}\langle(b^{i}_{n_{1}}-b^{i}_{n_{2}})u\rho,\nu\rho^{-1}\rangle ds,\quad J=\int_{0}^{T}\langle(q^{i}_{m_{1}}-q^{i}_{m_{2}})u\rho,\nu\rho^{-1}\rangle ds$
and then arguing as above, but using the global weighted $L^{2}$ convergence result of Lemma 11.1 and replacing the energy inequalities of Section 11 (i.e. Proposition 11.2 and Corollary 11.2) by their weighted counterparts with the weights $\rho$ and $\rho^{-1}$ (the last weight is not discussed in Section 11, but the arguments do not change since $|\nabla\rho^{-1}|$ is majorated by $\rho^{-1}$, see Section 2). This comes at expense of imposing condition $\langle\nu^{2r}\rho^{-\alpha}\rangle<\infty$ for $1<r<\frac{1}{\sqrt{\delta}}$ and some $\alpha>0$.
Proof of (vii)
Put
$R^{n}_{\mu}f(x):=\mathbb{E}_{\mathbb{P}_{x}(\tilde{b}_{n})}\int_{0}^{\infty}e^{-\mu s}f(\omega_{s})ds\quad\biggl(=(\mu+\Lambda(\tilde{b}_{n}))^{-1}f(x)\biggr),\quad f\in C_{c}^{\infty},$
$R^{Q}_{\mu}f(x):=\mathbb{E}_{\mathbb{Q}_{x}}\int_{0}^{\infty}e^{-\mu s}f(\omega_{s})ds,\quad\mu>0.$
It suffices for us to show that $(\mu+\Lambda(b))^{-1}f(x)=R^{Q}_{\mu}f(x)$ for all $x\in\mathbb{R}^{d}$ and all $\mu>0$ sufficiently large, where $\Lambda(b)$ is the Feller realization of $\Delta+b\cdot\nabla$ in $C_{\infty}$ constructed in (i). This will imply that $\{\mathbb{Q}_{x}\}_{x\in\mathbb{R}^{d}}=\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$.
Step 1:
$R^{n}_{\mu}f(x)\rightarrow R_{\mu}^{Q}f(x)\text{ as }n\rightarrow\infty\quad\forall\,x\in\mathbb{R}^{d},$
as follows right away from $\mathbb{Q}_{x}=w{\mbox{-}}\mathcal{P}(\mathbf{C})\mbox{-}\lim_{n}\mathcal{P}_{x}(\tilde{b}_{n})$.
Step 2:
$\|R_{\mu}^{Q}f\|_{2}\leqslant(\mu-\mu_{0})^{-1}\|f\|_{2}\quad\text{ for all $\mu>\mu_{0}$,}$
for some $\mu_{0}$ independent of $n$. Indeed, by the elliptic energy inequality (see e.g. Proposition 11.1), $\|R_{\mu}^{n}f\|_{2}\leqslant(\mu-\mu_{0})^{-1}\|f\|_{2}$ for all $n$. Now 2) follows from 1) by a weak compactness argument in $L^{2}$.
By Step 2, operators $R_{\mu}^{Q}$ admits unique extensions by continuity to $L^{2}$, which we denote by $R_{\mu,2}^{Q}$.
On the other hand, operators $(\mu+\Lambda(b))^{-1}|_{C_{c}^{\infty}\cap L^{2}}$ are bounded on $L^{2}$ and, in fact, constitute the resolvent $\Lambda_{2}$ of the generator of a strongly continuous semigroup in $L^{2}$, i.e.
$(\mu+\Lambda_{2}(b))^{-1}:=\bigg[(\mu+\Lambda(b))^{-1}|_{C_{c}^{\infty}\cap L^{2}}\biggr]_{L^{2}\rightarrow L^{2}}^{\rm clos}.$
One can also construct $\Lambda_{2}(b)$ directly (using e.g. quadratic forms), see [KS3].
Step 3:
$\|b\cdot\nabla(\mu-\Delta)^{-1}\|_{2\rightarrow 2}\leqslant\delta,$
which follows right away from $b\in\mathbf{F}_{\delta}$ and $\|\nabla(-\Delta)^{-\frac{1}{2}}\|_{2\rightarrow 2}=1$.
Step 4:
$(\mu+\Lambda_{2}(b))^{-1}f=(\mu-\Delta)^{-1}\big(1+b\cdot\nabla(\mu-\Delta)^{-1}\big)^{-1}f\text{ in }L^{2}.$ (8.33)
Indeed, since by our assumptions we have $\delta<1$, in view of 3) the right-hand side of the previous formula is well defined. Now, we have to appeal for a moment to a “good” approximation $\{b_{n}\}\in[b]$, i.e. an approximation that actually does converge to $b$ (recall that we do not require from $\{\tilde{b}_{n}\}$ any kind of convergence to $b$).
The sought identity (8.33) holds for every $b_{n}$. This follows by rearranging the usual Neumann series representation for $(\mu+\Lambda_{2}(b_{n}))^{-1}$ while taking into account the estimate of Step 3. So,
$(\mu+\Lambda_{2}(b_{n}))^{-1}f=(\mu-\Delta)^{-1}\big(1+b_{n}\cdot\nabla(\mu-\Delta)^{-1}\big)^{-1}f.$
It remains to pass to the limit in $n$. In the left-hand side one has $(\mu+\Lambda_{2}(b_{n}))^{-1}f\rightarrow(\mu+\Lambda_{2}(b))^{-1}f$ in $L^{2}$ (see [KS2], but it is not difficult to prove this directly, see e.g. the proof of Proposition 8.3; here we need a simpler version of this in $L^{2}$). In the right-hand side the denominator of the geometric series $b_{n}\cdot\nabla(\mu-\Delta)^{-1}g\rightarrow b\cdot\nabla(\mu-\Delta)^{-1}$ in $L^{2}$ for every $g\in L^{2}$. This is immediate on $g\in C_{c}^{\infty}$, so it remains to apply the estimate of Step 3. So, we can pass to the limit in the right-hand side, arriving at the identity (8.33).
Step 5:
$(\mu+\Lambda(b))^{-1}f=R_{\mu}^{Q}f\quad\text{ a.e.~on $\mathbb{R}^{d}$.}$
Indeed, since, by our assumptions, $\mathbb{Q}_{x}$ is a weak solution of the SDE (1.1), we have by Itô’s formula
$(\mu-\Delta)^{-1}h=R_{\mu}^{Q}[\big(1+b\cdot\nabla(\mu-\Delta)^{-1}\big)h],\quad h\in C_{c}^{\infty}.$
Since $1+b\cdot\nabla(\mu-\Delta)^{-1}\in\mathcal{B}(L^{2})$ by Step 3, we have, in view of Step 2,
$(\mu-\Delta)^{-1}g=R_{\mu,2}^{Q}[\big(1+b\cdot\nabla(\mu-\Delta)^{-1}\big)g],\quad g\in L^{2}.$
Take $g=\big(1+b\cdot\nabla(\mu-\Delta)^{-1}\big)^{-1}f$, $f\in C_{c}^{\infty}$, which is possible by Step 3 and $\delta<1$. Then, by Step 4, $(\mu+\Lambda_{2}(b))^{-1}f=R_{\mu,2}^{Q}f$. By the consistency property $(\mu+\Lambda(b))^{-1}|_{C_{c}^{\infty}\cap L^{2}}=(\mu+\Lambda_{2}(b))^{-1}|_{C_{c}^{\infty}\cap L^{2}}$ and the result follows.
Step 6: Fix some $r>2\vee(d-2)$ in the interval $2,\frac{2}{\sqrt{\delta}}[$ (here we use our hypothesis on $\delta$, which must be sufficiently small so that $r$ can be large enough). Since $R_{\mu}^{n}f=(\mu+\Lambda(\tilde{b}_{n}))^{-1}f$, we obtain by assertion (xii) that for all $\mu>\mu_{0}$
$\|\nabla R_{\mu}^{n}f\|_{\frac{rd}{d-2}}\leqslant K_{2}(\mu-\mu_{0})^{-\frac{1}{2}+\frac{1}{r}}\|f\|_{r}.$
By a weak compactness argument in $L^{rj}$, in view of Step 1, we have $\nabla R_{\mu}^{Q}f\in[L^{rj}]^{d}$, and there is a subsequence of $\{R_{\mu}^{n}f\}$ (without loss of generality, it is $\{R_{\mu}^{n}f\}$ itself) such that
$\nabla R_{\mu}^{n}f\overset{w}{\longrightarrow}\nabla R_{\mu}^{Q}f\quad\text{ in }[L^{rj}]^{d}.$
By Mazur’s lemma, there is a sequence of convex combinations of the elements of $\{\nabla R_{\mu}^{n}f\}_{n=1}^{\infty}$ that converges to $\nabla R_{\mu}^{Q}f$ strongly in $L^{rj}]^{d}$, i.e.
$\sum_{\alpha}c_{\alpha}\nabla R_{\mu}^{n_{\alpha}}f\overset{s}{\longrightarrow}\nabla R_{\mu}^{Q}f\quad\text{ in $[L^{rj}]^{d}$}.$
Now, in view of the latter, Step 1 and the Sobolev embedding theorem, we have $\sum_{\alpha}c_{\alpha}R_{\mu}^{n_{\alpha}}f\overset{s}{\longrightarrow}R_{\mu}^{Q}f$ in $C_{\infty}$. Therefore, by Step 5, $(\mu+\Lambda(b))^{-1}f(x)=R_{\mu}^{Q}f(x)$ for all $x\in\mathbb{R}^{d}$, $f\in C_{c}^{\infty}$, as claimed.
9. Proof of Theorem 6.1
The proof of Theorem 6.1 follows closely the proof of Theorem 5.1, with only one calculation done differently. We stay at the level of a priori estimates, so $b$, $q$ are additionally bounded and smooth. Assuming for simplicitly that $c_{\delta}=c_{\delta_{+}}=0$ in the conditions on $b$, $({\rm div\,}b)_{+}$, we consider Cauchy problem $(\partial_{t}-\Delta+(b+q)\cdot\nabla)v=0$, $v|_{t=0}=v_{0}\in C_{c}^{\infty}$ (without loss of generality, $v_{0}\geq 0$), multiply the parabolic equation by $v^{p-1}$ and integrate by parts:
$\frac{1}{p}\langle\partial_{t}v^{p}\rangle+\frac{4(p-1)}{p^{2}}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle+\frac{2}{p}\langle b\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\rangle=0,$
where we have used ${\rm div\,}q=0$. Thus,
$\partial_{t}\langle v^{p}\rangle+\frac{4(p-1)}{p}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle=-2\langle b\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\rangle.$
In turn, $\langle b\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\rangle=\langle bv^{\frac{p}{2}},\nabla v^{\frac{p}{2}}\rangle+\langle{\rm div\,}b,v^{p}\rangle,$ hence $\langle\partial_{t}v^{p}\rangle+\frac{4(p-1)}{p}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle=\langle{\rm div\,}b,v^{p}\rangle,$ and so
$\langle\partial_{t}v^{p}\rangle+\frac{4(p-1)}{p}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle\leq\langle({\rm div\,}b)_{+},v^{p}\rangle,$
Now, applying ${\rm div\,}b\in\mathbf{F}_{\delta_{+}}$, we obtain energy inequality
$\langle\partial_{t}v^{p}\rangle+\biggl(\frac{4(p-1)}{p}-\delta_{+}\biggr)\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle\leq 0$
with $\frac{4(p-1)}{p}-\delta_{+}>0$ provided that $p>\frac{4}{4-\delta_{+}}$.
10. Proof of Theorem 6.3(iii)
We repeat the proof of Theorem 5.1(i):
Proposition 8.2 (“A priori Hölder continuity”) is replaced by [K5, Theorem 5].
Proposition 8.3 (“Convergence”) is replaced by [K5, Theorem 3(v)].
Proposition 8.4 (“Separation property”) is replaced by [K5, Propositions 5 and 6] in the proof of [K5, Theorem 5].
Proposition 8.1 (“Embedding property”) is replaced by [K5, Theorem 5].
11. Auxiliary results used in the proofs of Propositions 8.1-8.4
11.1. Energy inequalities
In the next two propositions we assume that
$\left\{\begin{array}[]{l}b\in\mathbf{F}_{\delta}\text{ with }\delta<4,\\ q\in\mathbf{BMO}^{-1}\end{array}\right$. (11.1)
(that is, we are in the assumptions of Theorem 5.1), and fix some $\{b_{n}\}\in[b]$, $\{q_{m}\}\in[q]$.
Proposition 11.1.
Assume that hypothesis (11.1) holds. Let $u=u_{n,m}$ denote the classical solution of the elliptic equation
$\big(\mu-\Delta+(b_{n}+q_{m})\cdot\nabla\big)u=f,\quad\mu\geq 0,\quad f\in C_{c}^{\infty}.$ (11.2)
Fix some $s>\frac{2}{2-\sqrt{\delta}}$, $s\geq 2.$ Then the following are true:
  1. (ii)
    There exist positive constants $\mu_{0}$, $C_{1}$ independent of $n$, $m$ such that
    $(\mu-\mu_{0})\langle|u|^{s}\rangle+C_{1}\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rangle\leq\langle f,u|u|^{s-2}\rangle$ (11.3)
    for all $\mu\geq\mu_{0}$.
  2. (iiii)
    [Weighted variant] Provided that constant $\sigma>0$ in weight $\rho(y)=(1+\sigma|y|^{2})^{-\frac{d+\epsilon_{0}}{2}}$ is fixed sufficiently small, there exist positive constants $\mu_{0}$, $C_{1}$ independent of $n$, $m$ such that, for all $x\in\mathbb{R}^{d}$,
    $(\mu-\mu_{0})\langle|u|^{s}\rho_{x}\rangle+C_{1}\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rho_{x}\rangle\leq\langle f,u|u|^{s-2}\rho_{x}\rangle,\quad\rho_{x}(y)=\rho(x-y),$ (11.4)
    for all $\mu\geq\mu_{0}$.
Remark 11.1.
We have added condition $s\geq 2$ to save ourselves some efforts since to cover the values of the form-bound $\delta$ close to $4$ we need to select $s$ large anyway. But, generally speaking, $s\in]1,2[$ does not pose a substantial difficulty, see e.g. [KS3, proof of Theorem 4.2].
In the proof of assertion (ii) we will need to control a term resulting from the interaction between the weight $\rho$ and the stream matrix $Q$ of drift $q$. This will be done using Proposition 2.2 (a compensated compactness type estimate) and Lemma 2.1 on ${\rm BMO}(\mathbb{R}^{d})$ multipliers.
Proof of Proposition 11.1.
We will only prove (ii). We write for brevity $b=b_{n}$, $q=q_{m}$. In the proof we will need estimates
$|\nabla\rho_{x}(y)|\leq\frac{d+\epsilon_{0}}{2}\frac{2\sigma|x-y|}{1+\sigma|x-y|^{2}}\rho_{x}(y)\leq\frac{d+\epsilon_{0}}{2}\sqrt{\sigma}\rho_{x}(y).$ (11.5)
We multiply equation (11.2) by $u|u|^{s-2}\rho_{x}$ and integrate by parts:
$\mu\langle|u|^{s}\rangle+\frac{4(s-1)}{s^{2}}\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rho_{x}\rangle$ $+\frac{2}{s}\langle\nabla|u|^{\frac{s}{2}},|u|^{\frac{s}{2}}\nabla\rho_{x}\rangle$
$+\frac{2}{s}\langle b\cdot\nabla|u|^{\frac{2}{s}},|u|^{\frac{2}{s}},\rho_{x}\rangle+\frac{2}{s}\langle q\cdot\nabla|u|^{\frac{s}{2}},|u|^{\frac{s}{s}}\rho_{x}\rangle=\langle f,u|u|^{s-2}\rho_{x}\rangle,$
where, taking into account anti-symmetry of the stream matrix $Q$ of $q$,
$\frac{2}{s}|\langle q\cdot\nabla|u|^{\frac{s}{2}},|u|^{\frac{s}{s}}\rho_{x}\rangle$ $=-\frac{2}{s}\langle Q\cdot\nabla|u|^{\frac{s}{2}},|u|^{\frac{s}{2}}\nabla\rho_{x}\rangle$
$=\frac{2}{s}\langle Q\cdot\nabla(|u|^{\frac{s}{2}}\sqrt{\rho_{x}}),|u|^{\frac{s}{2}}\sqrt{\rho_{x}}\frac{\nabla\rho_{x}}{\rho_{x}}\rangle$
[…]
$\leq\frac{2C}{s}\|\frac{\nabla\rho_{x}}{\rho_{x}}\cdot Q\|_{\rm BMO}\|\nabla(|u|^{\frac{s}{2}}\sqrt{\rho_{x}})\|_{2}\|u^{\frac{s}{2}}\sqrt{\rho_{x}}\|_{2},$
Applying Lemma 2.1 to $\frac{\nabla_{i}\rho(y)}{\rho(y)}=\frac{2\sigma y_{i}}{1+\sigma|y|^{2}}$, we obtain
$K:=\sup_{x\in\mathbb{R}^{d}}\|\frac{\nabla\rho_{x}}{\rho_{x}}\cdot Q\|_{\rm BMO}<\infty,$
so we can conclude the previous estimate as
$\frac{2}{s}|\langle q\cdot\nabla|u|^{\frac{s}{2}},|u|^{\frac{s}{s}}\rho_{x}\rangle$ $\leq\frac{2CK}{s}\bigl(\langle|\nabla|u|^{\frac{s}{2}}\sqrt{\rho_{x}}|^{2}\rangle+\langle|u|^{s}\rho_{x}\rangle\bigr)^{\frac{1}{2}}\langle|u|^{s}\rho_{x}\rangle^{\frac{1}{2}},$ (11.6)
and then use Cauchy-Schwarz and the second inequality in (11.5) to estimate $\langle|\nabla|u|^{\frac{s}{2}}\sqrt{\rho_{x}}|^{2}\rangle\leq(1+\varepsilon_{1})\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rho_{x}\rangle+C_{\sigma}(1+\varepsilon_{1}^{-1})\langle|u|^{s}\rho_{x}\rangle$. Here we can fix any positive $\varepsilon_{1}$ because, going back to (11.6), in the end we will apply Cauchy-Schwart inequality which will allow us to make the constant in front of the term $\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rho_{x}\rangle$ in the resulting upper bound on $\frac{2}{s}|\langle q\cdot\nabla|u|^{\frac{s}{2}},|u|^{\frac{s}{s}}\rho_{x}\rangle$ as small as we want.
Next,
$\frac{2}{s}|\langle b\cdot\nabla|u|^{\frac{s}{2}},|u|^{\frac{2}{s}}\rho_{x}\rangle|$ $\leq\frac{2}{s}\biggl(\beta\langle|b|^{2},|u|^{s}\rho_{x}\rangle+\frac{1}{4\beta}\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rho_{x}\rangle\bigg)$ (11.7)
$(\text{use $b\in\mathbf{F}_{\delta}$})$
$\leq\frac{2}{s}\biggl(\beta(\delta\|\nabla(|u|^{\frac{s}{2}}\sqrt{\rho_{x}})\|_{2}^{2}+c_{\delta}\langle|u|^{s}\rho_{x}\rangle)+\frac{1}{4\beta}\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rho_{x}\rangle\bigg).$
Take $\beta=\frac{1}{2\sqrt{\delta}}$ and then apply inequality in the end of the previous paragraph, but with $\varepsilon_{1}$ chosen small.
Applying the last inequality in (11.5), we can replace in the previous inequalities all occurrences of $|\nabla\rho_{x}|$ by $C\sqrt{\sigma}\rho_{x}$. That way, we arrive at the inequality (11.4) with constant
$C_{1}=\frac{4(s-1)}{s^{2}}-\frac{2}{s}\sqrt{\delta}-\text{(constant terms proportional to $\sqrt{\sigma}$ and $\varepsilon_{1}$)},$
where $\mu_{0}$ is given in terms of $c_{\delta}$, $\frac{CK}{s}$ and $\varepsilon_{1}^{-1}$. Since $\frac{4(s-1)}{s^{2}}-\frac{2}{s}\sqrt{\delta}>0$ ($\Leftrightarrow s>\frac{2}{2-\sqrt{\delta}}$) by our choice of $s$, we can fix $\sigma$ and $\varepsilon_{1}$ sufficiently small so that $C_{1}>0$. This ends the proof of Proposition 11.1. ∎
Corollary 11.1.
In the assumptions and notations of Proposition 11.1, we also have
  1. (ii)
    There exist positive constants $\mu_{0}$, $C_{1}$, $C_{2}$ independent of $n$, $m$ such that
    $(\mu-\mu_{0})\langle|u|^{s}\rangle+C_{1}\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rangle\leq C_{2}\langle|f|^{s}\rangle$ (11.8)
    for all $\mu\geq\mu_{0}$.
  2. (iiii)
    [Weighted variant] Provided that constant $\sigma$ in the definition of weight $\rho$ is fixed sufficiently small, there exist positive constants $\mu_{0}$, $C_{1}$, $C_{2}$ independent of $n$, $m$ such that, for all $x\in\mathbb{R}^{d}$,
    $(\mu-\mu_{0})\langle|u|^{s}\rho_{x}\rangle+C_{1}\langle|\nabla|u|^{\frac{s}{2}}|^{2}\rho_{x}\rangle\leq C_{2}\langle|f|^{s}\rho_{x}\rangle$ (11.9)
    for all $\mu\geq\mu_{0}$.
Proof.
In the above proof of Proposition 11.1 we can take one step further and apply Young’s inequality in order to estimate
$\langle f,u|u|^{s-2}\rho_{x}\rangle\leq\frac{\varepsilon^{s}}{s}\langle|u|^{s}\rho_{x}\rangle+\frac{1}{\varepsilon^{s^{\prime}}s^{\prime}}\langle|f|^{s}\rho_{x}\rangle.$
The previous energy inequalities have their parabolic counterparts:
Proposition 11.2.
Assume that hypothesis (11.1) is satisfied. Let $v=v_{n,m}$ be the classical solution of Cauchy problem
$\big(\mu+\partial_{t}-\Delta+(b_{n}+q_{m})\cdot\nabla\big)v=0,\quad v(0)=f\in C_{c}^{\infty},\quad\mu\geq 0.$ (11.10)
Fix some $s>\frac{2}{2-\sqrt{\delta}}$, $s\geq 2.$ Then there exists $\mu_{0}\geq 0$ independent of $n$, $m$ such that for all $\mu\geq\mu_{0}$ the following are true:
  1. (ii)
    $(\mu-\mu_{0})\int_{0}^{t}\langle|v|^{s}\rangle+\frac{1}{s}\sup_{r\in[0,t]}\langle|v(r)|^{s}\rangle+C_{1}\int_{0}^{t}\langle|\nabla|v|^{\frac{s}{2}}|^{2}\rangle\leq\frac{2}{s}\langle|f|^{s}\rangle$ (11.11)
    for constant $C_{1}>0$ independent of $n$, $m$.
  2. (iiii)
    [Weighted variant] Provided that $\sigma$ is the definition of weight $\rho$ is chosen sufficiently small, we have
    $(\mu-\mu_{0})\int_{0}^{t}\langle|v|^{s}\rho_{x}\rangle+\frac{1}{s}\sup_{r\in[0,t]}\langle|v(r)|^{s}\rho_{x}\rangle+C_{1}\int_{0}^{t}\langle|\nabla|v|^{\frac{s}{2}}|^{2}\rho_{x}\rangle\leq\frac{2}{s}\langle|f|^{s}\rho_{x}\rangle$ (11.12)
    for constant $C_{1}>0$ independent of $n$, $m$ and $x\in\mathbb{R}^{d}$.
Proof of Proposition 11.2.
Let us prove (ii). We multiply the parabolic equation in (11.10) by $v|v|^{s-2}\rho_{x}$ and integrate over $0,r]\times\mathbb{R}^{d}$. All the terms in the resulting integral identity, except the one containing $\partial_{t}v$, are dealt with as in the proof of the previous proposition. In turn, the term containing $\partial_{t}v$ is evaluated as follows:
$\int_{0}^{r}\langle\partial_{t}v,v|v|^{s-2}\rho_{x}\rangle=\frac{1}{s}\int_{0}^{r}\langle\partial_{t}|v|^{s}\rho_{x}\rangle=\frac{1}{s}\bigl(\langle|v(r)|^{s}\rho_{x}\rangle-\frac{1}{s}\langle|f|^{s}\rho_{x}\rangle\bigr).$
This gives us
$(\mu-\mu_{0})\int_{0}^{r}\langle|v|^{s}\rho_{x}\rangle+\frac{1}{s}\langle|v(r)|^{s}\rho_{x}\rangle+C_{1}\int_{0}^{r}\langle|\nabla|v|^{\frac{s}{2}}|^{2}\rho_{x}\rangle\leq\frac{1}{s}\langle|f|^{s}\rho_{x}\rangle.$ (11.13)
for appropriate $\mu_{0}\geq 0$ and $C>0$, both independent of $n$, $m$. Let $\mu\geq\mu_{0}$. We have, in particular,
$\frac{1}{s}\langle|v(r)|^{s}\rho_{x}\rangle\leq\frac{1}{s}\langle|f|^{s}\rho_{x}\rangle,\quad(\mu-\mu_{0})\int_{0}^{r}\langle|v|^{s}\rho_{x}\rangle+C_{1}\int_{0}^{r}\langle|\nabla|v|^{\frac{s}{2}}|^{2}\rho_{x}\rangle\leq\frac{1}{s}\langle|f|^{s}\rho_{x}\rangle.$
We can pass to the supremum in $r$ in both inequalities since their right-hand side does not depend on $r$. Adding up the resulting inequalities, we arrive at (11.12). ∎
In the proof of assertion (v) of Theorem 5.1 we use energy inequality for the Fokker-Planck equation:
Corollary 11.2.
Assume that hypothesis (11.1) is satisfied with $\delta<1$. Let $\nu=\nu_{n,m}$ denote the classical solution to Cauchy problem
$\mu\nu+\partial_{t}\nu-\Delta\nu+{\rm div}\big[(b_{n}+q_{m})\nu\big]=0,\quad\nu(0)=\nu_{0}\in C_{c}^{\infty},\quad\mu\geq 0.$ (11.14)
Fix some $2\leq s<\frac{2}{\sqrt{\delta}}$. Then there exist positive constants $\mu_{0}$, $C_{1}$ independent of $n$, $m$ such that
$(\mu-\mu_{0})\int_{0}^{t}\langle|\nu|^{s}\rangle+\frac{1}{s}\sup_{r\in[0,t]}\langle|\nu(r)|^{s}\rangle+C_{1}\int_{0}^{t}\langle|\nabla|\nu|^{\frac{s}{2}}|^{2}\rangle\leq\frac{2}{s}\langle|\nu_{0}|^{s}\rangle$
for all $\mu\geq\mu_{0}$.
One also has a straightforward weighted counterpart of this inequality as in assertion (ii) of Proposition 11.2.
We will be applying Corollary 11.2 in the case when $\nu_{0}\geq 0$, $\langle\nu_{0}\rangle=1$.
Proof.
All the terms in the corresponding integral identity, except the next one, are dealt with in the same way as in the proof of Proposition 11.1. Let $b=b_{n}$. Then
$\langle{\rm div}(b\nu),\nu|\nu|^{s-2}\rangle=(s-1)\langle b\nu,|\nu|^{s-2}\nabla\nu\rangle=\frac{2}{s}(s-1)\langle b\cdot\nabla|\nu|^{\frac{s}{2}},|\nu|^{\frac{s}{2}}\rangle.$
Hence we arrive at the counterpart of (11.13) with the coefficient of the dispersion term
$C=\frac{4(s-1)}{s^{2}}-\frac{2}{s}(s-1)\sqrt{\delta}-\text{(constant terms proportional to $\varepsilon_{1}$)}$
that must be positive. We can fix $\varepsilon_{1}$ as small as needed. What matters is the value of $\delta$ that ensures that $\frac{4(s-1)}{s^{2}}-\frac{2}{s}(s-1)\sqrt{\delta}>0$. The latter is equivalent to $s<\frac{2}{\sqrt{\delta}}$, which is satisfied by our assumptions. ∎
11.2. Global weighted $L^{2}$ summability of a form-bounded drift
A form-bounded vector field $b\in\mathbf{F}_{\delta}$ is a priori only local summable: $|b|\in L^{2}_{{\rm loc}}$. In fact, condition $b\in\mathbf{F}_{\delta}$ implies global square summability of $|b|$, but with respect to weight $\rho_{x}$. Indeed, selecting $\sqrt{\rho_{x}}$ as the test function in the definition of $b\in\mathbf{F}_{\delta}$ and using (2.2), we obtain
$\langle|b|^{2}\rho_{x}\rangle$ $\leq\frac{\delta}{4}\langle\frac{|\nabla\rho_{x}|^{2}}{\rho_{x}}\rangle+c_{\delta}\langle\rho_{x}\rangle$
$\leq\biggl(\frac{\delta}{4}\frac{(d+\varepsilon_{0})^{2}}{4}\sigma+c_{\delta}\biggr)\langle\rho\rangle<\infty,$ (11.15)
where we have used $\langle\rho_{x}\rangle=\langle\rho\rangle$. Moreover, we have the following global convergence result:
Lemma 11.1.
Let $b\in\mathbf{F}_{\delta}$, $\{b_{n}\}\in[b]$. Then, for every $x\in\mathbb{R}^{d}$, $\langle|b_{n_{1}}-b_{n_{2}}|^{2}\rho_{x}\rangle\rightarrow 0$ as $n_{1}$, $n_{2}\rightarrow\infty$.
Proof.
First, let us show that
$\lim_{R\rightarrow\infty}\langle|b_{n_{1}}-b_{n_{2}}|^{2}\mathbf{1}_{\mathbb{R}^{d}\setminus B_{R+1}}\rho_{x}\rangle\rightarrow 0\quad\text{ uniformly in $n_{1}$, $n_{2}$.}$ (11.16)
Indeed, replacing $\mathbf{1}_{\mathbb{R}^{d}\setminus B_{R+1}}$ by greater function $\eta^{2}_{R}$, where
$\eta_{R}(y):=\xi_{R}(|y|),\quad\xi_{R}(r):=\left\{\begin{array}[]{ll}0,&0\leq r<R,\\ r-R,&R\leq r\leq R+1,\\ 1,&r>R+1,\end{array}\right$.
and noting that $|\nabla\eta_{R}(y)|\leq\mathbf{1}_{R<|y|<R+1}$, we estimate
$\langle|b_{n_{1}}-b_{n_{2}}|^{2}\mathbf{1}_{\mathbb{R}^{d}\setminus B_{R+1}}\rho_{x}\rangle$ $\leq\langle|b_{n_{1}}-b_{n_{2}}|^{2}\eta^{2}_{R}\rho_{x}\rangle$
$\leq 4\delta\|\nabla(\eta_{R}\sqrt{\rho_{x}})\|_{2}^{2}+4c_{\delta}\|\eta_{R}\sqrt{\rho_{x}}\|_{2}^{2},$
where $\|\eta_{R}\sqrt{\rho_{x}}\|_{2}\rightarrow 0$ as $R\rightarrow\infty$, and so, in view of the second estimate in (11.5), $\|\eta_{R}\nabla(\sqrt{\rho_{x}})\|_{2}\rightarrow 0$ as $R\rightarrow\infty$. Also, taking without loss of generality $x=0$,
$\|(\nabla\eta_{R})\sqrt{\rho}\|_{2}^{2}=\langle\mathbf{1}_{R\leq|\cdot|\leq R+1}\rho\rangle=CR^{-d-\varepsilon}R^{d}=R^{-\varepsilon}\rightarrow 0$
as $R\rightarrow\infty$. This yields (11.16).
In turn, inside the ball $B_{R}$ we have $\langle|b_{n_{1}}-b_{n_{2}}|^{2}\mathbf{1}_{B_{R}}\rho_{x}\rangle\rightarrow 0$ as $n_{1},n_{2}\rightarrow\infty$ since $b_{n_{1}}-b_{n_{2}}\rightarrow 0$ in $L^{2}_{\rm loc}$. This ends the proof. ∎
12. Proof of Proposition 8.1 (Embedding property)
We will need the following lemma.
Lemma 12.1 ([G, Lemma 7.1]).
If $\{z_{i}\}_{i=0}^{\infty}\subset\mathbb{R}_{+}$ is a sequence of positive real numbers such that
$z_{i+1}\leq NC_{0}^{i}z^{1+\alpha}_{i}$
for some $C_{0}>1$, $\alpha>0$, and
$z_{0}\leq N^{-\frac{1}{\alpha}}C_{0}^{-\frac{1}{\alpha^{2}}}.$
Then $\lim_{i}z_{i}=0.$
Throughout the proof, we write for brevity
$b=b_{n},\quad q=q_{m},\quad Q^{i}=Q_{m}^{i},\quad w=w_{n,m}.$
It suffices to estimate the positive part of $w$:
$\sup_{\mathbb{R}^{d}}w_{+}$ $\leq K_{1}(\mu-\mu_{0})^{-\frac{\beta}{p}}\big\langle|Q^{i}|^{p\theta^{\prime}}(|\nabla f|^{p\theta^{\prime}}+|f|^{p\theta^{\prime}})\big\rangle^{\frac{1}{p\theta^{\prime}}}$
$+K_{2}(\mu-\mu_{0})^{-\frac{1}{p\theta}}\langle|Q^{i}|^{p\theta}(|\nabla f|^{p\theta}+|f|^{p\theta})\rangle^{\frac{1}{p\theta}}.$ (12.1)
Step 1. First, we prove the following energy inequality: there exist generic constants $\mu_{0}\geq 0$ and $_{0}, CC$ such that, for every $k\geq 0$ and all $\mu>\mu_{0}$, the positive part $v:=(w-k)_{+}$ of wkw-k satisfies
$(\mu-\mu_{0})\|v^{\frac{p}{2}}\|_{2}^{2}+C_{0}\|\nabla v^{\frac{p}{2}}\|_{2}^{2}\leq C\bigl[\langle|Q^{i}|^{p}|\nabla f|^{p}\mathbf{1}_{v>0}\rangle+\langle(1+|Q^{i}|^{p})|f|^{p}\mathbf{1}_{v>0}\rangle\bigr].$ (12.2)
Proof of (12.2). We obtain from equation (8.1), using that both $\mu$ and $k$ are non-negative, $(\mu-\Delta+(b+q)\cdot\nabla)(w-k)\leq(b^{i}+q^{i})f.$ We now basically apply the energy inequality of Proposition 11.1(i) with $s=p$. The fact that we have an elliptic differential inequality instead of an elliptic equation does not change anything since we multiply it by a non-negative function $v^{p-1}$. So, we integrate, apply $b\in\mathbf{F}_{\delta}$ and use anti-symmetry of the stream matrix $Q$ of drift $q$:
$(\mu-\mu_{0})\langle v^{p}\rangle+C_{1}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle\leq\langle(b^{i}+q^{i})f,v^{p-1}\rangle.$ (12.3)
Let us estimate the terms in the right-hand side. First,
$|\langle b^{i}f,v^{p-1}\rangle|$ $\leq\varepsilon\langle|b^{i}|^{2},v^{p}\rangle+\frac{1}{4\varepsilon}\langle|f|^{2}v^{p-2}\rangle$
$\leq\varepsilon\biggl(\delta\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle+c_{\delta}\langle v^{p}\rangle\biggr)$
$+\frac{1}{4\varepsilon}\biggl(\frac{p-2}{p}\langle v^{p}\rangle+\frac{2}{p}\langle|f|^{p}\mathbf{1}_{v>0}\rangle\biggr).$
Second,
$|\langle q^{i}f,v^{p-1}\rangle|$ $=|\langle{\rm div\,}Q^{i},fv^{p-1}\rangle|$
$\leq|\langle Q^{i},(\nabla f)v^{p-1}\rangle|+\frac{2(p-1)}{p}|\langle Q^{i},fv^{\frac{p}{2}-1}\nabla v^{\frac{p}{2}}\rangle|=:K_{1}+\frac{2(p-1)}{p}K_{2},$
where
$K_{1}$ $\leq\frac{1}{2}\langle|Q^{i}|^{2}|\nabla f|^{2}v^{p-2}\rangle+\frac{1}{2}\langle v^{p}\rangle$
$\leq\frac{1}{2}\biggl(\frac{2}{p}\langle|Q^{i}|^{p}|\nabla f|^{p}\mathbf{1}_{v>0}\rangle+\frac{p-2}{p}\langle v^{p}\rangle\biggr)+\frac{1}{2}\langle v^{p}\rangle$
and
$K_{2}$ $\leq\varepsilon\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle+\frac{1}{4\varepsilon}\langle|Q^{i}|^{2}f^{2}v^{p-2}\rangle$
$\leq\varepsilon\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle+\frac{1}{4\varepsilon}\biggl(\frac{2}{p}\langle|Q^{i}|^{p}|f|^{p}\mathbf{1}_{v>0}\rangle+\frac{p-2}{p}\langle v^{p}\rangle\biggr).$
Thus,
$|\langle q^{i}f,v^{p-1}\rangle|\leq\frac{2(p-1)}{p}\varepsilon\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle+c_{1}\bigl(\langle|Q^{i}|^{p}|\nabla f|^{p}\mathbf{1}_{v>0}\rangle+\langle|Q^{i}|^{p}|f|^{p}\mathbf{1}_{v>0}\rangle\bigr)+c_{2}\langle v^{p}\rangle,$
where $c_{1}$, $c_{2}$ depend on $\varepsilon$ and $p$.
Substituting the resulting estimates in (12.3) and selecting $\varepsilon$ sufficiently small, we obtain estimate (12.2).
Step 2. In what follows, we will be selecting $k>0$. Then $|\{v>0\}|<\infty$. We obtain from (12.2), using the Sobolev embedding theorem,
$(\mu-\mu_{0})\|v\|_{p}^{p}+C_{S}\|v\|^{p}_{\frac{pd}{d-2}}\leq C\big\langle\big(|Q^{i}|^{p}|\nabla f|^{p}+(1+|Q^{i}|^{p})|f|^{p}\big)\mathbf{1}_{v>0}\big\rangle.$ (12.4)
We estimate the left-hand side of (12.4) using the interpolation inequality:
$(\mu-\mu_{0})^{\beta}\|v\|_{p\theta_{0}}^{p}\leq\beta(\mu-\mu_{0})\|v\|_{p}^{p}+(1-\beta)\|v\|_{L^{\frac{pd}{d-2}}}^{p},\quad 0<\beta<1,\quad\frac{1}{p\theta_{0}}=\beta\frac{1}{p}+(1-\beta)\frac{d-2}{pd},$
where $1<\theta_{0}<\frac{d}{d-2}$. So,
$(\mu-\mu_{0})^{\beta}\|v\|_{p\theta_{0}}^{p}\leq C_{2}\bigl[\big\langle\big(|Q^{i}|^{p}|\nabla f|^{p}+(1+|Q^{i}|^{p})|f|^{p}\big)\mathbf{1}_{v>0}\big\rangle\bigr]$
Let us fix $\beta$ small enough so that we have $\theta_{0}>\theta$. (Recall that $1<\theta<\frac{d}{d-2}$ was fixed in the statement of the proposition.) Applying Hölder’s inequality, we obtain
$(\mu-\mu_{0})^{\beta}\|v\|_{p\theta_{0}}^{p}\leq C_{3}H|\{v>0\}|^{\frac{1}{\theta}},$
where
$H:=\big\langle\big(|Q^{i}|^{p\theta^{\prime}}|\nabla f|^{p\theta^{\prime}}+(1+|Q^{i}|^{p\theta^{\prime}})|f|^{p\theta^{\prime}}\big)\mathbf{1}_{v>0}\big\rangle^{\frac{1}{\theta^{\prime}}}.$
On the other hand, again by Hölder’s inequality,
$\|v\|_{p\theta}^{p\theta}\leq\|v\|_{p\theta_{0}}^{p\theta}|\{v>0\}|^{1-\frac{\theta}{\theta_{0}}}.$
Therefore, we obtain
$\|v\|_{p\theta}^{p\theta}\leq\tilde{C}(\mu-\mu_{0})^{-\beta\theta}H^{\theta}|\{v>0\}|^{2-\frac{\theta}{\theta_{0}}}.$
Step 3. Now, put $v_{m}:=(w-k_{m})_{+}$, $k_{m}:=\xi(1-2^{-m})\uparrow\xi,$ where constant $\xi>0$ will be chosen later.
Remark 12.1.
We have $k_{0}=0$, but we will not encounter the volume of $\{u>0\}$ in the proof (clearly, $|\{u>0\}|$ can be infinite).
So,
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{p\theta}^{p\theta}\leq\tilde{C}\frac{1}{\xi^{p\theta}}(\mu-\mu_{0})^{-\beta\theta}H^{\theta}|\{u>k_{m+1}\}|^{2-\frac{\theta}{\theta_{0}}}.$
From now on, we require constant $\xi$ to satisfy $\xi^{p}\geq(\mu-\mu_{0})^{-\beta}H$, so
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{L^{p\theta}(B^{m+1})}^{p\theta}\leq\tilde{C}|\{u>k_{m+1}\}|^{2-\frac{\theta}{\theta_{0}}}.$
Now,
$|\{w>k_{m+1}\}|$ $=\bigg|\bigg\{\biggl(\frac{w-k_{m}}{k_{m+1}-k_{m}}\biggr)^{p\theta}>1\bigg\}\bigg|$
$\leq(k_{m+1}-k_{m})^{-p\theta}\langle v^{p\theta}_{m}\rangle=\xi^{-p\theta}2^{p\theta(m+1)}\|v_{m}\|_{p\theta}^{p\theta},$
so, applying the previous two inequalities, we obtain
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{p\theta}^{p\theta}\leq C2^{p\theta m(2-\frac{\theta}{\theta_{0}})}\biggl(\frac{1}{\xi^{p\theta}}\|v_{m}\|_{L^{p\theta}(B^{m})}^{p\theta}\biggr)^{2-\frac{\theta}{\theta_{0}}}.$
Step 4. Denote $z_{m}:=\frac{1}{\xi^{p\theta}}\|v_{m}\|_{p\theta}^{p\theta}.$ Then
$z_{m+1}\leq C\gamma^{m}z_{m}^{1+\alpha},\quad m=0,1,2,\dots,\quad\alpha:=1-\frac{\theta}{\theta_{0}},\;\;\gamma:=2^{p\theta(2-\frac{\theta}{\theta_{0}})}$
and $z_{0}=\frac{1}{\xi^{p\theta}}\langle w_{+}^{p\theta}\rangle\leq C^{-\frac{1}{\alpha}}\gamma^{-\frac{1}{\alpha^{2}}}$ provided that we fix $\xi$ by
$\xi^{p\theta}:=C^{\frac{1}{\alpha}}\gamma^{\frac{1}{\alpha^{2}}}\langle w_{+}^{p\theta}\rangle+(\mu-\mu_{0})^{-\beta\theta}H^{\theta}$
(so that it also satisfies the previous requirement $\xi^{p}\geq(\mu-\mu_{0})^{-\beta}H$). Hence, by Lemma 12.1, $z_{m}\rightarrow 0$ as $m\rightarrow\infty$. Therefore, $w_{+}\leq\xi.$ Thus, we obtain inequality
$\sup_{\mathbb{R}^{d}}w_{+}\leq K\biggl(\langle w_{+}^{p\theta}\rangle^{\frac{1}{p\theta}}+(\mu-\mu_{0})^{-\frac{\beta}{p}}\big\langle\big(|Q^{i}|^{p\theta^{\prime}}|\nabla f|^{p\theta^{\prime}}+(1+|Q^{i}|^{p\theta^{\prime}})|f|^{p\theta^{\prime}}\big)\mathbf{1}_{v>0}\big\rangle^{\frac{1}{p\theta^{\prime}}}\biggr).$ (12.5)
Step 5. It remains to estimate $\langle w_{+}^{p\theta}\rangle^{\frac{1}{p\theta}}$. We already did this in (12.2) (use $p\theta>p$):
$(\mu-\mu_{0})^{\frac{1}{p\theta}}\|w_{+}\|_{p\theta}\leq C^{\frac{1}{p\theta}}\bigl[\langle|Q^{i}|^{p\theta}|\nabla f|^{p\theta}\rangle+\langle(1+|Q^{i}|^{p\theta})f^{p\theta}\rangle\bigr]^{\frac{1}{p\theta}}.$
This inequality, applied in (12.5), yields (12.1) and thus ends the proof of Proposition 8.1. ∎
13. Proof of Proposition 8.2 (Hölder continuity)
The following is a well-known consequence of the John-Nirenberg inequality:
Proposition 13.1.
There exists constant $C=C(d)$ such that, for every $g\in{\rm BMO}$,
$\sup_{x\in\mathbb{R}^{d},R>0}\frac{1}{|B_{R}(x)|}\langle e^{\frac{C|g-(g)_{B_{R}(x)}|}{\|g\|_{\rm BMO}}}\mathbf{1}_{B_{R}(x)}\rangle\leq C.$ (13.1)
In particular, for every $1\leq s<\infty$, we have $g\in L^{s}_{{\rm loc}}$ and
$\sup_{x\in\mathbb{R}^{d},R>0}\frac{1}{|B_{R}(x)|}\langle|g-(g)_{B_{R}(x)}|^{s}\mathbf{1}_{B_{R}(x)}\rangle\leq C(d,s)\|g\|_{{\rm BMO}}^{s}.$ (13.2)
Put for brevity $b=b_{n}$, $q=q_{m}$, $Q=Q_{m}$ and $u=u_{n,m}$. So,
$\big(\mu-\Delta+(b+q)\cdot\nabla\big)u=f.$
Let us fix some $1<\theta<\frac{d}{d-2}$. Let $p>\frac{2}{2-\sqrt{\delta}}$. By our assumption $\delta<4$, so such $p$ exist.
Lemma 13.1 (Caccioppoli’s inequality).
Let $v:=(u-k)_{+}$, $k\in\mathbb{R}$. For all $0<r<R\leq 1$, we have
$\|\nabla v^{\frac{p}{2}}\|_{L^{2}(B_{r})}^{2}\leq\frac{K_{1}}{(R-r)^{2}}|B_{R}|^{\frac{1}{\theta^{\prime}}}(1+\|Q\|_{{\rm BMO}}^{2})\|v^{\frac{p}{2}}\|_{L^{2\theta}(B_{R})}^{2}+K_{2}\||f-\mu u|^{\frac{p}{2}}\mathbf{1}_{v>0}\|_{L^{2}(B_{R})}^{2}$ (13.3)
for constants $K_{1}$, $K_{2}$ independent of $k$, $r$, $R$ and $n$, $m$.
Proof.
Let $\{\eta=\eta_{r,R}\}$ be a family of $[0,1]$-valued smooth cut-off functions satisfying
$\eta=1\text{ in }B_{r},\quad\eta=0\text{ in }\mathbb{R}^{d}\setminus B_{R},$ (13.4)
$|\nabla\eta|\leq\frac{c}{R-r}\mathbf{1}_{B_{R}},\quad\frac{|\nabla\eta|^{2}}{\eta}\leq\frac{c}{(R-r)^{2}}\mathbf{1}_{B_{R}}$ (13.5)
with constant $c$ independent of $r$, $R$. We rewrite the equation for $u$ in the form $(-\Delta+(b+q)\cdot\nabla\big)u=f-\mu u,$ multiply it by $v^{p-1}\eta$ and integrate, obtaining
$\frac{4(p-1)}{p^{2}}\langle\nabla v^{\frac{p}{2}},(\nabla v^{\frac{p}{2}})\eta\rangle$ $+\frac{2}{p}\langle\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\nabla\eta\rangle$
$\leq-\frac{2}{p}\langle b\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\eta\rangle-\frac{2}{p}\langle q\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\eta\rangle+\langle f-\mu u,v^{p-1}\eta\rangle.$
Hence, by Cauchy-Schwarz,
$\left(\frac{4(p-1)}{p}-\frac{4}{p}\epsilon\right)\langle|\nabla v^{\frac{p}{2}}|^{2}\eta\rangle$ $\leq\frac{p}{4\epsilon}\big\langle v^{p}\frac{|\nabla\eta|^{2}}{\eta}\big\rangle-2\langle b\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\eta\rangle-2\langle q\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\eta\rangle+p\langle f-\mu u,v^{p-1}\eta\rangle$
$=:I_{1}+I_{2}+I_{3}+I_{4}.$ (13.6)
Let us estimate terms $I_{1}$-$I_{4}$. We start with the term $I_{3}$ containing the distribution-valued vector field $q=\nabla Q$. The other terms $I_{1}$, $I_{2}$ and $I_{4}$ will be estimated in such a way as to fit the estimate on $I_{3}$. The following argument was used in [H]. Set $\tilde{Q}^{ij}:=Q^{ij}-(Q^{ij})_{R}$, where, recall, $(Q^{ij})_{R}=(Q^{ij})_{B_{R}}$ is the average of $Q^{ij}$ over ball $B_{R}$. We have
$I_{3}=-2\langle q\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\eta\rangle$ $=-\sum_{i=1}^{d}\langle q^{i}\nabla_{i}v^{p},\eta\rangle$
$(\text{use identity $q^{i}=\sum_{j=1}^{d}\nabla_{j}Q^{ij}=\sum_{j=1}^{d}\nabla_{j}\tilde{Q}^{ij}$ and integrate by parts})$
$=\sum_{i,j=1}^{d}\langle\tilde{Q}^{ij}\nabla_{j}\nabla_{i}v^{p},\eta\rangle+\sum_{i,j=1}^{d}\langle\tilde{Q}^{ij}\nabla_{i}v^{p},\nabla_{j}\eta\rangle$
$=\sum_{i,j=1}^{d}\langle\tilde{Q}^{ij}\nabla_{j}\nabla_{i}v^{p},\eta\rangle+2\sum_{i,j=1}^{d}\langle\tilde{Q}^{ij}\nabla_{i}v^{\frac{p}{2}},v^{\frac{p}{2}}\nabla_{j}\eta\rangle.$
Due to the anti-symmetry of $Q$, the first sum on the right hand side is zero, so $I_{3}=2\langle\tilde{Q}\cdot\nabla v^{\frac{p}{2}},v^{\frac{p}{2}}\nabla\eta\rangle.$ Hence
$|I_{3}|\leq\varepsilon_{1}\langle|\nabla v^{\frac{p}{2}}|^{2}\eta\rangle+\frac{1}{\varepsilon_{1}}\langle|\tilde{Q}|^{2}v^{p}\frac{|\nabla\eta|^{2}}{\eta}\rangle.$ (13.7)
The second term in the RHS of (13.7) is bounded as follows, using (13.5):
$\langle|\tilde{Q}|^{2}v^{p}\frac{|\nabla\eta|^{2}}{\eta}\rangle$ $\leq\frac{c}{(R-r)^{2}}\left\langle|\tilde{Q}|^{2\theta^{\prime}}\mathbf{1}_{B_{R}}\right\rangle^{\frac{1}{\theta^{\prime}}}\langle v^{p\theta}\mathbf{1}_{B_{R}}\rangle^{\frac{1}{\theta}}$
(use (13.2))
$\leq\frac{C}{(R-r)^{2}}|B_{R}|^{\frac{1}{\theta^{\prime}}}\|Q\|_{{\rm BMO}}^{2}\langle v^{p\theta}\mathbf{1}_{B_{R}}\rangle^{\frac{1}{\theta}}.$ (13.8)
Thus, to summarize, we have
$|I_{3}|\leq\varepsilon_{1}\langle|\nabla v^{\frac{p}{2}}|^{2}\eta\rangle+\frac{1}{\varepsilon_{1}}\frac{C}{(R-r)^{2}}|B_{R}|^{\frac{1}{\theta^{\prime}}}\|Q\|_{{\rm BMO}}^{2}\langle v^{p\theta}\mathbf{1}_{B_{R}}\rangle^{\frac{1}{\theta}}.$
Remark 13.1.
If the entries of $Q$ are in $L^{\infty}$, then we can take $\theta=1$. Indeed, in this case we can obtain (13) directly: $\langle|\tilde{Q}|^{2}v^{p}\frac{|\nabla\eta|^{2}}{\eta}\rangle\leq\frac{4c}{(R-r)^{2}}\|Q\|^{2}_{\infty}\langle v^{p}\mathbf{1}_{B_{R}}\rangle$.
Now, we estimate the remaining terms $I_{1}$, $I_{2}$ and $I_{4}$. By (13.5),
$I_{1}\leq\frac{cp}{4\epsilon(R-r)^{2}}|B_{R}|^{\frac{1}{\theta^{\prime}}}\langle v^{p\theta}\mathbf{1}_{B_{R}}\rangle^{\frac{1}{\theta}}.$
Next,
$\frac{1}{2}|I_{2}|$ $\leq\langle|b||\nabla v^{\frac{p}{2}}|,v^{\frac{p}{2}}\eta\rangle\leq\alpha\langle|\nabla v^{\frac{p}{2}}|\eta\rangle+\frac{1}{4\alpha}\langle|b|^{2},v^{p}\eta\rangle,\quad\alpha=\frac{\sqrt{\delta}}{2}$
$(\text{use $b\in\mathbf{F}_{\delta}$})$
$\leq\frac{\sqrt{\delta}}{2}\langle|\nabla v^{\frac{p}{2}}|^{2}\eta\rangle+\frac{1}{2\sqrt{\delta}}\biggl(\delta\langle|\nabla(v^{\frac{p}{2}}\sqrt{\eta})|^{2}\rangle+c_{\delta}\langle v^{p}\eta\rangle\biggr)$
$\leq\frac{\sqrt{\delta}}{2}\langle|\nabla v^{\frac{p}{2}}|^{2}\eta\rangle+\frac{\sqrt{\delta}}{2}\left((1+\varepsilon_{0})\langle|\nabla v^{\frac{p}{2}}|^{2}\eta\rangle+\frac{1}{4}\big(1+\frac{1}{\varepsilon_{0}}\big)\langle v^{p}\frac{|\nabla\eta|^{2}}{\eta}\rangle\right)+\frac{c_{\delta}}{2\sqrt{\delta}}\langle v^{p}\eta\rangle$
$\leq\frac{\sqrt{\delta}}{2}(2+\varepsilon_{0})\langle|\nabla v^{\frac{p}{2}}|\eta\rangle+\left(\frac{\sqrt{\delta}}{8}(1+\frac{1}{\varepsilon_{0}})\frac{c}{(R-r)^{2}}+\frac{c_{\delta}}{2\sqrt{\delta}}\right)|B_{R}|^{\frac{1}{\theta^{\prime}}}\langle v^{p\theta}\mathbf{1}_{B_{R}}\rangle^{\frac{1}{\theta}}.$
Finally, by Young’s inequality, for every $\varepsilon_{2}>0$,
$|I_{4}|=|\langle f-\mu u,v^{p-1}\eta\rangle|\leq\frac{1}{p\varepsilon_{2}^{p}}\langle|f-\mu u|^{p}\mathbf{1}_{\{v>0\}}\mathbf{1}_{B_{R}}\rangle+\frac{\varepsilon_{2}^{p^{\prime}}}{p^{\prime}}|B_{R}|^{\frac{1}{\theta^{\prime}}}\langle v^{p\theta}\mathbf{1}_{B_{R}}\rangle^{\frac{1}{\theta}}.$
Now, applying these estimates on $I_{1}$-$I_{4}$ in (13), we obtain:
$\left(\frac{4(p-1)}{p}-\frac{4}{p}\varepsilon-(2+\varepsilon_{0})\sqrt{\delta}-\varepsilon_{1}\right)\langle|\nabla v^{\frac{p}{2}}|^{2}\eta\rangle$ $\leq\frac{C_{1}}{(R-r)^{2}}|B_{R}|^{\frac{1}{\theta^{\prime}}}(1+\|Q\|_{{\rm BMO}}^{2})\langle v^{p\theta}\mathbf{1}_{B_{R}}\rangle^{\frac{1}{\theta}}$
$+C_{2}\||f-\mu u|^{\frac{p}{2}}\mathbf{1}_{v>0}\mathbf{1}_{B_{R}}\|_{2}^{2},$
where $\varepsilon$, $\varepsilon_{0}$, $\varepsilon_{1}$ are fixed sufficiently small so that the expression in the brackets in the LHS is strictly positive. The latter is possible since $\delta<4$ and $p>\frac{2}{2-\sqrt{\delta}}$. This ends the proof of Lemma 13.1. ∎
Lemma 13.2.
Fix $\alpha>0$ by $\alpha(\alpha+1)=1-\frac{\theta(d-2)}{d}$. Then, for all $0<r<R\leq 1$,
$\sup_{B_{\frac{R}{2}}}u\leq C\left(\frac{1}{|B_{R}|}\langle u^{p\theta}\mathbf{1}_{B_{R}\cap\{u>0\}}\rangle\right)^{\frac{1}{p\theta}}\left(\frac{|B_{R}\cap\{u>0\}|}{|B_{R}|}\right)^{\frac{\alpha}{p\theta}}+R^{\frac{2}{p}}$ (13.9)
for $C$ independent of $n$, $m$, $r$ and $R$.
Proof.
Step 1. Fix a family of cut-off functions $\eta=\eta_{r,R}\in C_{c}^{\infty}$ such that
$\eta=1\quad\text{on }B_{r}\qquad\eta=0\quad\text{on }\mathbb{R}^{d}\setminus{{\overline{B}}_{\frac{R+r}{2}}},$ (13.10)
and
$|\nabla\eta|\leq\frac{c}{R-r}\mathbf{1}_{B_{\frac{r+R}{2}}},\quad\frac{|\nabla\eta|^{2}}{\eta}\leq\frac{c}{(R-r)^{2}}\mathbf{1}_{B_{\frac{r+R}{2}}}$ (13.11)
with constant $c$ independent of $r$, $R$. Set $v=(u-k)_{+}$, where $k\in\mathbb{R}$ will be chosen later. Using Sobolev’s embedding theorem, we have
$\langle v^{\frac{pd}{d-2}}\mathbf{1}_{B_{r}}\rangle^{\frac{d-2}{d}}$ $\leq\langle(v^{\frac{p}{2}}\eta^{\frac{1}{2}})^{\frac{2d}{d-2}}\mathbf{1}_{B_{\frac{R+r}{2}}}\rangle^{\frac{d-2}{d}}$
$\leq C_{S}^{2}\langle|\nabla(v^{\frac{p}{2}}\eta^{\frac{1}{2}})|^{2}\mathbf{1}_{B_{\frac{R+r}{2}}}\rangle$
$\leq 2C_{S}^{2}\left(\langle|\nabla v^{\frac{p}{2}}|^{2}\eta\mathbf{1}_{B_{\frac{R+r}{2}}}\rangle+\langle v^{p}\frac{|\nabla\eta|^{2}}{\eta}\mathbf{1}_{B_{\frac{R+r}{2}}}\rangle\right).$
We rewrite the latter, after applying Hölder’s inequality in the second term, as follows:
$\|v^{\frac{p}{2}}\mathbf{1}_{B_{r}}\|^{2}_{\frac{2d}{d-2}}\leq C_{1}\left(\|(\nabla v^{\frac{p}{2}})\mathbf{1}_{B_{\frac{R+r}{2}}}\|_{2}^{2}+\frac{|B_{R}|^{\frac{1}{\theta^{\prime}}}}{(R-r)^{2}}\|v^{\frac{p}{2}}\mathbf{1}_{B_{\frac{R+r}{2}}}\|^{2}_{2\theta}\right).$
Next, we apply Lemma 13.1 to the first term in the RHS, obtaining
$\|v^{\frac{p}{2}}\mathbf{1}_{B_{r}}\|^{2}_{\frac{2d}{d-2}}$ $\leq C_{2}\biggl(\frac{|B_{R}|^{\frac{1}{\theta^{\prime}}}}{(R-r)^{2}}\|v^{\frac{p}{2}}\mathbf{1}_{B_{R}}\|^{2}_{2\theta}+\||f-\mu u|^{\frac{p}{2}}\mathbf{1}_{v>0}\mathbf{1}_{B_{R}}\|_{2}^{2}\biggr).$
Next, applying Hölder’s inequality in the second term in the RHS, we estimate
$\||f-\mu u|^{\frac{p}{2}}\mathbf{1}_{v>0}\mathbf{1}_{B_{R}}\|_{2}^{2}\leq\|f-\mu u\|^{p}_{\infty}|B_{R}\cap\{v>0\}|^{\frac{1}{\theta}}|B_{R}|^{\frac{1}{\theta^{\prime}}},$
which gives us, upon noting that $\|f-\mu u\|_{\infty}\leq 2\|f\|_{\infty}$,
$\|v^{\frac{p}{2}}\mathbf{1}_{B_{r}}\|^{2}_{\frac{2d}{d-2}}\leq C_{2}\biggl(\frac{|B_{R}|^{\frac{1}{\theta^{\prime}}}}{(R-r)^{2}}\|v^{\frac{p}{2}}\mathbf{1}_{B_{R}}\|^{2}_{2\theta}+|B_{R}|^{\frac{1}{\theta^{\prime}}}2^{p}\|f\|_{\infty}^{p}|B_{R}\cap\{u>k\}|^{\frac{1}{\theta}}\biggr).$
Representing $|B_{R}|^{\frac{1}{\theta^{\prime}}}=|B_{R}|^{\frac{d-2}{d}+\frac{2}{d}-\frac{1}{\theta}}$, and dividing both sides of the previous inequality by $|B_{R}|^{\frac{d-2}{d}}$, we have
$\begin{split}\frac{1}{|B_{R}|^{\frac{d-2}{d}}}\|v^{\frac{p}{2}}\mathbf{1}_{B_{r}}\|^{2}_{\frac{2d}{d-2}}&\leq C_{2}\biggl(|B_{R}|^{\frac{2}{d}}\frac{1}{(R-r)^{2}}\frac{1}{|B_{R}|^{\frac{1}{\theta}}}\|v^{\frac{p}{2}}\mathbf{1}_{B_{R}}\|^{2}_{2\theta}\\ &\quad+|B_{R}|^{\frac{2}{d}}2^{p}\|f\|_{\infty}^{p}\left(\frac{|B_{R}\cap\{u>k\}|}{|B_{R}|}\right)^{\frac{1}{\theta}}\biggr).\end{split}$ (13.12)
Next, note that if $h<k$, then $(u-k)_{+}\leq(u-h)_{+}$. Therefore, by Chebyshev’s inequality,
$|B_{R}\cap\{u>k\}|^{\frac{1}{\theta}}\leq\frac{1}{(k-h)^{p}}\langle(u-h)_{+}^{p\theta}\mathbf{1}_{B_{R}}\rangle^{\frac{1}{\theta}}.$
Recalling that $v=(u-k)_{+}$ and applying the last inequality in (13.12), we obtain
$\bigg(\frac{\langle(u-k)_{+}^{\frac{pd}{d-2}}\mathbf{1}_{B_{r}}\rangle}{|B_{R}|}\bigg)^{\frac{d-2}{d}}\leq C_{3}|B_{R}|^{\frac{2}{d}}\biggl(\frac{1}{(R-r)^{2}}+\frac{1}{(k-h)^{p}}\bigg)\bigg(\frac{\langle(u-h)_{+}^{p\theta}\mathbf{1}_{B_{R}}\rangle}{|B_{R}|}\biggr)^{\frac{1}{\theta}}.$
Now, using Hölder’s inequality and then applying the previous estimate, we get
$\frac{\langle(u-k)^{p\theta}_{+}\mathbf{1}_{B_{r}}\rangle}{|B_{R}|}\leq\bigg(\frac{\langle(u-k)^{\frac{pd}{d-2}}_{+}\mathbf{1}_{B_{r}}\rangle}{|B_{R}|}\bigg)^{\frac{\theta(d-2)}{d}}\bigg(\frac{|B_{r}\cap\{u>k\}|}{|B_{R}|}\bigg)^{1-\frac{\theta(d-2)}{d}}$
$\leq C_{3}^{\theta}|B_{R}|^{\frac{2\theta}{d}}\left(\frac{1}{(R-r)^{2}}+\frac{1}{(k-h)^{p}}\right)^{\theta}\bigg(\frac{\langle(u-h)_{+}^{p\theta}\mathbf{1}_{B_{R}}\rangle}{|B_{R}|}\bigg)\bigg(\frac{|B_{R}\cap\{u>k\}|}{|B_{R}|}\bigg)^{1-\frac{\theta(d-2)}{d}}$
Multiplying this inequality by $\left(\frac{|B_{r}\cap\{u>k\}|}{|B_{R}|}\right)^{\alpha}\left[\leq\frac{1}{(k-h)^{p\theta\alpha}}\left(\frac{\langle(u-h)_{+}^{p\theta}\mathbf{1}_{B_{R}}\rangle}{|B_{R}|}\right)^{\alpha}\right]$, and selecting $\alpha>0$ as in the statement of the lemma, i.e. such that $\alpha(\alpha+1)=1-\frac{\theta(d-2)}{d}$, we get
$\frac{\left\langle(u-k)^{p\theta}_{+}\mathbf{1}_{B_{r}}\right\rangle}{|B_{R}|}\bigg(\frac{|B_{r}\cap\{u>k\}|}{|B_{R}|}\bigg)^{\alpha}$
$\leq C_{3}^{\theta}|B_{R}|^{\frac{2\theta}{d}}\left(\frac{1}{(R-r)^{2}}+\frac{1}{(k-h)^{p}}\right)^{\theta}\frac{1}{(k-h)^{p\theta\alpha}}\left(\frac{\langle(u-h)_{+}^{p\theta}\mathbf{1}_{B_{R}}\rangle}{|B_{R}|}\left(\frac{|B_{R}\cap\{u>k\}|}{|B_{R}|}\right)^{\alpha}\right)^{1+\alpha}$
At the next step, we are going to iterate this inequality.
Step 2. Now, define
$r_{m}:=\frac{R}{2}\left(1+\frac{1}{2^{m}}\right),\quad B_{m}:=B_{r_{m}}$
$k_{m}:=\xi(1-2^{-m}),$
for a positive constant $\xi$ to be determined later. Setting
$z_{m}\equiv z(k_{m},r_{m}):=\frac{\left\langle(u-k_{m})^{p\theta}_{+}\mathbf{1}_{B_{m}}\right\rangle}{|B_{R}|}\bigg(\frac{|B_{m}\cap\{u>k_{m}\}|}{|B_{R}|}\bigg)^{\alpha},$
we rewrite the previous inequality, upon selecting $k:=k_{m+1}$ and $h:=k_{m}$ there, as
$z_{m+1}\leq C_{3}^{\theta}\frac{|B_{R}|^{\frac{2\theta}{d}}}{R^{2\theta}}\left(2^{2m}+2^{mp}\frac{R^{2}}{\xi^{p}}\right)^{\theta}\frac{2^{mp\theta\alpha}}{\xi^{p\theta\alpha}}z^{1+\alpha}_{m}.$ (13.13)
In what follows, we restrict out choice of constant $\xi$ to those satisfying
$\xi^{p}\geq R^{2}.$ (13.14)
Then, since $p\geq 2$,
$z_{m+1}\leq\big(\frac{C_{4}}{\xi^{p\alpha}}\big)^{\theta}2^{mp\theta(1+\alpha)}z_{m}^{1+\alpha}.$
Therefore, setting $C_{0}=2^{p\theta(1+\alpha)}$ and $N=(\frac{C_{4}}{\xi^{p\alpha}})^{\theta}$, we have the first inequality in Lemma 12.1, i.e. $z_{m+1}\leq NC_{0}^{i}z^{1+\alpha}_{m}$. To apply this lemma, we need to verify the second inequality there, i.e.
$z_{0}\leq N^{-\frac{1}{\alpha}}C_{0}^{-\frac{1}{\alpha^{2}}},$
where, recall, $z_{0}=\frac{\langle u^{p\theta}\mathbf{1}_{B_{R}\cap\{u>0\}}\rangle}{|B_{R}|}\big(\frac{|B_{R}\cap\{u>0\}|}{|B_{R}|}\big)^{\alpha}.$ The previous inequality holds, by the definition of $N$, if we select $\xi$ satisfying
$\xi\geq 2^{\frac{1+\alpha}{\alpha^{2}}}C_{4}^{\frac{1}{p\alpha}}z_{0}^{\frac{1}{p\theta}}$ (13.15)
We combine (13.14) and (13.15) by taking $\xi=2^{\frac{1+\alpha}{\alpha^{2}}}C_{4}^{\frac{1}{p\alpha}}z_{0}^{\frac{1}{p\theta}}+R^{\frac{2}{p}}$. Now Lemma 12.1 yields $z(\xi,\frac{R}{2})=0$, i.e. $\sup_{B_{\frac{R}{2}}}u\leq\xi$. Hence, $\sup_{B_{\frac{R}{2}}}u\leq Cz_{0}^{\frac{1}{p\theta}}+R^{\frac{2}{p}},$ as claimed. The proof of Lemma 13.2 is completed. ∎
Set
$\text{osc}(u,R):=\sup_{y,y^{\prime}\in B_{R}}|u(y)-u(y^{\prime})|.$
Lemma 13.3.
Fix $k_{0}$ by $2k_{0}=M(2R)+m(2R):=\sup_{B_{2R}}u+\inf_{B_{2R}}u.$ Assume that $|B_{R}\cap\{u>k_{0}\}|\leq\gamma|B_{R}|$ for some $\gamma<1$. If
${\rm osc}(u,2R)\geq 2^{n+1}CR^{\frac{2}{p}},$ (13.16)
then, for $k_{n}:=M(2R)-2^{-n-1}{\rm osc}(u,2R)$,
$\frac{|B_{R}\cap\{u>k_{n}\}|}{|B_{R}|}\leq cn^{-\frac{d}{2(d-1)}}.$
Proof of Lemma 13.3.
Let $h\in(k_{0},k)$. Define
$w:=\begin{cases}(u-h)^{\frac{p}{2}}&\text{if }h<u<k,\\ (k-h)^{\frac{p}{2}}&\text{if }u\geq k,\\ 0&\text{if }u\leq h.\end{cases}$
Note that $w=0$ in $B_{R}\setminus(B_{R}\cap\{u>k_{0}\})$. The measure of this set is greater than $\gamma|B_{R}|$, so the Sobolev embedding theorem yields
$(k-h)^{\frac{p}{2}}\left|B_{R}\cap\{u>k\}\right|^{\frac{d-1}{d}}$ $\leq c_{1}\langle w^{\frac{d}{d-1}}\mathbf{1}_{B_{R}}\rangle^{\frac{d-1}{d}}$
$\leq c_{2}\langle|\nabla w|\mathbf{1}_{\Delta}\rangle$
$\leq c_{2}|\Delta|^{\frac{1}{2}}\langle|\nabla(u-h)^{\frac{p}{2}}|^{2}\mathbf{1}_{B_{R}\cap\{u>h\}}\rangle^{\frac{1}{2}},$
where $\Delta:=B_{R}\cap\{u>h\}\setminus(B_{R}\cap\{u>k\}).$ On the other hand, repeating the proof of Lemma 13.1, but estimating the term $I_{3}$ there via (13.7) rather than going all the way to (13), we obtain
$\langle|\nabla(u-h)^{\frac{p}{2}}|^{2}\mathbf{1}_{B_{R}\cap\{u>h\}}\rangle$ $\leq\frac{C_{1}}{R^{2}}\langle(u-h)^{p}\mathbf{1}_{B_{2R}\cap\{u>h\}}\rangle$
$+\frac{C_{2}}{R^{2}}\langle|Q-(Q)_{R}|^{2}(u-h)^{p}\mathbf{1}_{B_{2R}\cap\{u>h\}}\rangle$
$+C_{3}\langle|f-\mu u|^{p}\mathbf{1}_{B_{2R}\cap\{u>h\}}\rangle$
Applying the John-Nirenberg inequality (13.2) in the second term, we get
$\langle|\nabla(u-h)^{\frac{p}{2}}|^{2}\mathbf{1}_{B_{R}\cap\{u>h\}}\rangle$ $\leq C_{4}R^{d-2}(M(2R)-h)^{p}+\frac{C_{5}}{R^{2}}(M(2R)-h)^{p}\cdot R^{d}\|Q\|^{2}_{{\rm BMO}}$
$+C_{6}\|f-\mu u\|^{p}_{\infty}R^{d}$
$\leq CR^{d-2}(M(2R)-h)^{p}+CR^{d},$
(we have used $\|f-\mu u\|_{\infty}\leq 2\|f\|_{\infty}$). For $h\leq k_{n}$, we have $M(2R)-h\geq M(2R)-k_{n}=2^{-n-1}{\rm osc}(u,2R)\geq CR^{\frac{2}{p}}$, where we have used (13.16), in which case
$(k-h)^{\frac{p}{2}}|B_{R}\cap\{u>k\}|^{\frac{d-1}{d}}\leq C|\Delta|^{\frac{1}{2}}R^{\frac{d-2}{2}}(M(2R)-h)^{\frac{p}{2}}.$
Now, choosing increasing finite sequence $k=k_{i}:=M(2R)-2^{-i-1}{\rm osc}(u,2R)$ for $i\in\{1,2,\cdots,n\}$ and $h=k_{i-1}$. Then
$M(2R)-h=2^{-i}{\rm osc}(u,2R),\quad|k-h|=2^{-i-1}{\rm osc}(u,2R)$
so
$\left|B_{R}\cap\{u>k_{n}\}\right|^{\frac{2(d-1)}{d}}\leq\left|B_{R}\cap\{u>k_{i}\}\right|^{\frac{2(d-1)}{d}}\leq C|\Delta_{i}|R^{d-2},$
where $\Delta_{i}=B_{R}\cap\{u>k_{i}\}\setminus(B_{R}\cap\{u>k_{i-1}\})]$. Summing up over $i$, we obtain
$n\left|B_{R}\cap\{u>k_{n}\}\right|^{\frac{2(d-1)}{d}}\leq CR^{d-2}\left|B_{R}\cap\{u>k_{0}\}\right|\leq C^{\prime}R^{2(d-1)},$
and the claimed inequality follows. This ends the proof of Lemma 13.3. ∎
We are in position to end the proof of Proposition 8.2. Fix $k_{0}=\frac{1}{2}(M(2R)+m(2R))$. Without loss of generality, $|B_{R}\cap\{u>k_{0}\}|\leq\frac{1}{2}|B_{R}|$ (otherwise replace $u$ by u-u). Set $k_{n}=M(2R)-2^{-n-1}{\rm osc}(u,2R)>k_{0}$. By Lemma 13.2 applied to $u-k_{n}$ ($k_{n}$ adds constant term $\mu k_{n}$ in the equation, but since $\{k_{n}\}$ is bounded, the constant in Lemma 13.2 can be chosen to be independent of $n$), we have
$\sup_{B_{\frac{R}{2}}}(u-k_{n})$ $\leq C_{1}\bigg(\frac{1}{|B_{R}|}\langle(u-k_{n})^{p\theta}\mathbf{1}_{B_{R}\cap\{u>k_{n}\}}\bigg)^{\frac{1}{p\theta}}\bigg(\frac{|B_{R}\cap\{u>k_{n}\}|}{|B_{R}|}\bigg)^{\frac{\alpha}{p\theta}}+R^{\frac{2}{p}}$
$\leq C_{1}\sup_{B_{R}}(u-k_{n})\bigg(\frac{|B_{R}\cap\{u>k_{n}\}|}{|B_{R}|}\bigg)^{\frac{1+\alpha}{p\theta}}+R^{\frac{2}{p}}.$ (13.17)
Fix $n$ by $cn^{-\frac{d}{2(d-1)}}\leq(\frac{1}{2C_{1}})^{\frac{p\theta}{1+\alpha}}$. We consider two cases. First, let ${\rm osc}(u,2R)\geq 2^{n+1}R^{\frac{2}{p}}$. Then by Lemma 13.3 (with, say, $C=1$) applied in the RHS of (13),
$M(R/2)-k_{n}$ $\leq\frac{1}{2}(M(2R)-k_{n})+R^{\frac{2}{p}}$
so
$M(R/2)\leq M(2R)-\frac{1}{2^{n+1}}{\rm osc}(u,2R)+\frac{1}{2}\frac{1}{2^{n+1}}{\rm osc}(u,2R)+R^{\frac{2}{p}},$
which yields
$M(R/2)-m(R/2)$ $\leq M(2R)-m(2R)-\frac{1}{2}\frac{1}{2^{n+1}}{\rm osc}(u,2R)+R^{\frac{2}{p}}$
$=\left(1-\frac{1}{2^{n+2}}\right){\rm osc}(u,2R)+R^{\frac{2}{p}}.$
Next, if ${\rm osc}(u,2R)\leq 2^{n+1}R^{\frac{2}{p}}$, then
${\rm osc}(u,R/2)\leq(1-\frac{1}{2^{n+2}}){\rm osc}(u,2R)+\frac{1}{2}R^{\frac{2}{p}}.$
This result provides the desired Hölder continuity of $u$ by applying the next Lemma 13.4 with $\tau=\frac{1}{4}$, $\delta=\log_{\tau}(1-2^{-n-1})$, and $0<\beta<\frac{2-p}{p}\wedge\delta$. Note that the second inequality in Lemma 13.3 is satisfied when $q=1$ and $\varphi$ is non-decreasing, which is our situation. ∎
Lemma 13.4 ([G, Lemma 7.3]).
Let $\varphi(t)$ be a positive function, and assume that there exists a constant $q$ and a number $0<\tau<1$ such that for every $0<R<R_{0}$,
$\varphi(\tau R)\leq\tau^{\delta}\varphi(R)+BR^{\beta}$
with $0<\beta<\delta$, and
$\varphi(t)\leq q\varphi(\tau^{k}R)$
for every $t$ in the interval $(\tau^{k+1}R,\tau^{k}R)$. Then, for every $0<\rho<R<R_{0}$, we have
$\varphi(\rho)\leq C\left(\frac{\rho}{R}\right)^{\beta}\varphi(R)+B\rho^{\beta}$
with a constant $C$ that depends only on $q$, $\tau$, $\delta$, and $\beta$.
14. Proof of Proposition 8.4 (Separation property)
Step 1. Without loss of generality, $x=0$. We will need the following local estimate on solution $u=u_{n,m}$ of (8.3): for all $\mu\geq\mu_{0}>0$,
$\sup_{B_{\frac{1}{2}}}|u|\leq K\biggl(\langle|u|^{p\theta}\mathbf{1}_{B_{1}}\rangle^{\frac{1}{p\theta}}+\big\langle|f|^{p\theta^{\prime}}\mathbf{1}_{B_{1}}\big\rangle^{\frac{1}{p\theta^{\prime}}}\biggr).$ (14.1)
In fact, it suffices to prove the previous estimate for $\sup_{B_{\frac{1}{2}}}u_{+}$ in the LHS.
To that end, we first establish the following Caccioppoli’s inequality for $v:=(u-k)_{+}$, $k\geq 0$:
$\|v\|^{p}_{L^{\frac{pd}{d-2}}(B_{r})}\leq C\biggl[(R-r)^{-2}|B_{R}|^{\frac{1}{\theta^{\prime}}}\|v\|^{p}_{L^{p\theta}(B_{R})}+\|f\mathbf{1}_{u>k}\|_{L^{p}(B_{R})}^{p}\biggr].$ (14.2)
To prove (14.2), we argue as in the proof of Lemma 13.1, but treat the term $\mu u$ differently: since $\mu$ and $k$ are non-negative, we have
$\mu(u-k)-\Delta(u-k)+(b+q)\cdot\nabla(u-k)\leq f.$
Therefore, multiplying the previous inequality by $v^{p-1}\eta$, with the cutoff function $\eta$ defined by (13.4), (13.5), and repeating the proof of Lemma 13.1, we obtain, for all $0<r<R\leq 1$,
$\|v^{\frac{p}{2}}\|^{2}_{W^{1,2}(B_{r})}\leq C_{1}\bigg[\frac{1}{(R-r)^{2}}|B_{R}|^{\frac{1}{\theta^{\prime}}}\|v^{\frac{p}{2}}\|_{L^{2\theta}(B_{R})}+\|f\mathbf{1}_{u>k}\|_{L^{p}(B_{R})}^{p}\bigg].$
The Sobolev embedding theorem now yields (14.2).
Set
$R_{m}:=\frac{1}{2}+\frac{1}{2^{m+1}},\quad m\geq 0,$
so $B^{m}:=B_{R_{m}}$ is a decreasing sequence of balls converging to the ball of radius $\frac{1}{2}$. For the purposes of this proof, we can estimate $|B_{R}|^{\frac{1}{\theta^{\prime}}}\leq 1$, which will make the iterations below converge slower, but will not change the sought estimate (14.1). Estimate (14.2) gives us
$\|v\|^{p}_{L^{\frac{pd}{d-2}}(B_{r})}$ $\leq C_{1}2^{2m}\|v\|_{L^{p\theta}(B^{m})}^{p}+C_{2}\|f\mathbf{1}_{u>k}\|_{L^{p}(B^{m})}^{p}$
$\leq C_{1}2^{2m}\|v\|_{L^{p\theta}(B^{m})}^{p}+C_{2}H|B^{m}\cap\{v>0\}|^{\frac{1}{\theta}},$ (14.3)
where $H:=\langle|f|^{p\theta^{\prime}}\mathbf{1}_{B^{0}}\rangle^{\frac{1}{\theta^{\prime}}}$ ($B^{0}=B_{1}$ is the ball of radius 1). On the other hand, by Hölder’s inequality,
$\|v\|_{L^{p\theta}(B^{m+1})}^{p\theta}\leq\|v\|_{L^{\frac{pd}{d-2}}(B^{m+1})}^{p\theta}\biggl(|B^{m}\cap\{v>0\}|\biggr)^{1-\frac{(d-2)\theta}{d}}.$
Applying (14.3) in the first multiple in the RHS, we obtain
$\|v\|_{L^{p\theta}(B^{m+1})}^{p\theta}\leq C\biggl(2^{2\theta m}\|v\|_{L^{p\theta}(B^{m})}^{p\theta}+H^{\theta}|B^{m}\cap\{v>0\}|\biggr)\biggl(|B^{m}\cap\{v>0\}|\biggr)^{1-\frac{(d-2)\theta}{d}}.$
Put
$v_{m}:=(u-k_{m})_{+},\quad k_{m}:=\xi(1-2^{-m})\uparrow\xi,$
where constant $\xi>0$ will be chosen later. Using $2^{2\theta m}\leq 2^{p\theta m}$ and dividing by $\xi^{p\theta}$, we obtain
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{L^{p\theta}(B^{m+1})}^{p\theta}$
$\leq C\biggl(\frac{2^{p\theta m}}{\xi^{p\theta}}\|v_{m+1}\|_{L^{p\theta}(B^{m})}^{p\theta}+\frac{1}{\xi^{p\theta}}H^{\theta}|B^{m}\cap\{u>k_{m+1}\}|\biggr)\bigl(|B^{m}\cap\{u>k_{m+1}\}|\bigr)^{1-\frac{(d-2)\theta}{d}}.$
From now on, we require that constant $\xi$ satisfies $\xi^{p}\geq H$, so
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{L^{p\theta}(B^{m+1})}^{p\theta}$ (14.4)
$\leq C\biggl(\frac{2^{p\theta m}}{\xi^{p\theta}}\|v_{m+1}\|_{L^{p\theta}(B^{m})}^{p\theta}+|B^{m}\cap\{u>k_{m+1}\}|\biggr)\bigl(|B^{m}\cap\{u>k_{m+1}\}|\bigr)^{1-\frac{(d-2)\theta}{d}}.$
Now,
$|B^{m}\cap\{u>k_{m+1}\}|$ $=\big|B^{m}\cap\bigg\{\bigg(\frac{u-k_{m}}{k_{m+1}-k_{m}}\bigg)^{2\theta}>1\bigg\}\big|$
$\leq(k_{m+1}-k_{m})^{-p\theta}\langle v^{p\theta}_{m}\mathbf{1}_{B^{m}}\rangle=\xi^{-p\theta}2^{p\theta(m+1)}\|v_{m}\|_{L^{p\theta}(B^{m})}^{p\theta},$
so using $\|v_{m+1}\|_{L^{p\theta}(B^{m})}\leq\|v_{m}\|_{L^{p\theta}(B^{m})}$ in (14.4) and applying the previous inequality, we obtain
$\frac{1}{\xi^{p\theta}}\|v_{m+1}\|_{L^{p\theta}(B^{m+1})}^{p\theta}\leq C2^{p\theta m(2-\frac{\theta}{\theta_{0}})}\biggl(\frac{1}{\xi^{p\theta}}\|v_{m}\|_{L^{p\theta}(B^{m})}^{p\theta}\biggr)^{2-\frac{(d-2)\theta}{d}}.$
Denote $z_{m}:=\frac{1}{\xi^{p\theta}}\|v_{m}\|_{L^{p\theta}(B^{m})}^{p\theta}.$ Then
$z_{m+1}\leq C\gamma^{m}z_{m}^{1+\alpha},\quad m=0,1,2,\dots,\quad\alpha:=1-\frac{(d-2)\theta}{d},\;\;\gamma:=2^{p\theta(2-\frac{(d-2)\theta}{d})}$
and $z_{0}=\frac{1}{\xi^{p\theta}}\langle u_{+}^{p\theta}\mathbf{1}_{B^{0}}\rangle\leq C^{-\frac{1}{\alpha}}\gamma^{-\frac{1}{\alpha^{2}}}$ (recall: $B^{0}:=B_{R_{0}}\equiv B_{1}$) provided that we fix $c$ by
$\xi^{p\theta}:=C^{\frac{1}{\alpha}}\gamma^{\frac{1}{\alpha^{2}}}\langle u_{+}^{p\theta}\mathbf{1}_{B^{0}}\rangle+H^{\theta}.$
Hence, by Lemma 12.1, $z_{m}\rightarrow 0$ as $m\rightarrow\infty$. It follows that $\sup_{B_{1/2}}u_{+}\leq\xi,$ and the claimed inequality follows.
Step 2. Next, we bound $\langle|u|^{p\theta}\mathbf{1}_{B_{1}}\rangle^{\frac{1}{p\theta}}$ using Lemma 11.1 with $s=p\theta$, which allows us to conclude that if $\sigma$ (in the definition of weight $\rho$) is fixed sufficiently small, then for all $x\in\mathbb{R}^{d}$
$\sup_{B_{\frac{1}{2}}(x)}u_{+}\leq K\biggl(\langle|f|^{p\theta}\rho_{x}\rangle^{\frac{1}{p\theta}}+\big\langle|f|^{p\theta^{\prime}}\mathbf{1}_{B_{1}(x)}\big\rangle^{\frac{1}{p\theta^{\prime}}}\biggr).$
This ends the proof of Proposition 8.4. ∎
15. Proof of Proposition 8.3 (Convergence)
The proof given below is close to [KS3, Proof of Theorem 4.3]. This argument (in $L^{2}$) can, in principle, be replaced by an argument based on the Lions’ variational approach that handles well both $b\in\mathbf{F}_{\delta}$ and $q\in\mathbf{BMO}^{-1}$ (regarding the latter, see [QX]).
In order to prove the existence of the limit, we need to construct first intermediate semigroups $e^{-t\Lambda(b_{n},q)}$ in $L^{p}$. Here $b_{n}$ are bounded and smooth, but $q$ can be singular.
Step 1. At the first step, we construct $e^{-t\Lambda(0,q)}$ in $L^{2}$. Here we work over complex numbers (in this regard, see Remark 15.1). Define sesquilinear form
$\tau[v,w]:=\langle\nabla v,\nabla w\rangle+\langle q\cdot\nabla v,w\rangle,\quad D(\tau)=W^{1,2},$
where $\langle v,w\rangle=\langle v\bar{w}\rangle$.
a) $\tau$ is bounded: $|\tau[v,w]|\leq C\|\nabla v\|_{2}\|\nabla w\|_{2}.$ Indeed, by the compensated compactness estimate (Proposition 2.1), $|\langle q\cdot\nabla v,w\rangle|\leq C^{\prime}\|\nabla v\|_{2}\|\nabla w|_{2}$.
b) $\tau$ is a sectorial form:
${\rm Im}\,\tau[v,v]\leq K{\rm Re}\,\tau[v,v],\quad v\in D(\tau)$
for some constant $K>0$. Indeed, writing $v=r+ie$, where $r,e$ are real-valued elements of $W^{1,2}$, we have
$\tau[v,v]$ $=\langle|\nabla r|^{2}+|\nabla e|^{2}\rangle+\langle q\cdot\big[(\nabla r)r+(\nabla e)e\big]\rangle+i\langle q\cdot[(\nabla e)r-(\nabla r)e]\rangle,$
so, taking into that the second term vanishes due to the anti-symmetry of $Q$, we have
${\rm Re}\,\tau[v,v]=\langle|\nabla r|^{2}+|\nabla e|^{2}\rangle,\quad{\rm Im}\,\tau[v,v]=\langle q\cdot[(\nabla e)r-(\nabla r)e]\rangle.$
Now, invoking again the compensated compactness, we obtain ${\rm Im}\,\tau[v,v]\leq C\|Q\|_{\rm BMO}{\rm Re}\,\tau[u,u]$.
c) ${\rm Re}\,\tau[v,v]$ is a closed form, i.e. if $v_{k}\rightarrow v$ in $L^{2}$ and ${\rm Re}\,\tau[v_{k}-v_{l}]\rightarrow 0$ as $k,l\rightarrow\infty$, then $\tau[v_{k}-v]\rightarrow 0$ (we only need to look at the real part of $\tau$ due to its sectoriality). But in our case the latter is just a re-statement that $W^{1,2}$ is a complete space.
Therefore, there exists a unique ($m$-sectorial) operator $\Lambda(0,q)$ such that
$\langle\Lambda(0,q)v,w\rangle=\tau[v,w],\quad v\in D(\Lambda(0,q))\subset W^{1,2},\quad w\in D(\tau)=W^{1,2},$
see [Ka, Ch. VI, §2]. This operator, being $m$-sectorial, generates a holomorphic semigroup $e^{-t\Lambda(0,q)}$ in $L^{2}$.
Remark 15.1.
A property of holomorphic semigroups that we will need at Step 5 is as follows: for every $f\in L^{2}$, $t>0$, $e^{-t\Lambda(0,q)}f$ belongs to the domain $D(\Lambda(0,q))$.
Step 2. We need to construct $e^{-t\Lambda(b_{n},q)}$, i.e. to add the drift term $b_{n}\cdot\nabla$. So, we re-do what we did above for the sesquilinear form
$\tau[v,w]:=\langle\nabla v,\nabla w\rangle+\langle q\cdot\nabla v,w\rangle+\langle b_{n}\cdot\nabla v,w\rangle.$ (15.1)
In particular, applying Cauchy-Schwarz’ inequality to $\langle b_{n}\cdot\nabla v,w\rangle$, we obtain
${\rm Im}\,\tau[v,v]\leq K\big({\rm Re}\,\tau[v,v]+C\langle v,v\rangle\big),$ (15.2)
where $C=C(\|b_{n}\|_{\infty})\geq 0$. The cited results apply to the case (15.2), and we get ($m$-sectorial) generator $\Lambda(b_{n},q)$ of a holomorphic semigroup in $L^{2}$ such that
[…]
where $v\in D(\Lambda(b_{n},q))\subset W^{1,2}$, $w\in D(\tau)=W^{1,2}$. (Or we could appeal to the Hille perturbation theorem and work with the algebraic sum $\Lambda(b_{n},q):=\Lambda(0,q)+b_{n}\cdot\nabla$, $D(\Lambda(b_{n},q))=D(\Lambda(0,q))$, that still generates a holomorphic semigroup in $L^{2}$.)
Step 3. Let us now note that we have the following quasi-contraction estimate for $e^{-t\Lambda(b_{n},q_{m})}$, $n,m=1,2,\dots$, in $L^{p}$, $p>\frac{2}{2-\sqrt{\delta}}$, $p\geq 2$:
$\|e^{-t\Lambda(b_{n},q_{m})}f\|_{p}\leq e^{\omega t}\|f\|_{p},\quad t\geq 0,\quad f\in C_{c}^{\infty}.$ (15.4)
for some $\omega$ independent of $n,m$. Indeed, setting $u_{n,m}=e^{-t(\omega+\Lambda(b_{n},q_{m}))}f$, we multiply the corresponding parabolic equation $(\omega+\partial_{t}-\Delta+(b_{n}+q_{m})\cdot\nabla)u_{n,m}=0$ by $u_{n,m}|u_{n,m}|^{p-2}$ and repeat the proof of Lemma 11.1 (without the weight there) with the obvious modifications for the time derivative term (see Step 5 below for details). The constant $\omega>0$ needs to be chosen to account e.g. for the constant $c_{\delta}$ resulting from the use of condition $b\in\mathbf{F}_{\delta}$.
If we do include the weight $\rho_{x}$, then, again arguing as in the proof of Lemma 11.1, we obtain a weighted quasi contraction estimate
$\|e^{-t\Lambda(b_{n},q_{m})}f\|_{L_{\rho_{x}}^{p}}\leq e^{\omega^{\prime}t}\|f\|_{L_{\rho_{x}}^{p}},\quad t\geq 0,$ (15.5)
for any $x\in\mathbb{R}^{d}$, with $\omega^{\prime}$ independent of $n$, $m$ or $x$. (Since $\rho_{x}\leq 1$, the result of the next section implies that
$u_{m,n}=(\mu+\Lambda(b_{n},q_{m}))^{-1}f\rightarrow u_{n}=(\mu+\Lambda(b_{n},q))^{-1}f\quad\text{ in $L_{\rho_{x}}^{p}$}.$ (15.6)
Moreover, (15.5) ensures that the resulting semigroup $e^{-t\Lambda(b_{n},q)}$ is strongly continuous in $L^{p}_{\rho_{x}}$.)
Step 4. From now on, we work over reals. Let us show convergence
$e^{-t\Lambda(b_{n},q_{m})}\rightarrow e^{-t\Lambda(b_{n},q)}\text{ in $L^{2}$ loc.\,uniformly in $t\geq 0$}$
as $m\rightarrow\infty$. Here $n$ is fixed.
Since $n$ is fixed, it is easily seen that the operator norms of the resolvents $\|(\mu+\Lambda(b_{n},q_{m}))^{-1}\|_{2\rightarrow 2}$ are uniformly in $m$ bounded, provided $\mu=\mu(\|b_{n}\|_{\infty})$ is fixed sufficiently large. It suffices for us (see [Ka, Ch. XI, §5]) to show the convergence of the resolvents
$(\mu+\Lambda(b_{n},q_{m}))^{-1}f\rightarrow(\mu+\Lambda(b_{n},q))^{-1}f\quad\text{ in $L^{2}$ as $m\rightarrow\infty$}$
for all $f\in L_{c}^{2}$ (subscript $c$ means compact support).
The standard argument yields that $u_{n}=(\mu+\Lambda(b_{n},q))^{-1}f$ is the unique weak solution to the elliptic equation $(\mu-\Delta+(b_{n}+q)\cdot\nabla)u_{n}=f,$ where the former means that
$\mu\langle u_{n},\varphi\rangle+\langle\nabla u_{n},\nabla\varphi\rangle+\langle(b_{n}+q)\cdot\nabla u_{n},\varphi\rangle=\langle f,\varphi\rangle,\quad\varphi\in C_{c}^{\infty}.$
(The compensated compactness estimate $|\langle q\cdot\nabla u_{n},\varphi\rangle|=|\langle Q\cdot\nabla u_{n},\nabla\varphi\rangle|\leq C^{\prime}\|\nabla u_{n}\|_{2}\|\nabla\varphi\|_{2}$ of Proposition 2.1 allows us to pass to test functions $\varphi\in W^{1,2}$.) In turn, this uniqueness and the usual weak compactness argument shows that $u_{n,m}=(\mu+\Lambda(b_{n},q_{m}))^{-1}f$, i.e. solutions to the approximating elliptic equations $(\mu-\Delta+(b_{n}+q_{m})\cdot\nabla)u_{n,m}=f,$ converge weakly in $W^{1,2}$ to the same limit $u_{n}$. Now we can appeal to the Rellich-Kondrashov theorem to obtain $u_{n,m}\rightarrow u_{n}$ in $L^{2}_{{\rm loc}}$, and further to upgrade this convergence to $u_{n,m}\rightarrow u_{n}$ in $L^{2}$ by “cutting tails” of $u_{n,m}$ at infinity uniformly in $m$ using the upper Gaussian bound on the heat kernel of $\Delta+(b_{n}+q_{m})\cdot\nabla$, see [QX], and taking into account that $f$ has compact support. (The constants in the upper Gaussian heat kernel bound, which yields the bound on the integral kernel of the resolvents, will depend on $n$. Since each $b_{n}$ is bounded, adding the drift term in $b_{n}\cdot\nabla$ does not affect the proof of the upper bound in [QX] which employs Moser’s iterations and the Davies device. In fact, one can account for $b_{n}\cdot\nabla$ by once again introducing a constant term in the operator that will absorb the contribution from $b_{n}\cdot\nabla$ in Moser’s method.)
Now, since $e^{-t\Lambda(b_{n},q_{m})}$ are $L^{\infty}$ contractions, we obtain by interpolation that
$e^{-t\Lambda(b_{n},q_{m})}\rightarrow e^{-t\Lambda(b_{n},q)}\text{ in $L^{p}$ loc.\,uniformly in $t\geq 0$}$
for all $p\geq 2$. This implies the convergence of the resolvents: as $m\rightarrow\infty$,
$u_{n,m}=(\mu+\Lambda(b_{n},q_{m}))^{-1}f\rightarrow u_{n}=(\mu+\Lambda(b_{n},q))^{-1}f\quad\text{ in $L^{p}$}$ (15.7)
with $\mu$ independent of $n$, $m$.
Step 5. Having constructed the intermediate semigroups $e^{-t\Lambda(b_{n},q)}$, our goal now is to show that they converge as $n\rightarrow\infty$. For reader’s convenience, at this step we give a proof that assumes additionally that $b_{n}\rightarrow b$ in $L^{2}$. In this case we obtain that, every $f\in C_{c}^{\infty}$,
$\{v_{n}(t):=e^{-t(\omega+\Lambda(b_{n},q))}f\}_{n=1}^{\infty}\quad\text{ is a Cauchy sequence in $L^{\infty}([0,1],L^{p})$}$ (15.8)
for
$p>\frac{2}{2-\sqrt{\delta}},\quad p\geq 2,\text{ for some fixed $\omega$}.$
Taking into account Remark 15.1 about $v(t)\in D(\Lambda(b_{n},q))$ for all $t>0$ and the identity (15.3), we can write
$\langle\partial_{t}v_{n},\psi\rangle+\omega\langle v_{n},\psi\rangle+\langle\nabla v_{n},\nabla\psi\rangle+\langle b_{n}\cdot\nabla v_{n},\psi\rangle-\langle Q\cdot\nabla v_{n},\nabla\psi\rangle=0,$ (15.9)
for all $\psi(t,\cdot)\in W^{1,2}$, where, recall, $\nabla Q=q$. Set
$h:=v_{n_{1}}-v_{n_{2}}.$
Subtracting identities (15.9) for $v_{n_{1}}$ and $v_{n_{2}}$ from each other, we obtain
$\langle\partial_{t}h,\psi\rangle+\omega\langle h,\psi\rangle+\langle\nabla h,\nabla\psi\rangle+\langle b_{n_{1}}\cdot\nabla h,\psi\rangle-\langle Q\cdot\nabla h,\nabla\psi\rangle=-(b_{n_{1}}-b_{n_{2}})\cdot\nabla v_{n_{2}}.$
We are basically in the setting of Proposition 11.2(i) with the only difference that $Q$ is no longer smooth and we are dealing with weak solutions of the parabolic equations rather than classical solutions. However, the latter does not pose a difficulty: since $p\geq 2$ the standard result on the composition of Lipschitz functions with the elements of Sobolev spaces yield that $h(t)|^{p-2}, h(t)|h(t)|$$h(t)|^{\frac{p}{2}-1}$, $|h(t)|^{\frac{p}{2}}\in W^{1,2}$. Therefore, we can take $\psi=h|h|^{p-2}$, obtaining
$\frac{1}{p}\partial_{t}\langle|h|^{p}\rangle+\omega\langle|h|^{p}\rangle+\frac{4(p-1)}{p^{2}}\langle|\nabla|h|^{\frac{p}{2}}|^{2}\rangle$ $+\frac{2}{p}\langle\langle b_{n_{1}}\cdot\nabla|h|^{\frac{p}{2}},|h|^{\frac{p}{2}}\rangle$
$\leq|\langle(b_{n_{1}}-b_{n_{2}})\cdot\nabla v_{n_{2}},h|h|^{p-2}\rangle|,$ (15.10)
where we have used anti-symmetry of $Q$ and $h(0)=0$. We handle the term containing $b_{n_{1}}$ as in the proof of Proposition 11.1, i.e. first applying quadratic inequality and then $b_{n_{1}}\in\mathbf{F}_{\delta}$. Now, handling the time derivative term as in the proof of Proposition 11.2, we obtain
$\frac{1}{p}\sup_{s\in[0,t]}\langle|h(s)|^{p}\rangle+\biggl[\omega-2\frac{c_{\delta}}{\sqrt{\delta}}\biggr]\int_{0}^{t}\|h\|_{p}^{p}ds+$ $\biggl[\frac{4(p-1)}{p^{2}}-\frac{2}{p}\sqrt{\delta}\biggr]\int_{0}^{t}\langle|\nabla|h|^{\frac{p}{2}}|^{2}\rangle ds$
$\leq 2\int_{0}^{t}\|b_{n_{1}}-b_{n_{2}}\|_{2}\|\nabla v_{n_{2}}\|_{2}\|h\|_{\infty}^{p-1}ds.$ (15.11)
Take $\frac{1}{2}\omega:=2\frac{c_{\delta}}{\sqrt{\delta}}$. Since $p>\frac{2}{2-\sqrt{\delta}}$, the expressions in square brackets are strictly positive. In the right-hand side of (15.11), $\|b_{n_{1}}-b_{n_{2}}\|_{2}\rightarrow 0$ as $n_{1},n_{2}\rightarrow\infty$ and $\|h(s)\|_{\infty}^{p-1}\leq 2^{p-1}\|f\|^{p-1}_{\infty}$, $s\in[0,t]$, for all $n_{1}$, $n_{2}$. It remains to note that $\int_{0}^{t}\|\nabla v_{n_{2}}\|_{2}ds$ is uniformly in $n_{2}$ bounded. Indeed, by (15.9) with $n=n_{2}$ and $\psi=v_{n_{2}}$ upon noting that
$|\langle b_{n_{2}}\cdot\nabla v_{n_{2}},v_{n_{2}}\rangle|+|\langle Q\cdot\nabla v_{n_{2}},\nabla v_{n_{2}}\rangle|$
$=|\langle b_{n_{2}}\cdot\nabla v_{n_{2}},v_{n_{2}}\rangle|\leq\frac{1}{2}\|b_{n_{2}}\|^{2}_{2}\|v_{n_{2}}\|^{2}_{\infty}+\frac{1}{2}\|\nabla v_{n_{2}}\|_{2}^{2},$
where we have used again the anti-symmetry of $Q$. Now we use $\sup_{n_{2}}\|b_{n_{2}}\|^{2}<\infty$ and the obvious a priori estimate $\|v_{n_{2}}(t)\|_{\infty}\leq\|f\|_{\infty}$, $t\geq 0$. Thus, the right-hand side of (15.11) tends to 0 as $n_{1}$, $n_{2}\rightarrow\infty$. Hence $h\rightarrow 0$ in $L^{\infty}([0,1],L^{p})$, and (15.8) follows. The latter and the contractivity estimate (15.4) yield that the limit $v=L^{p}\mbox{-}\lim_{n}v_{n}$ (loc. uniformly in $t\geq 0$) determines a strongly continuous semigroup in $L^{p}$, say, $v(t)=:e^{-t\Lambda(b,q)}f$. In turn, the convergence of the semigroups yields the convergence of the resolvents
$u_{n}=(\mu+\Lambda(b_{n},q))^{-1}f\rightarrow u=(\mu+\Lambda(b,q))^{-1}f\quad\text{ in $L^{p}$}$
with $\mu$ independent of $n$ (proportional to $\omega$). In view of (15.7), the existence of the limit $u=L^{p}\mbox{-}\lim_{n}u_{n}=L^{p}\mbox{-}\lim_{n}\lim_{m}u_{n,m}$ follows.
Step 6. Finally, we prove convergence in the general case, i.e. we do not assume global convergence of $b_{n}$ to $b$ in $L^{2}$. We show that, for any $x\in\mathbb{R}^{d}$, for every $f\in C_{c}^{\infty}$,
$\{v_{n}(t):=e^{-t(\mu+\Lambda(b_{n},q))}f\}_{n=1}^{\infty}\quad\text{ is a Cauchy sequence in $L^{\infty}([0,1],L_{\rho_{x}}^{p})$}$ (15.12)
for $p>\frac{2}{2-\sqrt{\delta}}$, $p\geq 2$, for some fixed $\omega$. Then, repeating the argument in the end of Step 5 but using (15.5) and (15.6), we will obtain the claimed in Proposition 8.3 existence of the limit
$u\;\biggl(=L^{p}_{\rho_{x}}\mbox{-}\lim_{n}u_{n}\biggr)\;=L^{p}_{\rho_{x}}\mbox{-}\lim_{n}\lim_{m}u_{n,m}.$
Let us prove (15.12). This time, taking $\psi=h|h|^{p-2}\rho_{x}$, we obtain
$\sup_{s\in[0,t]}\langle|h(s)|^{p}\rho_{x}\rangle+C_{1}\int_{0}^{t}\langle|h|^{p}\rho_{x}\rangle ds+$ $C_{2}\int_{0}^{t}\langle|\nabla|h|^{\frac{p}{2}}|^{2}\rho_{x}\rangle ds$
$\leq C_{3}\int_{0}^{t}|\langle(b_{n_{1}}-b_{n_{2}})\cdot\nabla v_{n_{2}},h|h|^{p-2}\rho_{x}\rangle|ds$ (15.13)
for constants $C_{1}$-$C_{3}$ independent of $n$, $m$, constant $C_{1}$ being strictly positive provided that $\sigma$ in the definition of $\rho$ is fixed sufficiently small. Arguing as at the previous step, we obtain
$\int_{0}^{1}\sup_{n_{2}}\langle|\nabla v_{n_{2}}(s)|^{2}\rho_{x}\rangle ds<\infty.$
Therefore,
$\int_{0}^{T}|\langle(b_{n_{1}}-b_{n_{2}})\cdot\nabla v_{n_{2}},h|h|^{p-2}\rho_{x}\rangle|ds$ $\leq\langle|b_{n_{1}}-b_{n_{2}}|^{2}\rho_{x}\rangle\|h\|_{\infty}^{p-1}\int_{0}^{1}\langle|\nabla v_{n_{2}}|^{2}\rho_{x}\rangle ds$
[…]
$\rightarrow 0\quad\text{ as $n_{1},n_{2}\rightarrow\infty$.}$
Combining this with (15.13), we obtain the claimed convergence (15.12). ∎
Remark 15.2.
If we try to obtain a stronger result about the existence of the limit $s\mbox{-}L^{p}\mbox{-}\lim_{n,m}u_{m,n}$ by extending the proof of Lemma 8.3 to the sequence $h:=u_{n_{1},m_{1}}-u_{n_{2},m_{2}}$, then we get an extra term in the right-hand side of (15.10): $|\langle(Q_{m_{1}}-Q_{m_{2}})\cdot\nabla u_{n_{2},m_{2}},\nabla(h|h|^{p-2})\rangle|$. It can be dealt with in two ways:
(a) We can estimate
$|\langle(Q_{m_{1}}-Q_{m_{2}})\cdot\nabla u_{n_{2},m_{2}},\nabla(h|h|^{p-2})\rangle|\leq\|Q_{m_{1}}-Q_{m_{2}}\|_{s}\|\nabla u_{n_{2},m_{2}}\|_{s^{\prime}}(p-1)(2\|f\|_{\infty})^{p-2}\|\nabla h\|_{2},$
$\frac{1}{s}+\frac{1}{s^{\prime}}=\frac{1}{2}.$
So, assuming for the illustration purposes that we have global convergence $\|Q_{m_{1}}-Q_{m_{2}}\|_{s}\rightarrow 0$ as $m_{1},m_{2}\rightarrow\infty$, we need a bound on $\|\nabla u_{n_{2},m_{2}}\|_{s^{\prime}}$ for a $s^{\prime}>2$. In principle, $s^{\prime}$ can be chosen to be close to $2$. To obtain such an estimate, we can use Gehring-Giaquinta-Modica’s lemma as in the proof of Theorem 6.3, but this, at least in the present form of the argument, requires us to consider the equation in $L^{2}$, hence we need to require $\delta<1$ rather than $\delta<4$ as in (5.3).
(b) Another option is to use the estimate
$|\langle(Q_{m_{1}}-Q_{m_{2}})\cdot\nabla u_{n_{2},m_{2}},\nabla(h|h|^{p-2})\rangle|\leq\|Q_{m_{1}}-Q_{m_{2}}\|_{{\rm BMO}}\|\nabla u_{n_{2},m_{2}}\|_{2}(2\|f\|_{\infty})^{p-2}\|\nabla h\|_{2},$
where $\|\nabla u_{n_{2},m_{2}}\|_{2}$ can be estimated as in the proof of Theorem 8.3, but now to have convergence $\|Q_{m_{1}}-Q_{m_{2}}\|_{{\rm BMO}}\rightarrow 0$ as $m_{1},m_{2}\rightarrow\infty$ we need a stronger hypothesis of the matrix field $Q$ and thus on $q$, namely, that $Q$ has entries in ${\rm VMO}(\mathbb{R}^{d})$.
16. Proof of Lemma 8.3
It suffices to carry out the proof for $b_{n}$ and $q_{m}$, and then use the convergence result of Proposition 8.3. Thus, our goal is to show that
$\sup_{n,m}\|\nabla u_{n,m}\|_{2+\varepsilon}<\infty$ (16.1)
for some $\varepsilon>0$ independent of $n,m$.
Put for brevity $b=b_{n}$, $q=q_{m}$ and $u=u_{n,m}$, so
$\big(\mu-\Delta+(b+q)\cdot\nabla\big)u=f.$
Let us fix some $1<\theta<\frac{d}{d-2}$.
By Lemma 13.1 (with $p=2$ there, which is admissible since $\delta<1$), the function $v:=(u-k)_{+}$ ($k\in\mathbb{R}$) satisfies Caccioppoli’s inequality: for all $x\in\mathbb{R}^{d}$, $0<r<R<\frac{1}{2}$,
$\|\nabla v\|_{L^{2}(B_{r}(x))}^{2}\leq\frac{K_{1}}{(R-r)^{2}}|B_{R}|^{\frac{1}{\theta^{\prime}}}(1+\|Q\|_{{\rm BMO}}^{2})\|v\|_{L^{2\theta}(B_{R})}^{2}+K_{2}\|(f-\mu u)\mathbf{1}_{v>0}\|_{L^{2}(B_{R})}^{2},$ (16.2)
for constants $K_{1}$, $K_{2}$ independent of $k$, $r$, $R$ and $n$, $m$. (There is some abuse of notation: our $R$ here is not the radius of a fixed large ball in Lemma 8.3, but this should not cause a confusion.) We will obtain the sought bound (16.1) by applying a corollary of this Caccioppoli’s inequality in the Gehring-Giaquinta-Modica lemma:
Lemma 16.1.
Assume that there exist constants $K\geq 1$, $1<\nu<\infty$ such that, for given $0\leq g\in L_{{\rm loc}}^{q}$, $0\leq h\in L_{{\rm loc}}^{\nu}\cap L^{\infty}$ we have, for all $x\in\mathbb{R}^{d}$,
$\biggl(\frac{1}{|B_{R}|}\langle g^{\nu}\mathbf{1}_{B_{R}(x)}\rangle\biggr)^{\frac{1}{\nu}}\leq\frac{K}{|B_{2R}|}\langle g\mathbf{1}_{B_{2R}(x)}\rangle+\biggl(\frac{1}{|B_{2R}|}\langle h^{\nu}\mathbf{1}_{B_{2R}(x)}\rangle\biggr)^{\frac{1}{\nu}}$
for all $0<R<\frac{1}{2}$. Then $g\in L_{{\rm loc}}^{s}$ for some $s>\nu$ and, for all $x\in\mathbb{R}^{d}$,
$\biggl(\frac{1}{|B_{R}|}\langle g^{s}\mathbf{1}_{B_{R}(x)}\rangle\biggr)^{\frac{1}{s}}\leq C_{1}\biggl(\frac{1}{|B_{2R}|}\langle g^{\nu}\mathbf{1}_{B_{2R}(x)}\rangle\biggr)^{\frac{1}{\nu}}+C_{2}\biggl(\frac{1}{|B_{2R}|}\langle h^{s}\mathbf{1}_{B_{2R}(x)}\rangle\biggr)^{\frac{1}{s}}.$
Remark 16.1.
The authors of [KrS] proved Lemma 16.1 with explicit constants independent of the dimension $d$.
We are in position to prove (16.1). Put, for brevity, $x=0$.
Step 1. Set $(u_{n})_{B_{2R}}:=\frac{1}{|B_{2R}|}\langle u_{n}\mathbf{1}_{B_{2R}}\rangle$. Applying (16.2) to the positive and the negative parts of $u_{n}-(u_{n})_{B_{2R}}$, we obtain
$\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{R}}\rangle\leq\frac{K_{1}}{|B_{2R}|^{\frac{2}{d}}}|B_{2R}|^{\frac{1}{\theta^{\prime}}}(1+\|Q\|_{{\rm BMO}}^{2})\langle|u_{n}-(u_{n})_{B_{2R}}|^{2\theta}\mathbf{1}_{B_{2R}}\rangle^{\frac{1}{\theta}}+K_{2}\langle|f-\mu u_{n}|^{2}\mathbf{1}_{B_{2R}}\rangle,\quad 0<R<\frac{1}{2}.$ (16.3)
By the Sobolev-Poincaré inequality,
$\biggl(\frac{1}{|B_{2R}|}\langle(u_{n}-(u_{n})_{B_{2R}})^{2\theta}\mathbf{1}_{B_{2R}}\rangle\biggr)^{\frac{1}{2\theta}}\leq C|B_{R}|^{\frac{1}{d}}\biggl(\frac{1}{|B_{2R}|}\langle|\nabla u_{n}|^{\frac{2\theta d}{d+2\theta}}\mathbf{1}_{B_{2R}}\rangle\biggr)^{\frac{d+2\theta}{2\theta d}},$ (16.4)
i.e.
$\langle(u_{n}-(u_{n})_{B_{2R}})^{2\theta}\mathbf{1}_{B_{2R}}\rangle^{\frac{1}{\theta}}\leq C^{2}|B_{R}|^{\frac{2}{d}+\frac{1}{\theta}}\biggl(\frac{1}{|B_{2R}|}\langle|\nabla u_{n}|^{\frac{2\theta d}{d+2\theta}}\mathbf{1}_{B_{2R}}\rangle\biggr)^{\frac{d+2\theta}{\theta d}}.$
Plug in above estimate in (16.4), and divide both side by $|B_{R}|$ then for appropriate constants $C_{1}$ and $c$ were $C_{1}$ depends on $\|Q\|_{{\rm BMO}}$
$\frac{1}{|B_{R}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{R}}\rangle\leq C_{1}\biggl(\frac{1}{|B_{2R}|}\langle|\nabla u_{n}|^{\frac{2\theta d}{d+2\theta}}\mathbf{1}_{B_{2R}}\rangle\biggr)^{\frac{d+2\theta}{\theta d}}+\frac{c}{|B_{2R}|}\langle|f-\mu u_{n}|^{2}\mathbf{1}_{B_{2R}}\rangle,$
Then the condition of the Gehring-Giaquinta-Modica lemma is verified with $g=|\nabla u_{n}|^{\frac{2\theta d}{d+2\theta}}$, $g^{\nu}=|\nabla u_{n}|^{2}$ (so $\nu=\frac{d+2\theta}{\theta d}$) and $h=(c^{\frac{1}{2}}|f-\mu u_{n}|)^{\frac{2\theta d}{d+2\theta}}$, $h^{\nu}=c|f-\mu u_{n}|^{2}$. Hence there exists $s>\frac{d+2\theta}{\theta d}$ such that
$\biggl(\frac{1}{|B_{R}|}\langle|\nabla u_{n}|^{s\frac{2\theta d}{d+2\theta}}\mathbf{1}_{B_{R}}\rangle\biggr)^{\frac{1}{s}}\leq C_{1}\biggl(\frac{1}{|B_{2R}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{2R}}\rangle\biggr)^{\frac{\theta d}{d+2\theta}}+C_{2}\biggl(\frac{1}{|B_{2R}|}\langle|f-\mu u_{n}|^{s\frac{2\theta d}{d+2\theta}}\mathbf{1}_{B_{2R}}\rangle\biggr)^{\frac{1}{s}},$
where all constants are independent of $n$, or
$\frac{1}{|B_{R}|}\langle|\nabla u_{n}|^{s\frac{2\theta d}{d+2\theta}}\mathbf{1}_{B_{R}}\rangle\leq C_{1}^{\prime}\biggl(\frac{1}{|B_{2R}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{2R}}\rangle\biggr)^{s\frac{\theta d}{d+2\theta}}+C_{2}^{\prime}\frac{1}{|B_{2R}|}\langle|f-\mu u_{n}|^{s\frac{2\theta d}{d+2\theta}}\mathbf{1}_{B_{2R}}\rangle.$
Fix some $R$, say, $R=1$. We consider equally spaced grid $\frac{1}{2}\mathbb{Z}^{d}$ in $\mathbb{R}^{d}$ so that the smaller balls centered at the nodes of the grid cover $\mathbb{R}^{d}$, apply the previous estimate on each ball, and then sum up. We obtain a global estimate
$\|\nabla u_{n}\|^{s\frac{2\theta d}{d+2\theta}}_{s\frac{2\theta d}{d+2\theta}}\leq C_{3}\sum_{x\in c\mathbb{Z}^{d}}\biggl(\frac{1}{|B_{2}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{2}(x)}\rangle\biggr)^{s\frac{\theta d}{d+2\theta}}+C_{4}\|f-\mu u_{n}\|^{s\frac{2\theta d}{d+2\theta}}_{s\frac{2\theta d}{d+2\theta}}.$ (16.5)
To deal with the first term in the right-hand side, we split the grid into two parts: $I:=\{x\in c\mathbb{Z}^{d}\mid\frac{1}{|B_{2}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{2}(x)}\rangle>1\}$ and its complement $I^{c}$. For the nodes in the complement we have, taking into account that $s\frac{\theta d}{d+2\theta}>1$,
$\sum_{x\in I^{c}}\biggl(\frac{1}{|B_{2}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{2}(x)}\rangle\biggr)^{s\frac{\theta d}{d+2\theta}}$ $\leq\sum_{x\in c\mathbb{Z}^{d}}\frac{1}{|B_{2}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{2}(x)}\rangle$
$\leq C_{5}\langle|\nabla u_{n}|^{2}\rangle.$
In turn, there are only finitely many nodes in $I$. In fact, the cardinality of $I$ can be estimated in terms of $\langle|\nabla u_{n}|^{2}\rangle$:
$|I|<\sum_{x\in I}\frac{1}{|B_{2}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{2}(x)}\rangle\leq C_{6}\langle|\nabla u_{n}|^{2}\rangle.$
So,
$\sum_{x\in I}\biggl(\frac{1}{|B_{2}|}\langle|\nabla u_{n}|^{2}\mathbf{1}_{B_{2}(x)}\rangle\biggr)^{s\frac{\theta d}{d+2\theta}}$ $\leq\sum_{x\in I}\biggl(\frac{1}{|B_{2}|}\langle|\nabla u_{n}|^{2}\rangle\biggr)^{s\frac{\theta d}{d+2\theta}}$
$\leq C_{7}\langle|\nabla u_{n}|^{2}\rangle^{1+s\frac{\theta d}{d+2\theta}}.$
We arrive at a global estimate
$\|\nabla u_{n}\|^{s\frac{2\theta d}{d+2\theta}}_{s\frac{2\theta d}{d+2\theta}}\leq C_{8}\biggl(\|\nabla u_{n}\|_{2}^{2}+\|\nabla u_{n}\|_{2}^{2+s\frac{2\theta d}{d+2\theta}}\biggr)+C_{4}\|f-\mu u_{n}\|^{s\frac{2\theta d}{d+2\theta}}_{s\frac{2\theta d}{d+2\theta}}.$
Step 2. Let us show that in the right-hand side of the estimate of Step 1 we have $\sup_{n}\|\nabla u_{n}\|^{2}_{2}<\infty$. To this end, we multiply $(\mu-\Delta+b_{n}\cdot\nabla)u_{n}=f$ by $u_{n}$ and integrate, obtaining $\mu\|u_{n}\|_{2}^{2}+\|\nabla u_{n}\|_{2}^{2}+\langle b_{n}\cdot\nabla u_{n},u_{n}\rangle=\langle f,u_{n}\rangle,$ where
$\langle b_{n}\cdot\nabla u_{n},u_{n}\rangle=-\frac{1}{2}\langle{\rm div\,}b_{n},u_{n}^{2}\rangle\geq-\frac{1}{2}\langle({\rm div\,}b_{n})_{+},u_{n}^{2}\rangle.$
Hence, by our form-boundedness assumption on $({\rm div\,}b_{n})_{+}$,
$\left(\mu-\frac{c_{\delta_{+}}}{2}\right)\|u_{n}\|_{2}^{2}+\bigg(1-\frac{\delta_{+}}{2}\bigg)\|\nabla u_{n}\|_{2}^{2}\leq\langle f,u_{n}\rangle.$ (16.6)
So, applying the quadratic inequality in the right-hand side, we arrive at $(\mu-\frac{c_{\delta_{+}}}{2}-\frac{1}{2})\|u_{n}\|_{2}^{2}+(1-\frac{\delta_{+}}{2})\|\nabla u_{n}\|_{2}^{2}\leq\frac{1}{2}\|f\|_{2}^{2}$. Since $\delta_{+}<2$, $\sup_{n}\|\nabla u_{n}\|_{2}^{2}<\infty$ for $\mu\geq\mu_{0}:=\frac{c_{\delta_{+}}}{2}+\frac{1}{2}$.
Step 3. Next, $\|u_{n}\|_{2}\leq C\|f\|_{2}$ and a priori bound $\|u_{n}\|_{\infty}\leq\|f\|_{\infty}$ yield $\sup_{n}\|u_{n}\|_{s\frac{2\theta d}{d+2\theta}}<\infty$. Hence $\sup_{n}\|f-\mu u_{n}\|^{2}_{s\frac{2\theta d}{d+2\theta}}<\infty$.
Steps 1-3 give us the sought gradient bound $\sup_{n}\|\nabla u_{n}\|_{s\frac{2\theta d}{d+2\theta}}<\infty$, which thus ends the proof. ∎
17. Proof of Theorem 7.1
Proof of (i)
We modify the proof of Theorem 5.1, i.e. we verify conditions of the Trotter’s approximation theorem, but now for Feller generators
$\Lambda(a_{n},b_{n},q_{m}):=-a_{n}\cdot\nabla^{2}+(b_{n}+q_{m})\cdot\nabla,\quad D\big(\Lambda(a_{n},b_{n},q_{m})\big)=(1-\Delta)^{-1}C_{\infty}.$
Condition $1^{\circ}$) of Trotter’s theorem is obvious.
Let us verify conditions $2^{\circ}$) and $3^{\circ}$). To this end, we note that Propositions 8.2, 8.4 and 8.1, i.e. a priori Hölder continuity of solutions, separation and embedding properties, are still valid for operators $\Lambda(a_{m},b_{n},q_{m})$ (in fact, under more general condition $\delta<4\xi^{2}$) since we can put these operators in divergence form
$\Lambda(a_{n},b_{n},q_{m})=-\nabla\cdot a_{n}\cdot\nabla+(\tilde{b}_{n}+q_{m})\cdot\nabla,\quad\tilde{b}_{n}=\nabla a_{n}+b_{n}\in\mathbf{F}_{\delta},$
so De Giorgi’s method applies.
Remark 17.1.
Condition $\delta<4\xi^{2}$ is seen from the following calculation, which we have to repeat several times (also, with the cutoff function $\eta$) when extending Propositions 8.1, 8.2, 8.4 to include matrix fields $a_{n}$. We mutiply elliptic equation $(\mu-\nabla\cdot a_{n}\cdot\nabla+(\tilde{b}_{n}+q_{m})\cdot\nabla)u=0$ by $u^{p-1}\geq 0$ and integrate by parts, obtaining, after taking into accout ${\rm div\,}q_{m}=0$,
$\mu\langle u^{p}\rangle+\frac{4(p-1)}{p^{2}}\langle a_{n}\cdot\nabla u^{\frac{p}{2}},\nabla u^{\frac{p}{2}}\rangle+\frac{2}{p}\langle\tilde{b}_{n}\cdot\nabla u^{\frac{p}{2}},u^{\frac{p}{2}}\rangle=0.$
Since $a_{n}\in H_{\xi}$,
$\mu\langle u^{p}\rangle+\frac{4(p-1)}{p^{2}}\xi\langle|\nabla u^{\frac{p}{2}}|^{2}\rangle+\frac{2}{p}\langle\tilde{b}_{n}\cdot\nabla u^{\frac{p}{2}},u^{\frac{p}{2}}\rangle=0,$
Applying the quadratic inequality in the last term, we arrive at
$\mu\langle u^{p}\rangle+\frac{4(p-1)}{p^{2}}\xi\langle|\nabla u^{\frac{p}{2}}|^{2}\rangle\leq 2\biggl(\alpha\langle|\tilde{b}_{n}|^{2},v^{p}\rangle+\frac{1}{4\alpha}\langle|\nabla v^{\frac{p}{2}}|^{2}\rangle\biggr).$
Now, using $\tilde{b}_{n}\in\mathbf{F}_{\delta}$ and selecting $\alpha=\frac{1}{2\sqrt{\delta}}$, we obtain
$\mu\langle u^{p}\rangle+\biggl[\frac{4(p-1)}{p^{2}}\xi-\frac{2}{p}\sqrt{\delta}\biggr]\langle|\nabla u^{\frac{p}{2}}|^{2}\rangle\leq 0.$
So, $\delta<4\xi^{2}$ is exactly the condition that ensures that $\frac{4(p-1)}{p^{2}}\xi-\frac{2}{p}\sqrt{\delta}>0$ for some finite $p\geq 2$ and hence gives us an energy inequality; that is, we need $p>\frac{2}{2-\xi^{-1}\sqrt{\delta}}$.
Proposition 8.3 is replaced by a simpler convergence result
$u:=L^{2}_{{\rm loc}}\mbox{-}\lim_{n}\lim_{m}u_{n,m},$ (17.1)
where $u_{n,m}=(\mu+\Lambda(a_{n},b_{n},q_{m}))^{-1}f$, $f\in C_{c}^{\infty}$. (As is explained in the proof of Theorem 5.1(i), we need convergence in some topology to establish the approximation uniqueness.) Let us prove (17.1). Due to our more restrictive assumption $\delta<\xi^{2}$ we can work in $L^{2}$ rather than $L^{p}$. By the Steps 1-3 in the proof of Proposition 8.3, which extend easily to $a_{n}$ for each fixed $n$, the limit $u_{n}:=L^{2}\mbox{-}\lim_{m}u_{n,m}$ exists and satisfies the identity
$\mu\langle u_{n},\varphi\rangle-\langle a_{n}\cdot\nabla u_{n},\nabla\varphi\rangle+\langle(\nabla a_{n}+b_{n}+q)\cdot\nabla u_{n},\varphi\rangle=\langle f,\varphi\rangle,\quad\varphi\in C_{c}^{\infty}$ (17.2)
with $\mu$ independent of $n$. The standard energy inequality argument (cf. Section 11) and the compensated compactness estimate of Proposition 2.1 allow us to extend (17.2) to test functions $\varphi\in W^{1,2}$, i.e. $u_{n}$ is the standard weak solution to the elliptic equation $(\mu-\nabla\cdot a_{n}\cdot\nabla+(\nabla a_{n}+b_{n}+q)\cdot\nabla)u_{n}=f$ in $L^{2}$. Now, the convergence $\nabla a_{n}+b_{n}\rightarrow\nabla a+b$ in $L^{2}_{{\rm loc}}$ and the convergence $a_{n}\rightarrow a$ a.e. on $\mathbb{R}^{d}$, applied in the standard weak compactness argument in $L^{2}$, give us, via the uniqueness of the weak solution in $L^{2}$, the sought convergence $u_{n}\rightarrow u$ in $L^{2}_{{\rm loc}}$, where $u$ satisfies (17.2) with $a$, $b$ instead of $a_{n}$, $b_{n}$, moreover, this identity extends in the same way to $\varphi\in W^{1,2}$, so $u$ is the standard weak solution $(\mu-\nabla\cdot a\cdot\nabla+(\nabla a+b+q)\cdot\nabla)u=f$ in $L^{2}$. In the case $q=0$, this is essentially how the divergence form operator $\nabla\cdot a\cdot\nabla+b\cdot\nabla$ with form-bounded $b$ was treated in [KS3, Theorem 4.3]
Now, armed with the above analogues of Propositions 8.2, 8.3 and 8.1 for $\Lambda(a_{m},b_{n},q_{m})$, we verify condition $2^{\circ}$) of Trotter’s theorem in the same way as in the proof of Theorem 5.1.
Condition $3^{\circ}$) requires a comment. Fix $g\in C_{c}^{\infty}$. By the resolvent identity,
$\mu(\mu+\Lambda(a_{m},b_{n},q_{m}))^{-1}g-\mu(\mu-\Delta)^{-1}g$ $=\mu(\mu+\Lambda(a_{m},b_{n},q_{m}))^{-1}(a_{m}-I)\cdot\nabla^{2}(\mu-\Delta)^{-1}g$
$+\mu(\mu+\Lambda(a_{m},b_{n},q_{m}))^{-1}(b_{n}+q_{m})\cdot\nabla(\mu-\Delta)^{-1}g.$
Since $\mu(\mu-\Delta)^{-1}g\rightarrow g$ uniformly on $\mathbb{R}^{d}$ as $\mu\rightarrow\infty$, it suffices to show convergence
$\|(\mu+\Lambda(a_{m},b_{n},q_{m}))^{-1}(a_{m}-I)$ $\cdot\mu(\mu-\Delta)^{-1}\nabla^{2}g\|_{\infty}\rightarrow 0$ (17.3)
and
$\|(\mu+\Lambda(a_{m},b_{n},q_{m}))^{-1}(b_{n}+q_{m})$ $\cdot\mu(\mu-\Delta)^{-1}\nabla g\|_{\infty}\rightarrow 0$ (17.4)
$\text{ as }\mu\rightarrow\infty\quad\text{ uniformly in $n$, $m$}.$
To prove (17.4) we argue as in the proof of Theorem 5.1(i) and apply the discussed above analogue of Proposition 8.1 to $w_{n,m}:=(\mu+\Lambda(a_{m},b_{n},q_{m}))^{-1}(b^{i}_{n}+q^{i}_{m})f$ with $f$ chosen as $f:=\mu(\mu-\Delta)^{-1}\nabla_{i}g$. (Let us note in passing that Proposition 8.1 is valid for $b_{n}^{i}$ and $q_{n}^{i}$ in the RHS of the equation for $w_{n,m}$ replaced by the $i$-th components of general vector fields in $\mathbf{F}_{\nu}$, $\nu<\infty$, and $\mathbf{BMO}^{-1}$, i.e. the proof of Proposition 8.1 does not exploit any cancellations between the RHS of the equation and the drift term.)
The proof of (17.3) is even easier. Indeed, we can apply a straightforward analogue of Proposition 8.1 to $w_{n,m}:=(\mu+\Lambda(a_{m},b_{n},q_{m}))^{-1}f$ with bounded $f$ chosen as $f:=((a_{m})_{ij}-\delta_{ij})\mu(\mu-\Delta)^{-1}\nabla_{i}\nabla_{j}g$, use the uniform in $m$ boundedness of $a_{m}$ on $\mathbb{R}^{d}$ and then argue as in the proof of Theorem 5.1(i).
Proof of (ii)
The proof of the relaxed approximation uniqueness essentially does not change. Since $\delta<\xi^{2}$, we continue to work in the standard setting of weak solutions in $L^{2}$. We get an extra term in Step 2: the difference $h_{n}=u_{n}-u$ satisfies
$\mu\langle h_{n},\varphi\rangle+\langle a_{n}\cdot\nabla h_{n},\nabla\varphi\rangle+\langle b_{n}\cdot\nabla h_{n},\varphi\rangle+\langle q\cdot\nabla h_{n},\varphi\rangle=\langle\nabla\cdot(a_{m}-a)\cdot\nabla u,\nabla\varphi\rangle+\langle(b-b_{n})\cdot\nabla u,\varphi\rangle$
for all $\varphi\in W^{1,2}$, for all $\mu$ greater than some $\mu_{0}$ independent of $n$. Taking $\varphi=h_{n}\rho$ and repeating the proof of the energy inequality of Proposition 11.1(ii) for $s=2$, we obtain
$(\mu-\mu_{0})\langle|h_{n}|^{2}\rho\rangle+C_{1}\langle|\nabla h_{n}|^{2}\rho\rangle\leq\langle\nabla\cdot(a_{n}-a)\cdot\nabla u,\nabla(h_{n}\rho)\rangle+\langle(b-b_{n})\cdot\nabla u,h_{n}\rho\rangle,$
where the last term tends to zero as $n\rightarrow\infty$ by the argument in the proof of Theorem 5.1(ii). The term $\langle\nabla\cdot(a_{n}-a)\cdot\nabla u,(\nabla h_{n})\rho+h_{n}\nabla\rho\rangle\rightarrow 0$ tends to zero by an even simpler argument:
$|\langle\nabla\cdot(a_{n}-a)\cdot\nabla u,(\nabla h_{n})\rho\rangle|\leq\||a_{n}-a||\nabla u|\sqrt{\rho}\|_{2}\|(\nabla h_{n})\sqrt{\rho}\|_{2},$
where the second multiple is bounded uniformly in $n$ due to the energy inequality, and the first multiple tends to 0 by the Dominated convergence theorem (since $|\nabla u|\sqrt{\rho}\in L^{2}$, also by the energy inequality). (The proof that $\langle\nabla\cdot(a_{n}-a)\cdot\nabla u,h_{n}\nabla\rho\rangle\rightarrow 0$ as $n\rightarrow\infty$ is easier since $|\nabla\rho|$ is majorated by $\rho$.)
Proof of (iii)
The proof repeats the proof of Theorem 5.1(iii).
Proof of (iv)
We obtain in the same way as in the proof of Theorem 5.1(iv)
$X_{t}^{n,m}(\omega^{\prime})\rightarrow X_{t}(\omega^{\prime}),\quad t\geq 0,\quad\omega^{\prime}\in\Omega,$ (17.5)
where
$X_{t}^{n,m}=X_{0}-\int_{0}^{t}(b_{n}(X_{s}^{n,m})+q_{m}(X_{s}^{n,m}))ds+\int_{0}^{t}\sigma_{n}(X_{s}^{n,m})dB_{s},\quad n,m=1,2,\dots$
We only need to supplement the proof of Theorem 5.1(v) by the convergence
[…]
Form now on, we drop index $ij$ to lighten notations. Since $a_{n}\rightarrow a$ a.e. on $\mathbb{R}^{d}$, we have convergence of their square roots: $\sigma_{n}\rightarrow\sigma$ a.e. By Itô’s isometry, our task is to show that
$\mathbf{E}^{\prime}\int_{0}^{t}|\sigma_{n}(X_{s}^{n,m})-\sigma(X_{s})|^{2}ds\rightarrow 0\quad\text{ as }n,m\rightarrow\infty.$
In turn, this convergence follows from:
$\mathbf{E}^{\prime}\int_{0}^{t}|\sigma_{n}(X_{s}^{n_{0},m_{0}})-\sigma(X^{n_{0},m_{0}}_{s})|^{2}ds\rightarrow 0\text{ as }n\rightarrow\infty\text{ uniformly in $n_{0}$, $m_{0}\geq 1$},$
and
$\mathbf{E}^{\prime}\int_{0}^{t}|\sigma(X_{s}^{n,m})-\sigma(X_{s})|^{2}ds\rightarrow 0\quad\text{ as }n,m\rightarrow\infty.$
The latter is immediate from (17.5) via the Dominated convergence theorem, and the former follows right away from the strong Feller property of the resolvents (assertion (vii)):
$\mathbf{E}^{\prime}\int_{0}^{t}|\sigma_{n}(X_{s}^{n_{0},m_{0}})-\sigma(X^{n_{0},m_{0}}_{s})|^{2}ds$ $=\int_{\mathbb{R}^{d}}\nu_{0}(dx)\mathbb{E}_{\mathbb{P}^{n_{0},m_{0}}_{x}}\int_{0}^{t}|\sigma_{n}(\omega_{s})-\sigma(\omega_{s})|^{2}ds$
$=\int_{\mathbb{R}^{d}}\int_{0}^{t}\bigl(e^{-s\Lambda(a_{n_{0}},b_{n_{0}},q_{m_{0}})}|\sigma_{n}-\sigma|^{2}\bigr)(x)ds\nu_{0}(x)dx$
$=\int_{\mathbb{R}^{d}}\int_{0}^{t}e^{\mu s}e^{-\mu s}\bigl(e^{-s\Lambda}|\sigma_{n}-\sigma|^{2}\bigr)(x)ds\nu_{0}(x)dx$
$\leq e^{\mu t}\int_{\mathbb{R}^{d}}\int_{0}^{\infty}e^{-\mu s}\bigl(e^{-s\Lambda}|\sigma_{n}-\sigma|^{2}\bigr)(x)ds\nu_{0}(x)dx.$
The last term is, modulo $e^{\mu t}$, which is bounded anyway, is
$\int_{\mathbb{R}^{d}}(\mu-\Lambda)^{-1}|\sigma_{n}-\sigma|^{2}(x)\nu_{0}(x)dx\leq C\int_{\mathbb{R}^{d}\setminus B_{R}}\nu_{0}(x)dx+\|\mathbf{1}_{B_{R}}(\mu-\Lambda))^{-1}|\sigma_{n}-\sigma|^{2}\|_{\infty},$
where $C=\sup_{n}\|\sigma_{n}\|_{\infty}+\|\sigma\|_{\infty}$. The first integral can be made as small as needed by selecting $R$ sufficiently large. To estimate the second term, we invoke the strong Feller property (vii):
$\|\mathbf{1}_{B_{R}}(\mu-\Lambda))^{-1}|\sigma_{n}-\sigma|^{2}\|_{\infty}$
$\leq K\sup_{x\in\frac{1}{2}\mathbb{Z}^{d}\cap B_{R}}\biggl[\langle|\sigma_{n}-\sigma|^{2p\theta}\rho_{x}\rangle^{\frac{1}{p\theta}}+\langle|\sigma_{n}-\sigma|^{2p\theta^{\prime}}\rho_{x}\rangle^{\frac{1}{p\theta^{\prime}}}\biggr].$
It remains to apply the Dominated convergence theorem in $n$. (Strictly speaking, we are applying Feller resolvent to discontinuous functions, but since the former is a family of integral operators, a standard limiting argument addresses this.) ∎
Appendix A Weakly form-bounded drifts and Keller-Segel finite particles
In the previous sections we tested out results for SDEs with form-bounded drifts against the interacting particle system in Example 1.1. This, however, was limited to dimensions $d\geq 3$. In dimension $d=2$, which is of interest e.g. in the study of the Keller-Segel model of chemotaxis, the particle system in Example 1.1 is more difficult to handle since its drift (1.5) is not in $L^{2}_{{\rm loc}}(\mathbb{R}^{2})$ and, thus, is not form-bounded. We can address this issue, at least to some extent, by pursuing a different approach to proving weak well-posedness of SDEs. It works for substantially larger class of weakly form-bounded drifts.
Definition A.1.
A vector field $b\in[L^{1}_{{\rm loc}}]^{d}$ is said to be weakly form-bounded if there exists constant $\delta>0$ such that
$\langle|b|\varphi,\varphi\rangle\leq\delta\|(\lambda-\Delta)^{\frac{1}{4}}\varphi\|_{2}^{2},\quad\forall\,\varphi\in W^{1,2},$
for some $\lambda=\lambda_{\delta}\geq 0$. This will be abbreviated as $b\in\mathbf{F}_{\delta}^{\scriptscriptstyle\frac{1}{2}}$.
Example A.1.
1.  Morrey class $M_{1+\varepsilon}$ is a large subclass of $\mathbf{F}_{\delta}^{\scriptscriptstyle\frac{1}{2}}$ defined in elementary terms:
$\|b\|_{M_{1+\varepsilon}}:=\sup_{r>0,x\in\mathbb{R}^{d}}r\biggl(\frac{1}{|B_{r}|}\int_{B_{r}(x)}|b|^{1+\varepsilon}dx\biggr)^{\frac{1}{1+\varepsilon}}<\infty.$
The inclusion follows by D. R. Adams’ theorem [A1, Theorem 7.3]. The value of $\delta$ will be proportional to the Morrey norm, with a coefficient that depends on the constants in some fundamental inequalities of Harmonic Analysis.
2. The class of form-bounded drifts $\mathbf{F}_{\delta^{2}}$ considered in the previous section, i.e.
$\langle|b|^{2}\varphi,\varphi\rangle\leq\delta^{2}\|(-\Delta)^{\frac{1}{2}}\varphi\|_{2}^{2}+c_{\delta^{2}}\|\varphi\|_{2}^{2}\quad\biggl(=\delta^{2}\|(\lambda-\Delta)^{\frac{1}{2}}\varphi\|_{2}^{2},\;\;\lambda=\frac{c_{\delta^{2}}}{\delta^{2}}\biggr),$
is a proper subclass of $\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$. This is seen easily by appyling Heinz’ inequality. Alternatively, one can invoke the inclusion $\mathbf{F}_{\delta}~(\text{with $c_{\delta}=0$})\subset M_{2}$, see Examples 3.1 in Section 1, and, next, apply $M_{2}\subset M_{1+\varepsilon}$ if $\varepsilon<1$, and then use the previous example. That said, if one follows this path, one to a large extent loses the control over the value of the form-bound $\delta$, which is in our focus in this paper.
3. It is instructive to compare how $\mathbf{F}_{\delta}$ and $\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$ handle the weak $L^{d}$ class. Namely, for $|b|\in L^{d,\infty}$, we verify, using [KPS, Prop. 2.5, 2.6, Cor. 2.9],
$d\geq 2,\quad b\in\mathbf{F}_{\delta}^{\scriptscriptstyle\frac{1}{2}}\text{ with }\sqrt{\delta}$ $=\||b|^{\frac{1}{2}}(-\Delta)^{-\frac{1}{4}}\|_{2\rightarrow 2}\leqslant\|(|b|^{\ast})^{\frac{1}{2}}(-\Delta)^{-\frac{1}{4}}\|_{2\rightarrow 2}$
$\leqslant\biggl(\|b\|_{d,\infty}\Omega_{d}^{-\frac{1}{d}}\biggr)^{\frac{1}{2}}\||x|^{-\frac{1}{2}}(-\Delta)^{-\frac{1}{4}}\|_{2\rightarrow 2}=\biggl(\|b\|_{d,\infty}\Omega_{d}^{-\frac{1}{d}}\biggr)^{\frac{1}{2}}2^{-\frac{1}{2}}\frac{\Gamma\bigl(\frac{d-1}{4}\bigr)}{\Gamma\bigl(\frac{d+1}{4}\bigr)},$
where $\Omega_{d}=\pi^{\frac{d}{2}}\Gamma(\frac{d}{2}+1)$, and $|b|^{\ast}$ is the symmetric decreasing rearrangement of $|b|$. Similarly,
$d\geq 3,\quad b\in\mathbf{F}_{\delta_{1}}\text{ with }\sqrt{\delta_{1}}$ $=\||b|(-\Delta)^{-\frac{1}{2}}\|_{2\rightarrow 2}$
$\leqslant\|b\|_{d,\infty}\Omega_{d}^{-\frac{1}{d}}\||x|^{-1}(-\Delta)^{-\frac{1}{2}}\|_{2\rightarrow 2}$
$\leqslant\|b\|_{d,\infty}\Omega_{d}^{-\frac{1}{d}}2^{-1}\frac{\Gamma\bigl(\frac{d-2}{4}\bigr)}{\Gamma\bigl(\frac{d+2}{4}\bigr)}=\|b\|_{d,\infty}\Omega_{d}^{-\frac{1}{d}}\frac{2}{d-2}.$
In particular, using [KPS, Cor. 2.9], we have
$d\geq 2,\quad x|x|^{-2}\in\mathbf{F}^{\scriptscriptstyle\frac{1}{2}}_{\delta},\quad\sqrt{\delta}=2^{-\frac{1}{2}}\frac{\Gamma\bigl(\frac{d-1}{4}\bigr)}{\Gamma\bigl(\frac{d+1}{4}\bigr)},$ (A.1)
$d\geq 3,\quad x|x|^{-2}\in\mathbf{F}_{\delta_{1}},\quad\sqrt{\delta_{1}}=\frac{2}{d-2}.$ (A.2)
In fact, (A.2) coincides with the classical Hardy inequality.
4. An important proper subclass of $\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$ that is not contained in the Morrey class $M_{1+\varepsilon}$, regardless of how small $\varepsilon>0$ is, is the Kato class. The Kato class consists of vector fields $b\in[L^{1}_{{\rm loc}}]^{d}$ such that
$\|(\lambda-\Delta)^{-\frac{1}{2}}|b|\|_{\infty}\leq\sqrt{\delta}$
for some $\delta>0$ and $\lambda=\lambda_{\delta}\geq 0$ (The inclusion Kato class $\subset\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$ follows e.g. by duality and interpolation.) SDEs with Kato class drifts were treated by Bass-Chen [BC], who studied Brownian motion on fractals such as the Sierpinski gasket. To be more precise, they considered measure-valued $b$ with the total variation $|b|$ satisfying the Kato class condition. Moreover, when $\delta$ is sufficiently small, one obtains two-sided Gaussian bounds on the heat kernel of $\Delta+b\cdot\nabla$ [Za]. Note that the Kato class does not contain $L^{d}]^{d}$, but, for every fixed $\varepsilon>0$, it contains some vector fields that are not in $L^{1+\varepsilon}_{{\rm loc}}]^{d}$.
The proof of the next theorem is based on the resolvent representation (A.4) where we, crucially, work with the fractional powers $|b|^{\frac{1}{r}}$ for r>d1r>d-1 (so, here we take advantage of the fact that $b$ is not distributional or measure-valued).
Theorem A.1 ([K1, KS1]).
Let $d\geq 2$. Assume that $b\in\mathbf{F}^{\scriptscriptstyle 1/2}_{\delta}$ with weak form-bound $\delta$ satisfying
$\delta<m_{d}^{-1}\left\{\begin{array}[]{ll}\frac{4(d-2)}{(d-1)^{2}}&\text{ if $d\geq 4$,}\\ 1&\text{ if $d=2,3$},\end{array}\right$.
where $m_{d}:=\pi^{\frac{1}{2}}(2e)^{-\frac{1}{2}}d^{\frac{d}{2}}(d-1)^{\frac{1-d}{2}}$. The following are true:
  1. (i)
    (Weak solution to SDE) There exists a strong Markov family of probability measures $\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$ on the canonical space of continuous trajectories $\mathbf{C}$ that deliver, for every $x\in\mathbb{R}^{d}$, a weak solution to SDE
    $X_{t}=x-\int_{0}^{t}b(X_{r})dr+\sqrt{2}B_{t}.$ (A.3)
  2. (ii)
    (Feller semigroup)
    $(e^{-t\Lambda(b)}f)(x):=\mathbb{E}_{\mathbb{P}_{x}}[f(X_{t})],\quad x\in\mathbb{R}^{d},$
    is a strongly continuous Feller semigroup on $C_{\infty}$.
  3. (iii)
    (Uniqueness of weak solution to Kolmogorov backward PDE [KS5]) $v(t):=e^{-t\Lambda(b)}f$, $f\in C_{\infty}\cap L^{2}$, is the unique weak solution to Cauchy problem
    $(\partial_{t}-\Delta+b\cdot\nabla)v=0,\quad v|_{t=0}=f,$
    in the “shifted” triple of Bessel potential spaces $\mathcal{W}^{\frac{3}{2},2}\hookrightarrow\mathcal{W}^{\frac{1}{2},2}\hookrightarrow\mathcal{W}^{-\frac{1}{2},2}$. This result yields approximation uniqueness for $\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}$.
  4. (iv)
    (Another kind of approximation uniqueness) If $\{\mathbb{Q}_{x}\}_{x\in\mathbb{R}^{d}}$ is another weak solution to (A.3) such that
    $\mathbb{Q}_{x}=w{\mbox{-}}\mathcal{P}(\mathbf{C})\mbox{-}\lim_{n}\mathbb{P}_{x}(\tilde{b}_{n})\quad\text{for every $x\in\mathbb{R}^{d}$},$
    for some $\{\tilde{b}_{n}\}\subset\mathbf{F}^{\scriptscriptstyle 1/2}_{\delta_{1}}\cap[C_{b}\cap C^{\infty}]^{d}$ with $\delta<\frac{4(d-2)}{(d-1)^{2}}$ if $d\geq 4$ or $m_{d}\delta<1$ if $d=2,3$, and $\lambda_{\delta}$ independent of $n$, then $\{\mathbb{Q}_{x}\}_{x\in\mathbb{R}^{d}}=\{\mathbb{P}_{x}\}_{x\in\mathbb{R}^{d}}.$
  5. (v)
    (Elliptic gradient bounds) $u:=(\mu+\Lambda(b))^{-1}f$, $f\in C_{\infty}\cap L^{r}$, $r\in]d-1,\frac{2}{1-\sqrt{1-m_{d}\delta}}[$, satisfies
    $\|(\mu-\Delta)^{\frac{1}{2}+\frac{1}{2s}}u\|_{r}\leq K\|(\mu-\Delta)^{-\frac{1}{2}+\frac{1}{2\ell}}f\|_{r},\quad\text{ for all }1\leq\ell<r<s,$
    for all $\mu$ greater than a generic $\mu_{0}$ (cf. ([K6])). In particular, since r>d1r>d-1, we can select $s$ sufficiently close to $r$ so that by the Sobolev embedding theorem $u$ is Hölder continuous.
Remarks.
1. Replacing condition $b\in\mathbf{F}_{\delta}$ by more general condition $b\in\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$ comes at a cost. Although one can still include some diffusion coefficients, these may no longer be discontinuous (cf. [K3, Sect. 14]). Also, distributional drifts $q\in\mathbf{BMO}^{-1}$ are out of reach. It is, however, possible to consider drifts $b+q$, where $b\in\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$ and $q$ is measure-valued with total variation in the Kato class [K7].
2. In Theorem A.1, we construct the candidate for resolvent of the Feller generator a priori. It is the following formal Neumann series for
$\mu+\Lambda\supset\mu-\Delta+b\cdot\nabla,$
possibly after a modification on a measure zero set [K1]:
$(\mu+\Lambda)^{-1}f:=(\mu-\Delta)^{-1}f-(\mu-\Delta)^{-\frac{1}{2}-\frac{1}{2s}}Q_{r}(1+T_{r})^{-1}G_{r}(\mu-\Delta)^{-\frac{1}{2}+\frac{1}{2\ell}}f,$ (A.4)
where $f\in L^{r}\cap C_{\infty}$, and
$Q_{r}:=(\mu-\Delta)^{-\frac{1}{2}+\frac{1}{2s}}|b|^{\frac{1}{r^{\prime}}},\quad G_{r}$ $:=b^{\frac{1}{r}}\cdot\nabla(\mu-\Delta)^{-\frac{1}{2}-\frac{1}{2\ell}}\quad\text{ are bounded on }L^{r},$
$T_{r}$ $:=b^{\frac{1}{r}}\cdot\nabla(\mu-\Delta)^{-1}|b|^{\frac{1}{r^{\prime}}}\quad\text{ is bounded on }L^{r},$
where
$b^{\frac{1}{r}}:=|b|^{-1+\frac{1}{r}}b,\quad\ell,s\text{ satisfy }1\leq\ell<r<s,$
and, of course, one gets stronger regularity result by choosing $\ell$, $s$ close to $r$. The proof of the boundedness of $Q_{r}$, $G_{r}$ and $T_{r}$ is based on the Stroock-Varopoulos inequalities for symmetric Markov generators, i.e. this is an elliptic argument (to bound $T_{r}$, we first apply pointwise estimate $|\nabla_{x}(\mu-\Delta)^{-1}(x,y)|\leq m_{d}(\kappa_{d}\mu-\Delta)^{-\frac{1}{2}}(x,y)$ for appropriate $m_{d}$, $\kappa_{d}>0$). The smallness condition on $\delta$ in Theorem A.1 ensures $\|T_{r}\|_{r\rightarrow r}<1$, so that $(1+T_{r})^{-1}$ converges as geometric series in $L^{r}$. Earlier, similar estimates were employed in [BS, LS] to refine the $L^{2}$ theory of Schrödinger operators with the usual form-bounded potentials to an $L^{r}$ theory, which allowed the authors, for instance, to obtain additional information about the Sobolev regularity of the eigenfunctions of Schrödinger operators.
3. A priori, it is not clear why (A.4) should determine the resolvent of a strongly continuous semigroup in $L^{r}$. The latter is, in fact, true: the proof uses Hille’s theory of pseudoresolvents [K1]. When $r=2$, one can give a different proof using Lions’ variational approach, but it requires working in a quintuple of Hilbert spaces (instead of the usual triple) [KS3].
4. Having an explicit candidate for the limiting object, i.e. the Feller resolvent, greatly simplifies the approximation arguments. In Theorem 5.1, no such representation is available, so one must rely on Trotter’s approximation theorem, whose key feature is that it does not require any a priori representation of the limiting operator.
5. Although these broad assumptions on $b$ destroy the usual $L^{r}$ estimates for second-order derivatives of solution $u$ to $(\mu-\Delta+b\cdot\nabla)u=f$, one can still use (A.4) to obtain some $L^{r}$ bounds on $\nabla^{2}u$. However, either one needs restriction $r<d$ or these estimates are valid only in weighted space $L^{r}(\mathbb{R}^{d},(1+|b(x)|)^{-r+1}dx)$. (The latter follows by applying $(1+|b|)^{-\frac{1}{r^{\prime}}}(\mu-\Delta)$ to (A.4); note that with the information about the second derivatives of $u$ disappears at the points where $|b|$ is infinite, but in a controlled way, see [K3] for more detailed discussion.)
6. A straigthforward dual variant of the resolvent representation (A.4) produces, in particular, strongly continuous semigroup $e^{-t\Lambda^{\ast}}$ for the Fokker-Planck operator
$\Lambda^{\ast}\supset-\Delta-\nabla\cdot b,\quad b\in\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2},$ (A.5)
in $L^{r}$, where $r>1$ can be chose as close to $1$ as needed at expense of assuming that the weak form-bound $\delta$ is sufficiently small.
Now, let us recall the Ambrosio-Figalli-Trevisan superposition principle: one can construct a weak solution to SDE
$X_{t}=X_{0}-\int_{0}^{t}b(X_{r})dr+\sqrt{2}\int_{0}^{t}\sigma(X_{s})dB_{s}$ (A.6)
for bounded $b$ and $\sigma$, provided that
$t\mapsto\int_{\mathbb{R}^{d}}\varphi d\mu_{t}\text{ is continuous for every $\varphi\in C_{b}(\mathbb{R}^{d})$},$ (A.7)
where $\mu_{t}$ (a probability measure for each $t>0$) is a weak solution of the corresponding Fokker-Planck equation with $\mu_{0}={\rm Law\,}(X_{0})$; then one has ${\rm Law\,}(X_{t})=\mu_{t}$ for all $t>0$ [Tr]. For unbounded coefficients, the superposition principle is valid under additional integrablity condition on $b$ and $a=\sigma\sigma^{\top}$ with respect to $\mu_{t}$ due to [BRS]:
$\int_{0}^{T}\frac{|\langle b(x),x\rangle|+|a(x)|}{1+|x|^{2}}\mu_{t}(dx)dt<\infty.$ (A.8)
This principle, and the appropriate regularity results on the Fokker-Planck equations, can be used to treat SDEs with some quite singular drifts that are not covered by the results in the present paper, and some fairly degenerate diffusion coefficients, see [Gru] and references therein.
In light of what is written above, let us now consider SDE (A.6) with $\sigma=I$ (identity matrix) and $b\in\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$:
a) The dual variant of the resolvent representation (A.4) can be used to verify the hypothesis (A.8) of [BRS]; since we are interested in local singularities of $b$, let us assume that additionally $|b|\in L^{1}$. Then
LHS of (A.8) $\leq e^{\mu T}|b|(\mu+\Lambda^{\ast})\mu_{0}+T=e^{\mu T}|b|^{\frac{1}{r^{\prime}}}|b|^{\frac{1}{r}}(\mu+\Lambda^{\ast})\mu_{0}+T$
$\leq e^{\mu T}\||b|^{\frac{1}{r^{\prime}}}\|_{r^{\prime}}\||b|^{\frac{1}{r}}(\mu+\Lambda^{\ast})\mu_{0}\|_{r}+T<\infty.$
b) Next, we can employ semigroup $e^{-t\Lambda^{\ast}}$ in $L^{r}$ to verify the hypothesis (A.7) of the superposition principle (one will have to “cut tails” of $e^{-t\Lambda^{\ast}}$ at infinity, but this can be done e.g. for $b$ bounded ourside of a large ball). A priori, verifying the weak continuity in (A.7) via the strong continuity of $e^{-t\Lambda^{\ast}}$ seems to be an overkill, but recall that for operator semigroups the strong continuity is equivalent to the weak continuity.
a) and b) come at the cost of imposing conditions on the density of the law of the initial datum $X_{0}$, e.g. that it is in $L^{r}$, while, say, in Theorem A.1 we allow it to be a delta-function. In b) one can work directly in $L^{1}$ and construct there strongly continuous semigroup $e^{-t\Lambda^{\ast}}$, but at expense of imposing substantially more restrictive than $b\in\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$ Kato class condition on $b$.
Corollary A.1 (Finite particle approximation of the elliptic-parabolic Keller-Segel model).
In $\mathbb{R}^{2N}$, consider SDE
$X_{t}=x_{0}-\int_{0}^{t}b(X_{s})ds+\sqrt{2}B_{t},\quad x_{0}=(x_{0}^{1},\dots,x_{0}^{N})\in\mathbb{R}^{2N},$ (A.9)
where $B_{t}=(B_{t}^{1},\dots,B_{t}^{N})$ is a Brownian motion in $\mathbb{R}^{2N}$, and
$b_{i}(x^{1},\dots,x^{N}):=\frac{\sqrt{\kappa}}{N}\sum_{j=1,j\neq i}^{N}\frac{x^{i}-x^{j}}{|x^{i}-x^{j}|^{2}}.$ (A.10)
Then, provided that $\kappa<\frac{C}{N^{3}}$, the assertions of Theorem A.1 are valid for this particle system.
Proof.
Thus defined drift $b:\mathbb{R}^{2N}\rightarrow\mathbb{R}^{2N}$ is in $\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$. In fact, it is in the Morrey class $M_{1+\varepsilon}$, a subclass of $\mathbf{F}_{\delta}^{\scriptscriptstyle 1/2}$. To see this, it suffices to prove this inclusion for a single term
$\mathbb{R}^{2N}\ni(x^{1},\dots,x^{N})\mapsto\frac{x^{1}-x^{2}}{|x^{1}-x^{2}|^{2}}$
which, after a change of variable, reduces to proving that the scalar function $(x^{1},\dots,x^{N})\mapsto|x^{1}|^{-1}$ is in the Morrey class $M_{1+\varepsilon}$. Put $C_{r}(x)=D_{r}(x^{1})\times\dots\times D_{r}(x^{N})$ (the direct product of $N$ discs centered at $x^{i}$). We have
$\||x^{1}|^{-1}\|_{M_{1+\varepsilon}}$ $\leq c\sup_{r>0}r\biggl(\frac{1}{r^{2N}}\langle\mathbf{1}_{C_{r}(0)}|x^{1}|^{-(1+\varepsilon)}\rangle\biggr)^{\frac{1}{1+\varepsilon}}$
$=c\sup_{r>0}r\biggl(\frac{1}{r^{2}}\int_{D_{r}(0)}|x^{1}|^{-(1+\varepsilon)}dx^{1}\biggr)^{\frac{1}{1+\varepsilon}}$
$=c\sup_{r>0}r\biggl(\frac{1}{r^{2}}\int_{0}^{r}t^{-(1+\varepsilon)+1}dt\biggr)^{\frac{1}{1+\varepsilon}}=c\sup_{r>0}r\biggl(\frac{1}{r^{2}}\frac{r^{-\varepsilon+1}}{-\varepsilon+1}\biggr)^{\frac{1}{1+\varepsilon}}<\infty.$
The main, quite unacceptable drawback of Corollary A.1 is that the condition on $\kappa$ degenerates as $N\rightarrow\infty$. Fournier-Jourdain [FJ], Fournier-Tardy [FT] and Tardy [T] exploit the special form of the interaction kernel in (A.10) and establish for (A.9), among other results, weak existence and the existence of mean field limit as $N\rightarrow\infty$ for all $\kappa\in[0,16[$, where $16$ is the sticky collisions threshold for (A.9), (A.10). One can also apply the Dirichlet forms approach, see Cattiaux-Pédèches [CP]. Already the weak existence results of [FJ] are thus much stronger than Corollary A.1. Our point here, however, is different. Corollary A.1 shows that one, in fact, can reach the Keller-Segel finite particle system (A.9), (A.10) by applying results on general singular SDEs. (Fournier and Jourdain noted that, at the time of writing, the strongest known SDEs results for general singular drift did not apply (A.10). This was indeed true, but only until the preprint [KS1] appeared a few months later; unfortunately, at the time of writing [KS1] we were not aware of papers [CP, FJ].)
One advantage of Theorem A.1, compared to [CP, FJ, FT, T], is that we can easily modify the drift in Corollary A.1. For example, multiplying each interaction kernel by a function with $L^{\infty}$ norm less or equal to one does not affect the conclusion.
Appendix B Critical divergence, super-critical drift
Some results for the operator $\Delta+b\cdot\nabla$ depend only on ${\rm div\,}b$. For instance, if $({\rm div\,}b)_{+}^{1/2}\in\mathbf{F}_{\delta_{+}}$, $\delta_{+}<4$, then the solution $v$ to the Kolmogorov backward equation
$(\partial_{t}-\Delta+b\cdot\nabla)v=0,\quad v|_{t=0}=v_{0},$
satisfies, for all $\frac{2}{2-\sqrt{\delta_{+}}}<q\leq p\leq\infty$, the dispersion estimate
$\|v(t)\|_{p}\leq Ce^{\omega t}t^{-\frac{d}{2}(\frac{1}{p}-\frac{1}{q})}\|v_{0}\|_{q},\quad t>0.$
The proof is due to J. Nash, see [KS3] for details. While one needs smoothness and boundeness of $b$ and $v_{0}$ to carry out integration by parts, the constants $C$ and $\omega$ depend only on $d$ and $\delta_{+}$. They and do not depend on any integral characteristics of $b$.
However, to obtain more detailed information about the diffusion process with drift $b$, one must impose some conditions on $b$. For instance, in Theorem 6.1 we also required $b\in\mathbf{F}_{\delta}$.
Between these two types of assumptions, there are intermediate conditions, such as in assertion (i) of Theorem 6.3. In this assertion, selecting $\nu$ close to zero, one can treat $b$ that can be essentially twice more singular than the vector fields in $\mathbf{F}_{\delta}$. This is a super-critical condition on the drift in the sense of scaling. Let us recall the sub-critical/critical/super-critical classification of the spaces of vector fields. Given a vector field $b$, put $b_{\lambda}(x):=\lambda b(\lambda x)$. Let $Y$ be a translation-invariant Banach space of distribution-valued vector fields $b$ such that
$\|b_{\lambda}\|_{Y}=\lambda^{a}\|b\|_{Y}.$
Now,
$a>0$, then $Y$ is sub-critical, i.e. passing to the small scales decreases the norm,
$a=0$, then $Y$ is critical,
$\text{$a<0$, then $Y$ is super-critical}.$
In the last two cases “zooming in” does not change the norm or makes the norm larger. For example, $L^{p}$ is sub-critical, critical or super-critical according to whether $p>d$, $p=d$ or $p<d$. This classification is widely used in the study of Navier-Stokes equations. There one applies it to spaces of solutions or initial data.
The super-critical condition on $b$ appearing in assertion (i) was introduced in the work of Q.S. Zhang [Za2]. He considered the time-inhomogeneous counterpart of $|b|^{\frac{1+\nu}{2}}\in\mathbf{F}_{\delta}$, namely, $b\in[L^{1+\nu}_{{\rm loc}}(\mathbb{R}^{1+d})]^{d}$ and for a.e. $t\in\mathbb{R}$
$\langle|b(t,\cdot)|^{1+\nu}\varphi,\varphi\rangle\leq\delta\|\nabla\varphi\|_{2}^{2}+g_{\delta}(t)\|\varphi\|_{2}^{2}\qquad\forall\,\varphi\in W^{1,2}$ (B.1)
where $0\leq g_{\delta}\in L^{1}_{{\rm loc}}(\mathbb{R})$ describes how irregular $b$ can be in time. He established, among other results, local boundedness of any weak solution to the parabolic equation
[…]
provided that ${\rm div}\,b\leq 0$ and
$b\in[L^{2}_{{\rm loc}}(\mathbb{R}^{1+d})]^{d}.$ (B.2)
The last condition is satisfied if $b$ is taken to be a Leray-Hopf solution of 3D Navier-Stokes equations, which motivated [Za2].
The proof of Theorem 6.3(i) uses a tightness estimate for solutions of the approximating SDEs with bounded smooth drifts (cf. (6.12)). The proof of that estimate, in turn, uses the idea from [Za2] for handling cutoff functions in presence of $b$ satisfying (B.1).
The first result on SDEs with suprcritical divergence-free drifts belongs to X. Zhang and G. Zhao [ZZ1]. They considered
$X_{t}=x-\int_{s}^{t}b(r,X_{r})dr+\sqrt{2}(B_{t}-B_{s}),\quad t\geq s,$ (B.3)
with divergence-free drift $b$ additionally satisfying the square integrability condition (B.2), and included in the super-critical Ladyzhenskaya-Prodi-Serrin condition
$|b|\in L^{q}([0,T],L^{p}(\mathbb{R}^{d})),\quad p,q\geq 2,\quad\frac{d}{p}+\frac{2}{q}<2.$ (SLPS)
They proved that for every initial data $(s,x)\in\mathbb{R}^{1+d}$ the SDE (B.3) has a weak solution satisfying a Krylov type estimate. Moreover, using hypothesis (B.2), they proved that outside a measure zero set of $(s,x)$ one has approximation uniqueness and a.s. Markov property for these weak solutions. Consequently, the weak well-posedness result of [ZZ1] justifies the passive tracer model in the Leray-Hopf setting under the a priori assumption (SLPS). They also allow the positive part $({\rm div}\,b)_{+}$ of the divergence of $b$ to be singular, provided it satisfies condition (SLPS) (possibly with different exponents $p$, $q$). In recent paper [HZ], Z. Hao and X. Zhang extended the results in [ZZ1] to divergence-free super-critical distributional drifts.
Let us add that condition (B.2) is quite powerful. For instance, if one assumes only ${\rm div\,}b\leq 0$ and (B.2), then it is already sufficient to prove weak uniqueness results for the backward Kolmogorov equation in $L^{1}$ [GS].
It is easy to show, using Hölder’s inequality, that (SLPS) is a subclass of (B.1). It is a proper subclass. Indeed, (B.1) contains some vector fields having strong hypersurface singularities that are not covered by (SLPS).
The main focus of Theorem 6.3(i) was reaching the blow-up threshold for $\delta_{+}$ for $({\rm div}\,b)_{+}$. As a by-product, it also closes the gap (at the level of weak existence for SDE (B.3)) between the hypotheses on the drift in [ZZ1] and in [Za2].
There remains nontrivial (as it seems to us) work left to establish weak existence for SDEs whose drifts lie in an even larger class of super-critical divergence-free drifts, namely, those considered by Q.S. Zhang in [Za3]:
$\big\langle|b(t)|\log(1+|b(t)|)^{2}\varphi,\varphi\big\rangle\leq\delta\|\varphi\|_{2}^{2}+g_{\delta}(t)\|\varphi\|_{2}^{2}\qquad\forall\,\varphi\in W^{1,2}\quad\text{for a.e.\,$t\in\mathbb{R}$}.$ (B.4)
We conclude by returning to heat kernel bounds. Q.S. Zhang [Za2] and Qian-Xi [QX2] also obtained non-Gaussian upper bounds on the heat kernel of $\Delta+b\cdot\nabla$ with supercritical $b$. It would be interesting to “test” the optimality of these bounds and try to deduce the tightness estimate of [ZZ1] from them, in the same way as it was done above for $b\in\mathbf{MF}_{\delta}$ and the Gaussian upper bound. Of course, the calculations become more complicated due to the lack of a scaling invariance in these bounds.
Appendix C D.R. Adams’ estimates
The following are special cases of estimates proved by D.R. Adams. Let $V\geq 0$.
Lemma C.1 ([A1, Theorem 7.3]).
Let $0<\alpha<d$, $1<q<\infty$. Let $s>1$. If
$\sup_{x\in\mathbb{R}^{d},r>0}r^{\alpha q}\biggl(\frac{1}{|B_{r}|}\big\langle V^{s}\mathbf{1}_{B_{r}(x)}\big\rangle\biggr)^{\frac{1}{s}}<\infty,$
then, for all $\varphi\in\mathcal{S}$,
$\|V^{\frac{1}{q}}\varphi\|_{q}\leq C\|(-\Delta)^{\frac{\alpha}{2}}\varphi\|_{q}.$
Lemma C.2 ([A2]).
Let $1<p<q<\infty$, $p<d$. Then
$\|V^{\frac{1}{q}}\varphi\|_{q}\leq C\|(-\Delta)^{\frac{1}{2}}\varphi\|_{p},$
if and only if
$\sup_{x\in\mathbb{R}^{d},r>0}r^{-q\left(\frac{d}{p}-1\right)}\big\langle V\mathbf{1}_{B_{r}(x)}\big\rangle<\infty.$
Thus, Morrey class is responsible for the $L^{p}(\mathbb{R}^{d},dx)\rightarrow L^{q}(\mathbb{R}^{d},Vdx)$, $p<q$, estimates. It is only sufficient for the $L^{q}(\mathbb{R}^{d},dx)\rightarrow L^{q}(\mathbb{R}^{d},Vdx)$ estimates (already the larger Chang-Wilson-Wolff class shows that Morrey class is not necessary, see Section 3). In turn, the $L^{q}(\mathbb{R}^{d},dx)\rightarrow L^{q}(\mathbb{R}^{d},Vdx)$ estimates can be used e.g. to prove weak well-posedness of SDEs via (A.4), see Appendix A.
Appendix D Multiplicative form-boundedness and Morrey class $M_{1}$
The following result was proved in [M, Theorem 1.4.7]. We reproduce the proof for reader’s convenience.
Proposition D.1.
Let $b\in[L^{1}_{{\rm loc}}]^{d}$. Then
$\langle|b|\varphi,\varphi\rangle\leq\delta\|\nabla\varphi\|_{2}\|\varphi\|_{2}\quad\forall\,\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$ (D.1)
if and only if
$\langle|b|\mathbf{1}_{B_{r}(x)}\rangle\leq Kr^{d-1}$ (D.2)
for some constant $K$ independent of $r>0$ and $x\in\mathbb{R}^{d}$ (then $K$ is proportional to $\delta$).
In other words, $b\in\mathbf{MF}_{\delta}$ with $c_{\delta}=0$ if and only if $|b|$ belongs to the Morrey class $M_{1}$.
Proof of Proposition D.1.
Define weighted Lebesgue norm over ball $B_{r}(x)$:
$\|\varphi\|^{q}_{L^{q}(B_{r}(x),|b|)}:=\int_{B_{r}(x)}|\varphi(y)|^{q}|b(y)|dy.$
Define analogously $L^{q}(\mathbb{R}^{d},|b|)$. D.R. Adams’ Lemma C.2 will play a crucial role in the proof.
Proof of (D.2) $\Rightarrow$ (D.1). So, we assume that (D.2) holds, and prove the multiplicative inequality (D.1).
Lemma D.1.
For every $\varphi\in C(\overline{B_{r}})$,
$\|\varphi\|_{L^{2}(B_{r},|b|)}\leq C_{2}\left(r^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r})}+r^{-\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r})}\right),$ (D.3)
for constant $C_{2}$ independent of $\varphi$ or $r$.
Proof.
Any function $\varphi\in C^{\infty}(\overline{B_{1}})$ can be extended to a function in $C_{c}^{1}(B_{2})$, which we still denote, with some abuse of notation, by $\varphi$, so that
$\|\nabla\varphi\|_{L^{2}(B_{2})}\leq C_{1}\bigl(\varphi\|_{L^{2}(B_{1})}+\|\nabla\varphi\|_{L^{2}(B_{1})}\bigr),$
where $C_{1}$ does not depend on $\varphi$. After rescaling, we obtain
$\|\nabla\varphi\|_{L^{2}(B_{2r})}\leq C\left(\|\nabla\varphi\|_{L^{2}(B_{r})}+r^{-1}\|\varphi\|_{L^{2}(B_{r})}\right).$ (D.4)
Now that the integrals over $B_{r}$ do not depend on the choice of the extension of $\varphi$. The integrals over $B_{2r}$which do depend on the choice of the extension, like the one in the left-hand side of the previous inequality, will appear below, but only at the intermediate steps, and will not appear in the final estimates.
Now, put $q=2\frac{d-1}{d-2}$. Then our hypothesis $b\in M_{1}$ allows us to apply Lemma C.2 with thus selected $q$ and $p=2$. For $p=2$, we have $\|(-\Delta)^{\frac{1}{2}}u\|_{2}=\|\nabla u\|_{2}$. So,
$\biggl(\|\varphi\|_{L^{q}(B_{r},|b|)}\leq\biggr)\quad\|\varphi\|_{L^{q}(B_{2r},|b|)}\leq C\|\nabla\varphi\|_{L^{2}(B_{2r})},$ (D.5)
where, of course, it is crucial that $\varphi$ has support in the open ball $B_{2r}$, so integrating over $\mathbb{R}^{d}$ is equivalent to integrating over $B_{2r}$.
By Hölder’s inequality,
$\|\varphi\|_{L^{2}(B_{r},|b|)}\leq\langle|b|\mathbf{1}_{B_{r}}\rangle^{\frac{1}{2}-\frac{1}{q}}\|\varphi\|_{L^{q}(B_{r},|b|)},$ (D.6)
where $\frac{1}{2}-\frac{1}{q}=\frac{1}{2}\frac{1}{d-1}$.
Finally, assemblying estimates (D.6)-(D.4) and using our hypothesis (D.2) on $|b|$, we arrive at
$\|\varphi\|_{L^{2}(B_{r},|b|)}$ $\leq\langle|b|\mathbf{1}_{B_{r}}\rangle^{\frac{1}{2}-\frac{1}{q}}\|\varphi\|_{L^{q}(B_{r},|b|)}\leq CK^{\frac{1}{2}\frac{1}{d-1}}r^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{2r})}$
$\leq C_{2}\left(r^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r})}+r^{-\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r})}\right),$
i.e. we have proved (D.3). ∎
We are in position to complete the proof of implication (D.2) $\Rightarrow$ (D.1). Fix some $r_{0}>0$, and do the following for every $x\in\mathbb{R}^{d}$:
– If $r_{0}\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}\geq\|\varphi\|_{L^{2}(B_{r_{0}}(x))}$, then from (D.3)
$\|\varphi\|_{L^{2}(B_{r_{0}}(x),|b|)}$ $\leq C_{2}\left(r_{0}^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}+\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r_{0}}(x))}^{\frac{1}{2}}\right)$
$\leq C_{2}\left(r_{0}^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}+\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r_{0}}(x))}^{\frac{1}{2}}\right).$ (D.7)
- Otherwise, increase $r$ until one reaches equality $r\|\nabla\varphi\|_{L^{2}(B_{r}(x))}=\|\varphi\|_{L^{2}(B_{r}(x))}$, in which case (D.3) yields
$\|\varphi\|_{L^{2}(B_{r}(x),|b|)}$ $\leq C_{2}\left(\|\nabla\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}+\|\nabla\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\right)$
$=2C_{2}\|\nabla\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}.$ (D.8)
In any case, we obtain, for every $x\in\mathbb{R}^{d}$ and $r=r(x)$ (either $r_{0}$ or increased $r$, depending on $x$):
$\|\varphi\|_{L^{2}(B_{r}(x),|b|)}\leq 2C_{2}\bigg(\|\nabla\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}\|\varphi\|_{L^{2}(B_{r}(x))}^{\frac{1}{2}}+r_{0}^{\frac{1}{2}}\|\nabla\varphi\|_{L^{2}(B_{r_{0}}(x))}\bigg).$ (D.9)
Our goal now is to turn local estimate (D.9) into the multiplicative inequality (D.1). So, we cover $\mathrm{supp}|b|$ with balls $\{B_{r(x)}(x)\}$. Using [M, Theorem 1.2.1], i.e. a variant of the Besicovich covering theorem, one can extract at most countable subcover of finite multiplicity $M=M(d)$. We emphasize that while the covering depends on $\varphi$, the multiplicity $M$ does not. Let $\{B^{i}\}$ be this sub-cover. Squaring (D.9) (considered over each ball $B^{i}$ in the sub-cover), summing up in $i$, and using Hölder inequality, we have:
$\|\varphi\|^{2}_{L^{2}(\mathbb{R}^{d},|b|)}$ $\leq M\sum_{i}\|\varphi\|^{2}_{L^{2}(B^{i},|b|)}$
[…]
By Jensen’s inequality,
$\big(\sum_{i}\|\nabla\varphi\|_{L^{2}(B^{i})}^{2}\big)^{1/2}\leq\sum_{i}\|\nabla\varphi\|_{L^{2}(B^{i})}\leq M\|\nabla\varphi\|_{2}$
and similarly for $\|\varphi\|_{2}$, which gives us
$\|\varphi\|^{2}_{L^{2}(\mathbb{R}^{d},|b|)}\leq M^{2}C_{3}(\|\nabla\varphi\|_{2}\|\varphi\|_{2}+r_{0}\|\nabla\varphi\|^{2}_{2})$
Since $r_{0}$ was fixed arbitrarily, letting $r_{0}\downarrow 0$ yields the multiplicative inequality (D.1).
The reverse direction (D.1) $\Rightarrow$ (D.2) is easier. Choosing in (D.1) test functions $\varphi=\varphi_{r}\in C_{c}^{\infty}(B_{r}(x))$ such that $\|\nabla\varphi\|_{2}\leq cr^{\frac{d-2}{2}}$, $\|\varphi\|_{2}\leq cr^{d/2}$, we obtain $\langle|b|,\mathbf{1}_{B_{r}(x)}\rangle\leq Cr^{d-1}$, i.e. $b\in M_{1}$. This completes the proof. ∎
Appendix E Vanishing of stream matrix at infinity
Lemma E.1.
Assume that $q\in\mathbf{BMO}^{-1}$ has compact support in $B_{1}(0)$. Then we can find a stream matrix $Q=-Q^{\scriptscriptstyle\top}\in[{\rm BMO}]^{d\times d}$ for $q$, i.e. $q=\nabla Q$, that decays polynomially at infinity:
$|Q(x)|\leq C_{R}|x|^{-d+2}\quad\forall\,|x|\geq R>2.$
It suffices carry out the proof for scalar distributions. Recall that a tempered distribution $h\in\mathcal{S}^{\prime}$ belongs to ${\rm BMO}^{-1}$ if and only if
$\sup_{x\in\mathbb{R}^{d},R>0}\frac{1}{|B_{R}|}\int_{B_{R}(x)}\int_{0}^{R^{2}}|e^{t\Delta}h|^{2}dtdy<\infty.$
Now, given $h\in{\rm BMO}^{-1}$, one can find a vector field $H=(H_{j})_{j=1}^{d}\in[{\rm BMO}]^{d}$ such that
$h={\rm div\,}H$
by arguing as follows (see [KT]). Put
$h_{kj}:=\nabla_{k}\nabla_{j}(-\Delta)^{-1}h.$
By [KT, Lemma 4.1], $\|h_{kj}\|_{{\rm BMO}^{-1}}\leq\|h\|_{{\rm BMO}^{-1}}$. For each fixed $1\leq j\leq d$, we have $\nabla_{r}h_{kj}=\nabla_{k}h_{rj}$ for all $1\leq k,r\leq d$, i.e. $h_{\cdot j}$ is curl-free. Therefore, there exists $H_{j}$ such that $\nabla H_{j}=h_{\cdot j}$, e.g. take
$H_{j}=(-\Delta)^{-1}{\rm div\,}h_{\cdot j}.$ (E.1)
This $H_{j}\in{\rm BMO}$ by the Carleson’s characterization of $\rm BMO$} (Section 2). We have $\nabla_{j}H_{j}=h_{jj}$, hence
${\rm div\,}H\equiv\sum_{j=1}^{d}\nabla_{j}H_{j}=\sum_{k=1}^{d}h_{jj}=h.$
Proof of Lemma E.1.
Assume that $h\in{\rm BMO}^{-1}$ has compact support in $B_{1}(0)$. We will show that, by following the above procedure, we obtain a “primitive” vector field $H$ for $h$ which decays at infinity as $|x|^{-d+2}$. Indeed, since $h$ has compact support, $h_{kj}(x)=O((1+|x|)^{-d})$ as $|x|\rightarrow\infty$, and so ${\rm div\,}h_{\cdot j}(x)=O((1+|x|)^{-d-1})$. Therefore, by (E.1),
$|H_{j}|\leq C(-\Delta)^{-1}(1+|\cdot|)^{-d-1}.$
Invoking the Sobolev embedding property of $(-\Delta)^{-1}$, we obtain that
$(-\Delta)^{-1}(1+|\cdot|)^{-d-1}\in L^{\frac{d}{d-2}}.$
So, given that $(-\Delta)^{-1}(1+|\cdot|)^{-d-1}$ is a bounded rotationally-invariant function, we obtain the sought polynomial rate of decay of $(-\Delta)^{-1}(1+|\cdot|)^{-d-1}$ and hence of $H_{j}$. (It is also not difficult to estimate the rate of decay of $(-\Delta)^{-1}(1+|\cdot|)^{-d-1}$ directly.) ∎